A Bibliography of Publications by, and about, Ingrid Daubechies

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

06 November 2024 Version 1.08

Abstract

This bibliography records publications of Ingrid Daubechies.

Title word cross-reference

$1.5M [Duk16]. 2 [LK91, LLC08, LQLC10a, LLC11, MMC^{+}13, RB98a, YLB^{+}15]. $59.95 [Lun92]. $69.95 [Lun92]. [0, 3] [Pol92a, Pol92b, Pol92c]. C [ZCY16]. db8 [Kar03, Kar04]. δ [XZCM06]. ℓ_{1} [LNDD06, SP17]. J [WWD^{+}15]. l^{∞} [AM08]. M [BDS98, HSZ97, Sun99]. n [LP01, SZH97]. {ψ_{jk}^{(2)}}_{j,k∈Z} [Red15]. q [BAE11].

-analogue [BAE11]. -ary [SZH97]. -band [BDS98, HSZ97, Sun99]. -D [LK91, LLC08, LLC11, MMC^{+}13, YLB^{+}15]. -dimensional [LP01]. -means [ZCY16]. -norm [LNDD06, LNDD07]. -sequence [XZCM06]. -stability [AM08].

0-12-174590-2 [Lun92]. 0-86720-225-4 [Lun92].

1 [XTZ10]. 12-lead [WWD^{+}15]. 161 [DL01b]. 1992 [Ano93]. 1D [XLY^{+}23].

2 [DAR13, DA15]. 2011 [Ano11a]. 2012 [DKRS12]. 2015 [Ano15, DKRS15].

3-band [WP05].

4-tap [MMC^{+}13]. 44 [RRPT09].

5/3 [Ara13].

6-tap [MMC^{+}13]. 650-year-old [Dau16].

8-Tap [CKE17].

9/7 [TDK15]. 93f [DL01b]. 97g [CD97a].

A/D [CD00, CDL07, DDGV02, DDGV06]. Abstracts [DKRS12, DKRS15]. Academic [Lun92]. Academy [Ano98]. Accelerated [DFL07, DFL08]. accuracy [CD00, CDL07, KLC05]. accurate [BGS14]. acoustic [Zha00]. Acquisitions [CYG^{+}17]. AdaBoost [RDS04a, RDS04b, RSD07b]. Adaptive [DP02, HDH16]. addendum [DL01b]. Address [Dau15]. Adds [Ano98]. Advances [TSS04]. affect [DH95]. Affine [Dau94a, DKP87, TB94]. Air [YKIK04]. al [Wan01]. Alberto [DM99]. Algebra [DL01b]. algebraic [Kla97, MMC^{+}13]. algebras [AD79c]. Algorithm [STAV09, VD15, ZGSD04, BGS14, DDD03, DDD04, Dur96, LL00, LKC05, VD17, Wan07, ZCY16, ZGSD06]. algorithme [Dur96]. Algorithms [BLS^{+}11a, BLS^{+}11b, Dau93a, Nie99a, Nie12, AM08, CDJV93, Dur93, RSD07a, RSD08]. Aligning [BPG^{+}15]. Alignment [HD06]. also [Ano14]. Altarpiece [CRG^{+}13, PPR^{+}15, RCP^{+}11]. Alternatives [HD06]. amas [APD06]. American [Dau93b]. Among [Ano93]. Analog [DY06]. Analog-to-Digital [DY06]. analogue [BAE11]. Analysis [Ano08, CHT98, Dau97b, DS15a, DS15b, Grü92, Lun92, P15, RRD05, RSD07a, RSD08, Sin13, WPS^{+}14, YLB^{+}15, ZGSD04, Abo94, BDV00, CT15, CDJV93, CDDD03, Dau87, Dau89b, Dau90b, DMW92, Dau98e, DG99, Dau06c, DRT^{+}09, DKRS12, DKRS15, DKRS18, DKRS21, DVDD98, Gao14, HDH16, JPB^{+}09, JHB^{+}08, KAB11, KLC05, Lin97a, LQLC10a, NKM12, SS96, WPS^{+}13, Yos15, ZGSD06]. analysis/synthesis [Abo94]. Analytic [Dau80c, Yos10]. analyzing [AIK10a, AIK10b, XLY^{+}23, lZqJmTjZ08]. anatomical [BLS^{+}11a, BLS^{+}11b]. angle [PR05]. angles [AD83]. answers [WFD11]. Antonio [Dau93b]. any [DH04]. Appl. [DL01b]. Application [CWC04, DT14, DMC^{+}17, DC11, GLG94, yGjZsC11, GYD^{+}18, Koz06, Lun92, RRPT09, YLB^{+}15, APD06, Dau78a, Dau78b, DM99, DMC^{+}16b, Fin04a, Fin04b, RRD12, SLBD11a, SLBD11b]. Applications [Ano11e, CKE17, Dau08, RV09b, RBC^{+}92, Wah11, DP87, Dau95d, Dau95c, Dau98c, Dau98d, DG99, HDH16, Lin98, Yos10, Dau93a, Lun92, Lun92]. Applied [DKRS12, DKRS15, DKRS18, DKRS21, AHK13]. Approach [BPG^{+}15, HN15, PZC^{+}12, PZC^{+}15, RGMD15a, STAV09, AM08, CXS04, Dau88b, DP88, DDGV02, HN17, RGMD15b, WP11, XZCM06]. approaches [PBGD13]. Approximate [Tay08]. Approximating [DD03]. approximation [BD03, CDDD01, CDGO02, DRS04, DKRS12, DKRS15, DDF^{+}22, DDD^{+}23, EGL11, EGL13, Koz06, UD97]. April [SC02]. Arbitrary [CD93c, DD03]. arc [RSD07b]. arc-gv [RSD07b]. architecture [LK91]. architectures [MMC^{+}13]. arithmetic [Bon16]. Art [ACD^{+}13, DMC^{+}17, Dau16, DMC^{+}16b, YLB^{+}15, YDC^{+}14, YMH^{+}16]. Artifacts [YCF^{+}16]. artist [JHB^{+}08]. ary [SZH97]. Aspects [LPD11, LPD13]. Assessment [GYD^{+}18]. associated [Kar10, Lai95]. Astronomy [KK10, KKT10]. asymmetry [Dor94]. Asymptotic [BDS98, KLR95a, LS00, SS96, Sun99]. Asymptotics [Nov02a, Nov02b, Nov02c, SS98, Tem96, Tem97]. Atherosclerosis [OWW^{+}16]. attachment [DDK05]. auditory [DM96]. August [DKRS15]. Authentication [SM12]. author [Dau16]. automata [CD97b]. Automated [BPG^{+}15, GYD^{+}18, PBGD13, WWD^{+}15]. Automatic [FGDB17]. automatically [BLS^{+}11a, BLS^{+}11b]. autonomous [DDK05]. Award [Keh13]. awarded [Ano12]. Awards [Duk16].

B [LD16]. B-spline [LD16]. background [CK96, PAHD05]. balance [Gao14]. banco [VP13]. band [BDS98, HSZ97, Sun99, WP05]. Bandlimited [DS15a, DS15b, CDL02, DD03]. bank [VP13]. banking [AAI13]. banks [EGL11, EGL13, SK12]. Bargmann [Cob01, DG88]. Bartlett [Lun92]. Based [BBN^{+}10b, CWC04, yGjZsC11, XYD16, XYD18, Ara13, BDV00, CXS04, CT15, CDGO02, DM96, DHRS03, DT04, DAR13, Gao14, HŚ02b, HŚ02a, Jam96, KTJ09, LQLC10b, LLC11, LD16, PM13a, PR05, PM96, RM95, RSD04, SK12, Sud16, TB94, UD97, WWFW98, WP11, XLY^{+}23, ZCY16]. Bases [Dau88a, Dau90a, Kai10, YD16, YGLD16, YD17, YGLD17, CDF92, CD92, CD93a, CD93b, CD93d, Dau89a, Dau93c, Dau93d, Dau94b, Dau06b, KLC05, LW09, YD18]. basic [AM08]. Basis [DJJ91b, Lin97a, CT15, DJJ91a, GNG^{+}08a, GNG^{+}08b, LKC05, LD16, NG05, NG06]. Bayesian [CYV^{+}13a, CYV^{+}13b, PZC^{+}12, PZC^{+}15, RRD05, RRD12]. BBVA [Keh13]. be [Dau16]. beam [DMV09, WP11]. Beamlet [CWC04]. Beamlets [WC02]. behavior [BDS98, KLR95a, RDS04a, RDS04b]. bekende [Gre11]. bending [Koz06]. Bernstein [Nov02a, Nov02b, Nov02c]. Beta [DDGV02]. Beth [ADG^{+}24, Lun92]. Better [Dau95a, CD93d]. between [AD79c, AD83, AADL13, AIK10a, Dau11, JDB^{+}14, AIK10b]. Beylkin [Lun92]. bidimensional [CD93a, CD93b]. BigDFT [GVO^{+}11]. Biological [GYD^{+}18]. biologically [FGDB17]. Biomedical [CKE17]. Biorthogonal [CDF92, CD93c, KM12, CD92, Dau94b, VBU05b]. birth [Gre11]. bit [CD00, CDL07]. bit-rate [CD00]. Bivariate [ACV01]. block [HŚ02b, HŚ02a]. block-oriented [HŚ02b, HŚ02a]. Book [Bat93a, C.93, Dau93a, Gri95, Grü92, Hei92, Hei93, Lun92]. Boosting [RDS04c, RSD04, RSD07a, RSD08]. Boston [Lun92]. boundary [XZCM06]. bounded [Dau80b]. Boundedness [CS99]. boxes [GB95b, GB95a]. Brace [Lun92]. Brain [Ano08, DRT^{+}09]. Brainbow [KLT^{+}10]. brushstrokes [JHB^{+}08]. Brussels [Ano05]. Burgemeester [Gre11]. Butterworth [Abo94]. Butterworth/Daubechies [Abo94]. BV [CDDD03]. BVPs [Fin04a, Fin04b].

CA [Lun92]. calculation

[LXDS11, NG05, NG06]. Calculations [RGMD15a, GNG^{+}08a, GNG^{+}08b, GVO^{+}11, RGMD15b]. Calderón [DM99]. can [Dau16, Sta15]. canonical [Dau80a, DH02]. Canvas [CYG^{+}17, YLB^{+}15]. Capturing [CDD^{+}12]. Cardiovascular [OWW^{+}16]. cares [BBJ^{+}09]. Carlo [PS95]. carrying [KLT^{+}10]. cascade [Dur93, Dur96]. cascading [Dur96]. case [AM08, CRG^{+}13]. cassette [KLT^{+}10]. Cauchy [LL00, PM13b]. CDF [Ara13]. celebration [Dau95d]. change [Sta15]. Chantal [Coo11]. characteristics [HŚ02b, HŚ02a]. Characterization [Qix12, Yan12, AD79b, LP01]. characterize [PCR^{+}11]. Charles [Grü92, Lun92]. children [EBJ^{+}14]. Chui [Lun92, Grü92]. Chyzak [Wan01]. CIRM [Ano16b]. Citra [SG13]. class [HŚ02b, HŚ02a, HN17, LW09]. classificacão [VP13]. Classification [Ano08, BB07, VP13]. clicks [LG08]. clustering [DAR13, DA15, ZCY16]. clusters [APD06, PAHD04, PAHD05]. CMB [PAHD04]. coarsely [DD03]. coded [MDSW92]. coding [ABMD90, ABMD92, BDV00, CD92, CDGO02, MDSW92]. Coefficients [ADGT17, BBN^{+}10a, ADGT16, ČF04, EGL11, EGL13, LNDD06, LNDD07, OMOE14a, OMOE14b, Red15, Str92, Wan07]. Cohen [Ara13]. Coherent [Dau80a, Dau87, Dau91, Dau94a]. Coiflets [SYSP11, SYSP12]. Coifman [Lun92]. collocation [MD06]. Colon [MMN^{+}11]. Color [NY15, ST15]. Combination [BBL^{+}11]. combined [AHK13]. combining [CDGO02]. come [Dau96]. coming [Ano05]. Communities [OWW^{+}16]. Commutation [DGS01]. Compactly [Dau88a, CDF92, CD93d, Dau93c, Dau06b, GB95b, GB95a, PKG13, WP05]. compactly-supported [PKG13]. Companion [GBGL08]. Comparative [GYD^{+}18, P15]. Comparing [BPG^{+}15, BBL^{+}11, SYSP11, SYSP12, LD11a, LD11b]. Comparison [BB07, CVN^{+}13, Gao14, LD09, SHN10, AIK10a, AIK10b]. Complex [CLG04, GLG94, LM95, Lin98, TB94, CGB^{+}15, KTJ09, KKJ^{+}10, LM94, Lin97b, PM15, XTZ10]. Complexities [CL05]. Complexity [Wah11]. Component [P15, RRD05, DRT^{+}09]. Compression [SC00, SC02, TDK15, Wel99b, Ara13, AC14, CDSY97, DVDD98, KAB11, RB98b]. computable [Tas00]. computation [ČF04, CHXL06, LKC05]. Computational [MYN07, LPD11, LPD13]. computations [RL97]. Computerized [JHB^{+}08]. Computing [Daa93, Old92, PSB^{+}16, Wan07, WP11]. ConceFT [DWtW15, DWtW16]. Concentration [DWtW15, DWtW16]. concepts [DTV08]. condition [AD79d]. conditionally [CDDL21]. Conference [SC00, SC02, TSS04]. Conformal [yGjZsC11, LD11a, LD11b, LPD11, LPD13]. conjecture [Wan01]. connected [CDDL21]. Connecting [Kar03, Kar04]. Connection [Dau90a, AD79c, Dau83a, Dau89a, Wan07]. connectivity [DDK05]. cons [CD02]. conservation [PPR^{+}15]. conserve [AD83]. constants [Nov98]. Constrained [RGMD15a, RGMD15b]. Constraint [Tas00, DDD03, DDD04]. Constraint-selected [Tas00]. Constraints [DFL07, DFL08, CVN^{+}13, DTV07]. construct [HN17]. Constructing [DK82, WP05]. Construction [CWC04, GBM09a, GBM09b, HN15, LW09, WLW06, XTZ10, Dur96, Nov02a, Nov02b]. Constructions [YD16, YD17, DHRS03, YD18]. Content [WWFW98]. Content-based [WWFW98]. contents [RM95]. Continuation [WLW06, Yos10]. Continuity [Dau83a]. Continuous [AADL13, DM96, Dau11]. Contributions [Ano00]. Converge [DL92a, DL01b, DL01a]. Convergence [Dur93, Dur96, RDS04a, RDS04b, RSD07b]. Conversion [DY06, CD00, CDL07, DDGV02, DDGV06]. convex [DTV07]. Convolution [YGLD16, YGLD17, UD97]. convolution-based [UD97]. corrected [DDGV02]. corresponding [Dau80b, FGDB17]. Corrigenda [CD97a]. Corrigendum [DL01b]. Corrigendum/addendum [DL01b]. cosmic [PAHD05]. Coulomb [DL83, DL84, DL05, LT05]. counterintuitive [Dau83a]. coupled [DMC^{+}16c, DMC^{+}16a]. course [Dau93b]. Crack [CRG^{+}13, RCP^{+}11, CYV^{+}13a, CYV^{+}13b]. Cradle [FCY^{+}17, YCF^{+}16, YDC^{+}14]. criterion [CD92]. Crowns [BBL^{+}11]. cubic [Gao14]. curves [DRS04]. cycle [ZCY16]. Cyclic [RDS04b, RDS04a].

D [CD00, CDL07, DDGV02, DDGV06, LK91, LLC08, LQLC10a, LLC11, MMC^{+}13, RB98a, YLB^{+}15]. damping [CVN^{+}13]. dan [P15]. dans [Dur96]. Data [DVDD98, SC00, SC02, ZQW^{+}19, AIK10a, ABD^{+}11, Ara13, DD03, DKRS18, DKRS21, HDH16, SLBD11a, SLBD11b, AIK10b]. Data-Dependent [ZQW^{+}19]. Dataset [GYD^{+}18]. Daubechies [Ano11a, Ara13, Gre11, Grü92, Hei92, Lun92, STAV09, WC02, AAI13, ACV01, Abo94, AIK10a, AM08, Ano93, Ano99, Ano05, Ano08, Ano11e, Ano11b, Ano11c, Ano11d, Ano12, Ano16a, Ano16b, Ano17, Ano18, Anoxx, AC14, AHK13, BB07, BGS14, Bat93a, Bat93b, BPC23, BS94, BAE11, BBN^{+}10a, BBN^{+}10b, BDS98, Boe01, Bon16, Bow03, BSP98, C.93, CS99, ČF04, CGB^{+}15, CWC04, CXS04, CHXL06, CT15, CLG04, Cob01, Coo11, CKE17, CD97b, Daa93, DT14, DMV09, DC11, DAR13, DA15, DJF11, Dor94, DKB99, DW00a, DW00b, DW01, Du01, DL96, Dur93, Dur96, DKLR12, DKLR14, EBJ^{+}14, EGL11, EGL13, Fin04a, Fin04b, GLG94, GL94, yGjZsC11, Gao14, GNG^{+}08a, GNG^{+}08b, GVO^{+}11, GB95b, GB95a, GBM09a, GBM09b, Gre11, Gri95, GEV12, HŚ02b, HŚ02a, HN15, HN17, Hei93]. Daubechies [HPH09, Hop17, HSZ97, Huy08, Jam96, JZL98, Kai10, Kar03, Kar04, Kar07, Kar10, KM12, KLR95a, KLR95b, KLR97, Keh13, KAB11, KTJ09, KKJ^{+}10, Kla97, KK10, KKT10, KLC05, Koz06, Lai95, LS00, LG08, LK91, LP01, LW09, LXDS11, Lin97a, LL00, LKC05, LM93, LM94, LM95, LD96, Lin97b, Lin98, LLC08, LQLC10b, LQLC10a, LLC11, LD16, Lu97, Ma16, MMC^{+}13, MD06, MYN07, MRB^{+}14a, MRB^{+}14b, MMN^{+}11, Mor23, Mor24, NKM12, NG05, NG06, Nie99a, Nie99b, Nie12, NY15, NM13, Nov95, Nov98, Nov02a, Nov02b, Nov02c, OMOE14a, OMOE14b, Old92, P15, PMK16, PKG13, PM11, PM13a, PM13b, PM15, PSB^{+}16, PR05, PM96, PS95, Pol92a, Pol92b, Pol92c, Qix12, RRPT09, RB98b, RB98a, RGMD15a, RGMD15b, Red15, RLS96, RL97, RF09a, RF09b, RM95, RA95, RV09b, RV09a]. Daubechies [SK12, SP17, SHN10, SM12, SS96, SS98, ST15, Sin13, SYSP11, SYSP12, Duk14, Duk16, Sta15, Str92, Sud16, SG13, SZH97, Sun99, Tas99, Tas00, Tay08, Tem96, Tem97, TB94, TDK15, Van08, VP13, VBU05b, VBU05a, VBU07, WLW06, AIK10b, Wah11, WWFW98, Wan01, WP05, Wan07, WP11, Wel99a, Won02, Won11, XTZ10, XZCM06, XWL07, YKIK04, Yan12, Yos10, Yos15, Zei93, Zha00, ZCY16, lZqJmTjZ08]. Daubechies-2 [DAR13, DA15]. Daubechies-based [Jam96, RM95]. Daubechies-Lagarias [STAV09]. Davey [DLM^{+}07]. day [Dau97a]. DCC [SC00, SC02]. De-Noising [MMN^{+}11]. Deblurring [Sin13, DT05, KTJ09]. Decay [DJJ91b, DJJ91a, DH94]. Decimation [DS15a, DS15b]. Decomposition [DLW09, DT04, DT05, DLW11, PS95]. decomposition-like [DLW11]. deconvolution [APD06, APD06]. deep [DDF^{+}22]. deformation [LQLC10a]. Delta [CXS04, DS15a, DS15b, DD03]. Delta-sequence [CXS04]. democracy [CD02]. Denoising [NM13, Sin13, DT05, GEV12]. Density [MRB^{+}14a, RGMD15a, GNG^{+}08a, GNG^{+}08b, MRB^{+}14b, RGMD15b]. Dependent [ZQW^{+}19]. derivative [Dau94b]. derivatives [LKC05]. describe [AD78b]. description [BDV00, BS94]. Design [Nie99b, BDV00, Lin98]. Despeckling [KKJ^{+}10]. Desyat [Dau01]. detailed [DDK05]. Detection [EBJ^{+}14, PJB^{+}09, RCP^{+}11, WWD^{+}15, CYV^{+}13a, CYV^{+}13b, CRG^{+}13]. determinants [Dau95b]. Deterministic [DS15a, DS15b]. Deux [APD06]. Development [GYD^{+}18]. Deviation [DC11]. diagnostics [RRPT09]. diagonalizing [Dau94b]. Dictionary [DMC^{+}17, DMC^{+}16b, DMC^{+}16c, DMC^{+}16a]. Diego [Lun92]. Dietary [BBL^{+}11]. Diffeomorphism [XYD16, XYD18]. Diffeomorphism-Based [XYD16, XYD18]. difference [DL91, DL92b]. Different [Dau93b]. Differential [CMDAF97, DA15]. differentiation [Jam96]. Diffracted [WP11]. Diffusion [SD23]. Digital [ACD^{+}13, CYG^{+}17, DY06, PMK16, PPR^{+}15, SG13, WWD^{+}15, YDC^{+}14, Lai95]. digitally [DDGV02]. dilations [DP88]. Dimensional [GKBD19, ZQH^{+}17, CCD16a, CCD16b, LP01, PM96]. Dimensions [ADGY16, CDD^{+}12, ADGY19]. Diminishing [Dor94]. d’Ingrid [Dur96]. Direct [LKC05]. Directional [YD16, YD17, YD18]. Dirichlet [BBL^{+}11]. Discharge [YKIK04, XTZ10]. Discrete [Ano11e, Ano16a, Dau87, Dau90a, RV09b, Dau84, Dau89a, ST15]. disjoint [Tas99]. Distance [LCDF10, LPD11, AADL13, Dau11, LPD13]. distances [LD11a, LD11b]. Distinction [YKIK04]. distribution [Boe01, Kar10]. distributions [Dau80b]. do [Dau96]. does [DH95, DRT^{+}09]. Domain [CWC04, ABMD90, KLC05]. drawings [YMH^{+}16]. dual [Dau95a, DH02, DH04]. Dutch [Ano05, Gre11, Huy08]. Dyadic [YD16, YD17, RF09a, RF09b, YD18]. dynamic [CHXL06]. Dynamics [RDS04c, WHB^{+}14, PS95, RDS04a, RDS04b].

ECG [AC14, BGS14, NM13]. ecomorphology [FGDB17]. Editor [DLM^{+}07]. editors [Lun92]. EEG [Ano08]. efficacy [SYSP11, SYSP12]. Efficiency [AHK13]. efficient [FVD21, NG05, NG06]. eigenelements [PSB^{+}16]. eigenfunctions [DW01, Du01]. Eigenvalues [DW01, Du01, Yos10]. Ekstraksi [P15]. elastic [LLC08, LQLC10b, LLC11, PKG13]. elasticity [LD16]. elastostatics [DL96]. electrical [BSP98]. Electrocardiogram [WWD^{+}15]. Electrocardiographic [OWW^{+}16]. Electromagnetic [yGjZsC11, LXDS11]. Electron [DL83, DL05, LT05, Dau84]. electronic [GVO^{+}11]. element [PKG13]. elements [DMV09, PM96, WMJ^{+}11]. elétricos [VP13]. Embedding [LCDF10]. Empirical [DLW09, DLW11]. Encoder [DGWY08, DGWY10]. encoding [BD03, CDDD01]. Energy [BBL^{+}11, Dau83b, Dau84, NG05, NG06]. Enestrom [Kar04, Kar03]. enforcing [DDD16]. engineering [BDKP14, Dau93e]. enhancement [NY15]. entire [DG88]. entretien [BPC23, Mor23, Mor24]. equation [BAE11, PM13b, PR05]. equations [DL91, DL92b, DT14, GL94, MD06]. Erratum [DJJ91a]. ESPRIT [XWL07]. estimate [CD96, CD97a, Dau95b, Sun99]. Estimates [RF09a, RF09b, WLD^{+}16]. Estimation [DC11]. Étude [Dur96]. Evaluating [WHB^{+}14]. Event [NM13]. Exact [Ma16]. examples [Dau83a]. Existence [DL91]. Expansions [DGM86, DJ93, DDGV02, DGM06]. Experimental [ZGSD04, ZGSD06]. experiments [PAHD04, PAHD05]. explicit [AM08]. exponent [Sun99]. Exponential [DJJ91b, DY06, CD00, CDL07, DJJ91a, DKLR12, DKLR14]. expressed [NG05, NG06]. expression [KLT^{+}10, SHN10]. Extensions [LPD11, LPD13]. Extracão [VP13]. Extraction [NM13, P15, BGS14, VP13]. Extrapolation [CWC04]. extremal [DJF11].

Face [BBN^{+}10a, BBN^{+}10b, Hop17]. Facial [P15, SHN10]. Factored [LCDF10]. Factoring [DS98, DS00]. factorization [Tas00]. failures [VP13]. falhas [VP13]. families [Tas00, VBU07]. family [DD03, PM96, SHN10]. fast [CDJV93, CDV93, NKM12, AIK10b, AIK10a]. faster [DDFG08c]. fault [RRPT09]. Feature [NM13, P15, Sud16, BGS14]. features [FGDB17, PCR^{+}11]. Feauveau [Ara13]. Fellows [Ano93]. Female [Duk14]. fermions [Dau83b]. fewest [FVD21]. field [BS94, LXDS11, OMOE14a, OMOE14b]. Fields [Duk14]. filter [EGL11, EGL13, Lai95, SK12, VP13]. Filters [CKE17, Dau90a, Lin98, RB98a, Dau89a, DJF11, HSZ97, KLR95a, KLR97, Kla97, MMC^{+}13, RB98b, SS98, SZH97, Tas00]. filtres [KLR95b]. filtros [VP13]. Finite [Dau90a, CD97b, Dau89a, DMV09, PKG13, PM96, Zha00]. First [Duk14, EBJ^{+}14]. Fitur [P15]. Flanders [Hop17]. Flemish [Sta15]. fluctuations [OMOE14a, OMOE14b]. fluorescence [ABD^{+}11, ABD^{+}13]. fMRI [DRT^{+}09]. Foreword [ACD^{+}13, Dau06a]. forgery [PJB^{+}09]. formalism [DL94a, DL94b]. formula [DM99, PM11]. formulation [PKG13]. Foundation [Keh13, Duk16]. Four [SK12]. Fourier [AIK10a, ZGSD06, BDV00, Red15, AIK10b, ZGSD04]. Fractal [Wel99b]. fractals [DL92b]. Fragment [RGMD15a, RGMD15b]. frame [BDV00, DH02]. Framelets [DHRS03, YGLD16, YGLD17]. framer [Dau97a]. Frames [ADGT17, DG88, ADGT16, Dau97a, DHRS03, DH04]. Framework [ZQW^{+}19, WP11]. Fredholm [Dau95b]. Free [CKE17, BBJ^{+}09, Ano05]. French [APD06, BPC23, Dur96, Mor24]. frequencies [WFD11]. Frequency [Dau97b, DWtW15, CHT98, CD93d, Dau88b, DP88, Dau89b, Dau90b, Dau95a, DLL95, Dau97a, Dau98e, Dau06c, DWtW16, NKM12, YLB^{+}15]. Fully [BPG^{+}15, GYD^{+}18]. Function [Ano08, Pol92b, Pol92c, CDR96, CDP97, DH95, DD03, Lin97a, PM11, PM13a, PM13b, Pol92a, RSD07a, RSD08]. Functional [Grü92, MRB^{+}14a, RGMD15a, GNG^{+}08a, GNG^{+}08b, MRB^{+}14b, RGMD15b, RRD12]. functionals [DT04]. Functions [ACV01, ADGT17, Ano11b, CDD^{+}12, Dau80c, DS15a, DS15b, yGjZsC11, HN15, RV09a, XYD16, XYD18, ADGT16, BDS98, CD96, CD97a, CDL02, DG88, DH94, DL94a, DL94b, Dau95a, Dau95b, DH04, DDD^{+}23, DW00a, HN17, LS00, LKC05, MYN07, PM15, SS98, Sun99, Wan01]. Fundamental [HW06]. Fusion [BBN^{+}10a, BBN^{+}10b, CGB^{+}15]. future [PAHD04, PAHD05]. fuzzy [ZCY16].

Gabor [WC02, BB07, CHT98, CWC04, Cob01, Dau95a, DLL95, DP02]. Gabor-Daubechies [WC02]. galaxies [APD06]. Gauss [PM15]. Gauss-type [PM15]. Gaussian [DW00a, DW00b, GKBD18, GKD19, GKBD19, WP11]. geboortedorp [Gre11]. General [OWW^{+}16, DTV07]. Generalized [VBU05b, VBU05a, VBU07, Dau83b, DKLR12, DKLR14]. genomes [KLT^{+}10]. Geometric [GYD^{+}18, GKBD19, BLS^{+}11a, BLS^{+}11b, Dau88b, DP88, PBGD13]. geophysical [SLBD11a, SLBD11b]. Geostatistics [Pil09]. gets [Gre11]. Ghent [CRG^{+}13, PPR^{+}15, RCP^{+}11]. GIS [Pil09]. given [CDR96]. Global [CVN^{+}13, DL91, SLN^{+}11a, SLN^{+}11b]. Globally [GYD^{+}18]. Gogh [JHB^{+}08]. Golden [DGWY08, DGWY10]. Gordon [Bat93b]. Gradient [DFL07, DFL08]. Grant [Duk16]. graph [DDK05]. Gregory [Lun92]. grooved [lZqJmTjZ08]. growth [Nov95]. gv [RSD07b].

Haar [AAI13]. Hamiltonians

[DK85, KD84]. Hammerstein [MD06]. Harcourt [Lun92]. hardcover [Lun92]. Harmonic [CDDD03, DG99, GBM09a, GBM09b, DKRS12, DKRS15, DKRS18, DKRS21, DVDD98, Gao14]. Heart [WLD^{+}16, EBJ^{+}14]. Heisenberg [BD03, GB95b, GB95a]. held [DKRS12, DKRS15]. Henriksen [DLM^{+}07]. Henryk [Ano11a]. Herpesviruses [KLT^{+}10]. heterogeneity [SLN^{+}11a, SLN^{+}11b]. hidden [ABD^{+}13]. hiding [Ara13]. High [CDD^{+}12, GYD^{+}18, YKIK04, CYV^{+}13a, CYV^{+}13b, GVO^{+}11, KAB11]. Higher [SP17]. highly [BGS14]. Hilbert [Tay08, AD79c, CCD16a, CCD16b, Dau80c]. Hilbert-Pairs [Tay08]. holomorphic [Dau78a, Dau78b]. Homogenization [DRZ07]. Honored [Ano00]. Human [BBN^{+}10a, BBN^{+}10b]. hybrid [DKRS21]. hyperdifferential [Dau78a, Dau78b]. hypersingular [PM13a].

I. [Dur93]. Iberoamericana [CD97a]. Identification [HŚ02b, HŚ02a, HPH09, DM99, FVD21, JHB^{+}08]. Identity [Zei93, DLL95]. II [DGR83, DK85, DP88, DL92b, Dau93c, KKT10, LPD11, LPD13]. III [CD93d]. Image [ACD^{+}13, ABMD90, ABMD92, BBN^{+}10b, DMC^{+}17, HD06, JHB^{+}08, KK10, KKT10, Lin97b, PMK16, PPR^{+}15, PZC^{+}12, PZC^{+}15, Sin13, Wel99b, YGLD16, YGLD17, CDSY97, CGB^{+}15, CLG04, DT04, DT05, DTV08, DMC^{+}16b, DMC^{+}16c, DMC^{+}16a, GEV12, MDSW92, NY15, RB98b, SM12, ST15, Sud16, WWFW98]. imagery [Ara13]. Images [BB07, BBN^{+}10b, MMN^{+}11, YCF^{+}16, ABD^{+}13, CYV^{+}13a, CYV^{+}13b, FCY^{+}17, KTJ09, KKJ^{+}10, TDK15, YDC^{+}14]. Imaging [Wah11, WC02, ABD^{+}11]. imperfect [DDGV06]. Implementasi [SG13]. Implementation [CHT98, CKE17, Wah11, LL00]. Implementing [CD97b]. implications [FGDB17]. Importance [LD96, CDGO02]. improved [DAR13]. IMU [CDW14, Dau15]. incoming [KLT^{+}10]. independence [DRT^{+}09]. Independent [DRT^{+}09, RRD05]. indexing [WWFW98]. Indonesian [P15]. Inference [BBL^{+}11]. Infinite [ADGY16, DL92a, ADGY19, CCD16a, CCD16b, DL92b, DL01b, DL01a]. infinite-dimensional [CCD16a, CCD16b]. Information [HD06, TSS04]. Ingrid [Ano05, Ano11a, Bat93a, C.93, Gre11, Gri95, Grü92, Hei93, Lun92, Ano05, Ano16b, Ano17, Ano18, Anoxx, BPC23, Bon16, Coo11, Dur96, Gre11, Hei92, Hop17, Huy08, Mor23, Mor24, Duk16, Sta15]. Inner [RL97]. Inpainting [RCP^{+}11, CRG^{+}13]. Instability [CD93c]. Instantaneous [WLD^{+}16]. insurance [AIK10a, AIK10b]. integer [CDSY97]. integers [CDSY98, MMC^{+}13]. Integral [DG80, GBM09a, GBM09b, HSZ97, Dau78c, DGR83, LL00, PM13b, SZH97]. Integrals [DK82, DK83, KD82a, DK85, DK86, DKP87, Dau91, KD82b, KD84, MYN07, PM13a, PM15]. Interaction [DL83, DL05, LT05, DL84]. Interactions [JDB^{+}14]. Interfacing [Pil09]. internationaal [Gre11]. International [JDB^{+}14]. internationally [Gre11]. interplanetary [OMOE14a, OMOE14b]. Interpolation [CDL02, UD97]. interpretation [AD79d]. Interspecific [GYD^{+}18]. interval [CDJV93, CDV93, Dau93f, Dau94b, Jam96, LL00, Zha00]. intervals [Fin04a, Fin04b]. Interview [Ano16b, Bon16, Huy08, BPC23, Mor24, Huy08]. Intrepid [Bon16]. Intrinsic [ZQW^{+}19]. introduced [Dur93]. Introduction [DMW92, Grü92]. invariant [TB94]. Inverse [DFL07, DFL08, DDD03, DDD04, DTV07]. inversion [LNDD06, LNDD07, SLBD11a, SLBD11b]. Investigation [DMC^{+}17, DMC^{+}16b, YLB^{+}15]. involving [MYN07, PM15]. IR [HPH09]. irregular [CDL02, DGS99, DGSS99, DGS01]. ISBN [Lun92]. Ischemic [NM13]. Ising [PS95]. Ising-model [PS95]. Isolation [BBL^{+}11]. isometry [Cob01]. Issue [ACD^{+}13, DMW92, KD96]. ISTA [DDD16]. iterative [DDD03, DDD04, DTV08]. Iteratively [DTV07, DDFG08a, DDFG08b, DDFG08c, DDFG10, VD15, VD17]. iTWIST’14 [JDB^{+}14]. Iwaniec [Ano11a].

J [OWW^{+}16]. January [Dau93b]. John [Ano11a]. joint [AD78b]. Jones [Lun92]. Jovanovich [Lun92]. JPEG2000 [KAB11]. June [DKRS12]. justification [AD78b].

Kakeya [Kar03, Kar04]. kernel [PM13b, XLY^{+}23]. kernel-based [XLY^{+}23]. Kernels [Dau80c]. kind [PM13b]. kinematic [DKP87]. kinetic [Dau83b, Dau84]. Klauder [GBM09a, GBM09b]. Klein [Bat93b]. known [Gre11]. Komparasi [P15]. Kompresi [SG13]. komt [Ano05]. krijgt [Gre11]. Kuramoto [DT14].

l [LNDD07]. Lagarias [STAV09, BAE11]. landmark [BBJ^{+}09]. landmark-free [BBJ^{+}09]. Landmarking [GKBD18, GKD19, GKBD19]. large [LQLC10a]. lattice [DJ93]. lattices [Dau95a, DLL95]. LDMNet [ZQH^{+}17]. lead [WWD^{+}15]. Learning [DMC^{+}17, RRD05, ZQW^{+}19, DMC^{+}16b, DMC^{+}16c, DMC^{+}16a, JPB^{+}09, PJB^{+}09, RRD12]. Least [DDFG08a, DDFG08c, DDFG10, Tas99, VD15, DDFG08b, UD97, VD17]. Lectures [Ano15, Bat93a, C.93, Dau92, Dau01, Gri95, Grü92, Hei92, Hei93]. LeGall [Ara13]. lektsij [Dau01]. Letter [DLM^{+}07]. lifting [CHXL06, DS98, DS00]. like [DLW11, SLBD11a, SLBD11b]. limited [KLT^{+}10]. Linear [DL01b, DFL07, DFL08, MRB^{+}14a, AM08, Dau80a, DDD03, DDD04, DTV07, DDFG08c, HŚ02a, MRB^{+}14b]. Liouville [PSB^{+}16]. Local [CWC04, YGLD16, YGLD17, DL92b]. Local-Nonlocal [YGLD16, YGLD17]. localisation [DP88]. localization [Cob01, Dau88b, Dau90b, Dau95e, Dau98e, Dau06c, Nov95, Yos10, Yos15]. locally [CDDL21]. Lorentz [DKB99]. Lossless [CDSY97]. Louise [Lun92]. Low [Wah11, ZQH^{+}17, Kla97].

MA [Lun92]. MacArthur [Ano93]. machine [JPB^{+}09, RRPT09]. magnetic [OMOE14a, OMOE14b]. making [Dau93e]. Mallat [Lun92, LG08]. mammography [AHK13]. Manifold [ZQH^{+}17]. Manifolds [GKBD18, GKD19]. map [CDSY98]. maps [AD79a, AD78a, AD83, SD23]. March [SC00]. margin [RSD04, RSD07a, RSD08]. margins [RDS04a, RDS04b]. Markovic [DLM^{+}07]. Markowitz [BDD^{+}07, BDD^{+}09]. Mary [Lun92, ADG^{+}24]. mask [DH95]. Mass [LD09]. Masterpiece [Dau16]. Mat. [CD97a]. Matching [LG08]. Math [Duk14]. Mathematica [RA95]. Mathematical [Dau93b, AD79d, CK96, DGR83, Dau16]. Mathematician [Bon16, Gre11, Huy08, Ano05]. Mathematics [CL05, Dau16, GBGL08, BDKP14, Dau93e, Dau95d]. Maths [Ano14, Sta15]. Matrices [DL92a, DL92b, DL01b, DL01a, EGL11, EGL13]. Matrix [CWC04, Jam96]. Matzinger [DKB99]. Mayor [Gre11]. means [DT05, ZCY16]. measure [DK85, Dau91]. measurements [CDDL21, XTZ10]. Measures [DK82, DK83, KD82a, DK86, DKP87, KD82b, KD84]. Mechanical [Dau80c, AD79a, AD78a, AD83, DK85, KD82b, KD84]. Medal [Duk14]. Medical [Ano00, CGB^{+}15, Wah11, KTJ09, KKJ^{+}10]. meeting [DKRS21]. Memorizing [ZQW^{+}19]. memory [ADG^{+}24].

Menggunakan [P15]. meshless [CT15, LXDS11, LLC08]. Method [Ara13, DFL07, DRZ07, DFL08, yGjZsC11, WLW06, CHXL06, CT15, Gao14, LXDS11, LLC08, LQLC10b, LLC11, LD16, MD06, WP05]. méthodes [APD06]. Methods [BBL^{+}11, GYD^{+}18, AAI13, APD06, Dau95e, MYN07, PKG13]. Metrics [HD06]. Meyer [Dau93a, Lun92, Dau10]. microwave [PAHD05]. Milnor [Ano11a]. minimal [GB95b, GB95a]. Minimization [DDFG08a, DDFG10, DDFG08b, DDFG08c]. mixed [Abo94]. Modal [DMC^{+}17, DMC^{+}16b]. Mode [DLW09, DLW11]. Model [GLG94, NY15, PS95]. modeling [AAI13]. modelling [CLG04, PR05]. Models [WPS^{+}14, DM96, HŚ02b, HŚ02a, JDB^{+}14, SLN^{+}11a, SLN^{+}11b, WPS^{+}13]. Modified [Nov95, Nov98, Nov02a, Nov02b, Nov02c]. modulators [DD03]. Molar [BBL^{+}11]. Molecules [DL83, DL05, LT05, Dau84, DL84]. moments [Sud16]. Monte [PS95]. Mori [PS95]. morphological [FGDB17]. Morphometric [GYD^{+}18]. Morphometrics [GKBD19, PBGD13]. most [Tas99]. mother [Anoxx, AC14]. motor [VP13]. motores [VP13]. MR [BB07]. MR1142737 [DL01b]. MR1402677 [CD97a]. MRA [DHRS03]. MRA-based [DHRS03]. MRT [TDK15]. MRTD [yGjZsC11]. Multi [DMC^{+}16b, DMC^{+}17, LLC11]. Multi-Modal [DMC^{+}17, DMC^{+}16b]. Multi-scale [LLC11]. multifractal [DL94a, DL94b]. multilevel [KTJ09]. Multimodal [HD06, CYV^{+}13a, CYV^{+}13b]. multiple [BDV00]. Multiplier [CKE17, Qix12, Yan12]. multipliers [LK91]. Multiresolution [BBN^{+}10b, CDJV93, AM08, DMW92, DRS04, KLC05]. Multiscale [DRZ07, SP17, CHXL06]. Multisegmentation [SYSP11, SYSP12]. multitapered [DWtW15, DWtW16]. Mumford [Keh13]. Musculoskeletal [GEV12].

naar [Ano05]. near [CGB^{+}15]. needs [CDW14]. Nemmers [Ano12]. nerve [DM96]. network [DDD^{+}23, YMH^{+}16]. Networks [ZQH^{+}17, DDF^{+}22, FVD21]. Neumann [AD79c]. Neural [DDD^{+}23, ZQH^{+}17, FVD21, TSS04, YMH^{+}16]. neuroimaging [RRD12]. NIPS [TSS04]. NLS [GL94]. no. [CD97a]. nodes [PM15]. Noise [DC11, KTJ09, XTZ10]. Noising [MMN^{+}11]. Non [CD93a, CD93b, HŚ02a, XLY^{+}23]. non-linear [HŚ02a]. Non-separable [CD93a, CD93b]. non-stationary [XLY^{+}23]. Nonlinear [DDF^{+}22, GLG94, LDN^{+}08, LDN^{+}10, CDGO02, DM96, HŚ02b]. Nonlocal [YGLD16, YGLD17]. Nonorthogonal [DGM86, DGM06]. Nonparametric [PZC^{+}12, PZC^{+}15]. norm [LNDD06, LNDD07]. Normal [BBL^{+}11, DRS04]. Note [HN15, PM11]. number [Wan01]. numbers [KLT^{+}10]. Numerical [GLG94, GL94, LXDS11, LD16, NG05, NG06]. numerics [Tem96, Tem97].

Object [YMH^{+}16, TB94]. ocean [Zha00]. old [Dau16]. ondelettes [Dur96]. One [DL83, Dau84, DL05, LT05, PM96, WFD11, AD78b, PM11, PR05]. One-dimensional [PM96]. One-Electron [DL05, LT05, Dau84]. one-periodic [PR05]. one-point [PM11]. Opening [Dau15]. operator [Dau94b, Yos10]. Operators [Dau80c, SP17, Won02, Cob01, Dau78a, Dau78b, Dau80b, Dau88b, DP88, DW00a, DW00b, DW01, Du01, DL96, Yos15, Dau93a]. Optimal [BD03, CDDD01, Str92]. optimization [XLY^{+}23]. optimized [Tas00]. Optimizing [WLD^{+}16]. Order [KLC05, SP17, DD03, Kla97]. orientation [BBJ^{+}09]. oriented [HŚ02b, HŚ02a]. original [Dau97a]. Orthogonal [Dau89a, HN15, HN17, Kar10, RF09a, RF09b]. Orthonormal [CD93d, Dau88a, Dau90a, DJJ91b, Dau93c, Dau06b, Kai10, DJJ91a, Dau93d, DJF11]. Oscillator [GBM09a, GBM09b]. Other [ADGT17, ADGT16, Dau95e, DJF11]. Outcomes [OWW^{+}16]. oversampled [CD00, CDL07]. overview [Dau95c].

Packets [CD93c]. pada [P15, SG13]. Painless [DGM86, DGM06]. Painting [WPS^{+}13, WPS^{+}14, JHB^{+}08, PPR^{+}15]. Paintings [CYG^{+}17, YCF^{+}16, ABD^{+}13, CYV^{+}13a, CYV^{+}13b, CRG^{+}13, FCY^{+}17, JPB^{+}09, PCR^{+}11, PJB^{+}09, YDC^{+}14]. Pairs [DH04, Tay08]. panel [FCY^{+}17]. Papers [HW06]. parâmetros [VP13]. para [VP13]. parabolic [PR05]. parallel [WP11]. parameter [SD23]. Parameterizations [LM94]. parameters [AHK13, DKLR14, VP13]. Parametrizations [LM93]. part [BPC23, Mor24, XTZ10]. partial [PM11, XTZ10]. partie [BPC23, Mor23, Mor24]. Patches [YGLD16, YGLD17, LD16]. Path [DK82, DK83, GBM09a, GBM09b, KD82a, DK85, DK86, DKP87, Dau91, KD82b, KD84]. path-integrals [KD82b]. Pattern [ZQW^{+}19]. Patterns [CYG^{+}17]. Pauli [Ano15]. PCA [P15, P15]. PDE [CXS04]. pearls [PCR^{+}11]. Pengenalan [P15]. perfect [EGL11, EGL13]. Performance [Sin13, AC14, GVO^{+}11, SHN10]. periodic [PR05]. Periodized [RLS96, RL97]. personal [Dau96]. Perspectives [Dau93b]. pg [DL01a]. Phase [ADGY16, CCD16a, CCD16b, Dau91, GBM09a, GBM09b, LD96, ADGY19, CDDL21, Dau88b, DP88, Dau95e, KLR95b, KLR97]. Physical [AD78b, AD79d]. Physics [Ano00, AD79b]. Physiological [WHB^{+}14]. PiPs [XLY^{+}23]. plane [DL96]. plate [DMV09, Koz06]. platforms [WP11]. Point [CDD^{+}12, Dau96, DGSS99, PM11, XZCM06]. points [JZL98]. Polyharmonic [KK10, KKT10]. polynomial [DK85, KD84, LD11a, LD11b]. polynomials [Boe01, DKB99, DKLR12, DKLR14, Kar07, Kar10, KM12, Kla97, Nov02a, Nov02b, Nov02c, SS96, Tem96, Tem97]. Population [OWW^{+}16]. portfolios [BDD^{+}07, BDD^{+}09]. Portuguese [VP13]. potential [NG05, NG06]. power [BSP98, UD97]. pp [Lun92]. Practical [CHT98, DY06]. Pratt [DLM^{+}07]. Precise [RSD07b]. Precision [DY06]. Prediction [WHB^{+}14]. Preface [Dau98a, Dau98b]. presence [KTJ09]. present [Dau97a]. present-day [Dau97a]. preserving [AD79a, AD78a, AD83, DAR13, DA15, Nov95]. President [Dau15]. Press [Lun92]. Previous [BBL^{+}11]. Princeton [GBGL08]. Principal [P15, P15]. principle [Dau83b]. privacy [DAR13, DA15]. Prize [Ano12]. Prizes [Ano11a]. Probabilistic [WPS^{+}14, WPS^{+}13]. Problem [Bow03, XZCM06, Zha00]. Problems [DFL07, DRZ07, DFL08, DDD03, DDD04, Dau05, DTV07, LLC08, LQLC10b, LLC11, LD16, PSB^{+}16]. Proceedings [JDB^{+}14, SC00, SC02, TSS04]. Process [GKBD18, GKD19, GKBD19]. Processing [ACD^{+}13, KK10, KKT10, PPR^{+}15, DKRS18, JHB^{+}08, Lin97b, TSS04]. Procrustes [AADL13, Dau11]. product [AD78b, LP01, RL97]. Products [DL92a, Du01, Won02, DL92b, DL01b, DL01a]. professor [Sta15]. project [GVO^{+}11]. Projected [DFL07, DFL08]. projective [Dau80a]. Proof [DDFG08c, AD83]. Propagation [GLG94, WC02, PKG13, PR05, Bat93b]. Properties [CLG04, DGR83, Dau84, DJF11, Kla97]. property [SD23]. propositional [AD79a, AD78a, AD79c, AD79d, AD83]. pros [CD02]. pseudopotential [GNG^{+}08a, GNG^{+}08b]. Pulse [WLD^{+}16]. Pursuit [LG08].

Quadratic [DK83]. quadrature [NG05, NG06, PM11, PM13a, PM15]. quality [BSP98, KAB11]. Quantification [BBL^{+}11, BBJ^{+}09]. quantify [BLS^{+}11a, BLS^{+}11b]. Quantitative [YLB^{+}15]. Quantization [DG80, DS15a, DS15b, ABMD90, Dau78a, Dau78b, Dau78c, Dau80b, Dau83a, DGR83, MDSW92]. quantized [CDL02, DD03]. quantizers [DDGV06]. Quantum [Dau80c, DK85, KD84, AD79a, AD78a, AD78b, AD83, KD82b]. Quantum-mechanical [DK85, KD82b]. Queries [CDD^{+}12]. Quincunx [YD16, YD17, YD18]. Quotient [BBN^{+}10b].

radial [DW01]. random [Ara13]. Randomized [ZGSD04, ZGSD06]. Ranks [Ano98]. Raphael [Lun92]. Rate [WLD^{+}16, CD00, CDL07, DDFG08c]. Ratio [DGWY08, DGWY10]. rational [EGL11, EGL13]. Ray [YCF^{+}16, ABD^{+}11, ABD^{+}13, DMC^{+}16c, DMC^{+}16a, FCY^{+}17, YDC^{+}14]. Raz [DLL95]. Re [DDFG08a, DDFG08c, DDFG08b]. Re-weighted [DDFG08a, DDFG08c, DDFG08b]. Reading [Bon16]. Real [ADGT17, ADGT16, DK86, DKLR14, LG08]. real-time [LG08]. Real-Valued [ADGT17, ADGT16, DKLR14]. reassignment [CMDAF97]. Receive [Keh13]. Recognition [BBN^{+}10a, BBN^{+}10b, P15, SHN10, TB94, YMH^{+}16, ZCY16]. Reconstructing [ADGT16, ADGT17, PAHD04, PAHD05]. reconstruction [ABD^{+}11, APD06, EGL11, EGL13]. Recovery [DDFG08a, DDFG10, DDFG08b, DDFG08c]. Recursive [XYD16, XYD18]. redundant [CS99]. refinable [CDR96, CD96, CD97a, CDP97, DH94, DH95, Dau95b, DH04, DDD^{+}23]. reflex [Huy08]. Regression [XYD16, XYD18]. regularisation [DDD16]. Regularity [CDP97, DGS99, DKLR12, DKLR14, CD96, CD97a, DL91, DL92b, LS00, Sun99]. Regularization [SP17, VD15, YGLD16, YGLD17, ZQW^{+}19, CVN^{+}13, LNDD06, LNDD07, LDN^{+}08, LDN^{+}10, VD17]. Regularized [ZQH^{+}17]. Related [DG80, Boe01, CD92, Dau78c, DGR83, Nov02c, Tem96, Tem97]. Relationship [SP17]. Relativistic [DL83, DL84, DL05, LT05, Dau84]. ReLU [DDF^{+}22]. Removal [CYG^{+}17, FCY^{+}17, YDC^{+}14]. Removing [YCF^{+}16]. Repair [Dau16]. replication [KLT^{+}10]. Reply [DLM^{+}07]. represent [AD79d]. Representation [Dau80c, DL96, LD96, Dau80a, HSZ97, Lu97, SZH97, TB94]. representations [RRD12]. reproducing [DKLR12, DKLR14]. Research [yGjZsC11]. Resolution [PZC^{+}15, CD93d, CYV^{+}13a, CYV^{+}13b, PZC^{+}12]. resolving [SLN^{+}11a, SLN^{+}11b]. resource [FVD21]. Respect [ADGT17, ADGT16, Red15]. Restoration [ABD^{+}13, RCP^{+}11, CRG^{+}13, DT05, DTV08]. results [Dau94b, Dau98c, Dau98d]. Retrieval [ADGY16, ADGY19, CCD16a, CCD16b, CDDL21, Sud16]. Rev. [CD97a]. reveal [KLT^{+}10]. Review [Bat93a, C.93, Gri95, Grü92, Hei93]. Reviews [Dau93a, Hei92, Lun92]. revisited [DDD16]. revitalize [Dau16]. Reweighted [DDFG10, VD15, VD17]. Ridge [CDD^{+}12]. Risk [OWW^{+}16]. RNS [SK12]. Robust [DY06, FVD21]. Rock [HPH09]. Ronald [Lun92]. root [Tas99]. roots [Kar07, KM12, Nov02a, Nov02b]. rule [PM15]. rules [AM08, PM13a]. Ruskai [Lun92, ADG^{+}24]. Russian [Dau01].

S [CHT98]. SADT [SM12]. samples [CDL02, FVD21]. sampling [JZL98]. San [Dau93b, Lun92]. scale [BAE11, DL91, DL92b, Dau97a, LLC11, PM11, PM13a, PM13b, PM15, SLN^{+}11a, SLN^{+}11b]. Scaling [ACV01, Ano11b, yGjZsC11, HN15, MRB^{+}14a, Pol92b, Pol92c, RV09a, BDS98, ČF04, HN17, LS00, MRB^{+}14b, Pol92a, SS98, Sun99, Wan01]. scanning [Ara13]. Scattering [CWC04, yGjZsC11, Zha00]. scheme [CD92, XLY^{+}23]. schemes [BDV00, BD03]. Schneider [Ano93]. science [BDKP14, DKRS21]. Sciences [Ano98]. search [Tas00]. search-optimized [Tas00]. searching [WWFW98]. Second [JDB^{+}14, EBJ^{+}14, PM13b]. secret [Ara13]. sector [AAI13]. segmentation [FGDB17]. seismic [CVN^{+}13, LDN^{+}08, LDN^{+}10, SLN^{+}11a, SLN^{+}11b]. select [DRT^{+}09]. selected [Tas00]. Self [SM12]. selfreciprocal [Kla97]. semigroup [SD23]. Sensitive [YKIK04]. separable [CD93a, CD93b]. Separation [DMC^{+}17, APD06, DMC^{+}16b, DMC^{+}16c, DMC^{+}16a, APD06]. sequence [CXS04, XZCM06]. Series [Lu97, AIK10a, DW00b, Lun92, AIK10b]. set [CGB^{+}15, GNG^{+}08a, GNG^{+}08b]. Sets [DL92a, DL01b, DL01a, BD03, Dau87, DGSS99, Tas99]. shallow [FVD21, Zha00]. Shannon [JZL98]. Shape [BBL^{+}11, XYD16, XYD18, BBJ^{+}09]. Shapes [BPG^{+}15]. ship [CT15]. short [Dau93b]. shows [Dau16]. shrinkage [KTJ09]. Sigma [DS15a, DS15b, DD03]. Signal [GLG94, NM13, Dau87, Dau90b, DMW92, Dau98e, Dau06c]. Signals [LD96, WLD^{+}16, XLY^{+}23]. Similarity [HD06, BLS^{+}11a, BLS^{+}11b]. Simons [Duk16]. Simple [AD83, DJJ91b, DJJ91a]. Simpler [HD06]. Simplified [STAV09]. Simulation [GLG94, PKG13]. simultanées [APD06]. Simultaneous [DT05, APD06]. Single [CD00, CDL07]. Single-bit [CD00, CDL07]. singular [LL00, PM13a]. singularities [PM13a]. Sivashinsky [DT14]. Sleep [Ano08]. smooth [CDR96, RSD04, RSD07a, RSD08]. smoothest [CDR96]. smoothing [CVN^{+}13]. smoothness [Dau95b, Nov95, RF09a, RF09b]. Snowbird [SC00, SC02]. Sobolev [Sun99]. Society [Dau93b]. Solution [PM13b, GL94, Zha00]. solutions [DL91]. Solving [SLN^{+}11a, SLN^{+}11b, CXS04, DTV07, DT14]. Some [DJF11, DTV08, DGR83]. sounds [EBJ^{+}14]. source [BDV00]. Space [GBM09a, GBM09b, CDR96, CDDD03, DG88, Dau88b, DP88, Dau91, Dau95e]. Spaces [Dau80c, Qix12, AD79c, CCD16a, CCD16b, Yan12]. Sparse [BDD^{+}07, BDD^{+}09, DRZ07, DDFG08a, DDFG10, JDB^{+}14, VD15, ZGSD04, DDFG08b, DDFG08c, DKRS12, DKRS15, RRD12, VD17, ZGSD06]. Sparsity [DFL07, DFL08, DDD16, CVN^{+}13, DDD03, DDD04, SLN^{+}11a, SLN^{+}11b]. Sparsity-enforcing [DDD16]. Spatiogram [PCR^{+}11]. speaker [DM99]. Special [ACD^{+}13, DMW92, KD96]. Spectra [HPH09]. Spectral [DRZ07, PKG13, Yos15, Tas00]. spectrum [Dau84]. speed [Dur96]. sperm [LG08]. sphere [SLBD11a, SLBD11b]. spherical [SLN^{+}11a, SLN^{+}11b]. Spline [HN15, Gao14, LD16]. Squares [DDFG08a, DDFG08c, DDFG10, VD15, DDFG08b, UD97, VD17]. squeezing [DM96]. stability [AM08, CD92]. Stable [ADGY16, ADGY19, CDDL21, BDD^{+}07, BDD^{+}09, DD03]. Stage [Ano08]. Standard [DC11]. statements [Dau83a, RSD07b]. states [Dau80a, Dau87, Dau91, Dau94a]. stationary [XLY^{+}23]. statistical [AHK13, BS94, CLG04]. Steele [Ano11a]. steganography [ST15]. STEM [BDKP14].

Stephane [Lun92]. steps [DS98, DS00]. stochastic [BD03]. Stop [ZQW^{+}19]. straat [Gre11]. strategies [CDGO02, DDK05]. street [Gre11]. structure [AD79a, AD78a, AD83, FGDB17, GVO^{+}11]. structure-preserving [AD79a, AD78a]. structures [CT15]. Studi [P15]. Study [CT15, P15, DDK05, OMOE14a, OMOE14b, PPR^{+}15, Dur96, OWW^{+}16]. Sturm [PSB^{+}16]. style [WMJ^{+}11]. Stylistic [JPB^{+}09]. subband [CD92]. Subbanding [XWL07]. subdivision [DGS99, DGS01]. sublattice [AD79d]. Subsampling [YD16, YD17, YD18]. subsystem [AD79d]. subsystems [AD79b]. Sunyaev [PAHD04, PAHD05]. Super [PZC^{+}12, PZC^{+}15]. Super-Resolution [PZC^{+}15, PZC^{+}12]. supervised [PJB^{+}09]. Support [Dau90a, Dau89a, PM11]. Supported [Dau88a, CDF92, CD93d, Dau93c, Dau06b, GB95b, GB95a, PKG13, WP05]. Supporting [PPR^{+}15]. Suppressing [XTZ10]. Surface [BBL^{+}11, LD09]. surfaces [AADL13, BLS^{+}11a, BLS^{+}11b, Dau11, LD11a, LD11b]. surprises [Gre11]. symbols [DW01]. Symmetric [GL94, LD96, WP05]. Symmetry [LCDF10]. Synchrosqueezed [DLW09, DLW11, YLB^{+}15, DWtW15, DWtW16]. Synchrosqueezing [WHB^{+}14, WLD^{+}16, WFD11]. synthesis [Abo94]. system [AD79a, Abo94, AD78a, AD78b, AD79d, Red15, Sud16]. systematics [FGDB17]. systems [AD78b, AD79c, AD83, DDK05, HŚ02b, HŚ02a, TSS04, WP05].

table [RM95]. Tale [YGLD16, YGLD17]. Tap [CKE17, MMC^{+}13, SK12]. Technique [BBL^{+}11, CD96, CD97a, DAR13, SM12, ST15, TDK15]. Techniques [Wel99b, Dau16, LDN^{+}08, LDN^{+}10]. Technology [JDB^{+}14, BDKP14]. Ten [Dau92, Dau01, Hei92, Bat93a, C.93, Gri95, Grü92, Hei93]. tensor [AD78b, LP01]. Texas. [Dau93b]. Their [Lun92, RBC^{+}92, SP17, CD92, Dau87, SLBD11a, SLBD11b]. theme [Dau93c]. theorem [CS99, DH94, Kar03, Kar04]. theorems [DJ93]. Theoretic [HD06]. Theoretical [ZGSD04, ZGSD06]. theories [BS94]. Theory [HW06, Lun92, MRB^{+}14a, RGMD15a, CT15, HDH16, MRB^{+}14b, RGMD15b]. Thermal [BBN^{+}10b]. thermodynamic [DL94a, DL94b]. Thought [Dau05]. Three [GKBD19]. Three-Dimensional [GKBD19]. threshold [XTZ10]. thresholding [DDD03, DDD04, TDK15]. Time [CHT98, Dau88b, DP88, Dau97b, DWtW15, AIK10a, DK86, Dau89b, Dau90b, Dau95a, DLL95, Dau97a, Dau98e, Dau06c, DWtW16, KLC05, LG08, LD11a, LD11b, NKM12, AIK10b, YLB^{+}15]. time-domain [KLC05]. Time-Frequency [Dau97b, CHT98, Dau89b, Dau95a, DLL95, Dau98e, Dau06c, NKM12, YLB^{+}15]. time-scale [Dau97a]. Tomographic [LNDD06, LNDD07, SLN^{+}11a, SLN^{+}11b]. tomography [CVN^{+}13, LDN^{+}08, LDN^{+}10]. Tool [DLW09, Dau89b, DLW11, OMOE14a, OMOE14b]. Tools [Dau97b]. Tooth [BBL^{+}11, BBJ^{+}09]. Topic [WPS^{+}14, WPS^{+}13]. traces [Du01]. Transfer [YMH^{+}16]. Transform [BBN^{+}10b, DG80, DC11, MMN^{+}11, SM12, WLD^{+}16, YKIK04, ZGSD04, Ano16a, ABMD90, ABMD92, CD97b, Dau78c, DGR83, Dau90b, DM96, Dau98e, Dau06c, DWtW15, DWtW16, DAR13, DA15, KTJ09, KKJ^{+}10, LK91, MDSW92, NKM12, NY15, ST15, Sud16, ZGSD06]. Transformasi [SG13]. Transformation [Ano11e, JZL98, RV09b, AIK10a, Lin97a, AIK10b, AIK10a]. Transformations [Ano11d, Van08, Dau80a]. Transforms [CHT98, DLW09, RB98a, YLB^{+}15, CDSY97, CDSY98, CDV93, DMW92, Dau93d, DS98, DS00, DP02, DLW11, Du01, SLBD11a, SLBD11b]. Transient [NKM12]. Transitive [GYD^{+}18]. Transportation [LD09]. Traveling [JDB^{+}14]. treatment [PPR^{+}15]. Tree [CDDD01]. trellis [MDSW92]. True [DK86]. truncation [DH95]. Tumors [MMN^{+}11]. tuning [SD23]. Tutorial [Lun92]. TV [SP17]. Two [DL91, DL92b, DJ93, Dau94b, YGLD16, YGLD17, AD78b, AADL13, BAE11, DH04, Dau11, WFD11, XZCM06, APD06]. two-point [XZCM06]. Two-scale [DL91, DL92b, BAE11]. Type [HN15, BDS98, DT14, HSZ97, KK10, KKT10, LP01, LW09, PM13b, PM15, RF09a, RF09b]. typical [Huy08]. typische [Huy08].

ultra [CYV^{+}13a, CYV^{+}13b]. Ultrasound [MMN^{+}11, GEV12, KKJ^{+}10]. Uncertainty [Nov98, Dau83b]. Uncovering [WMJ^{+}11]. underpainting [ABD^{+}11]. underwater [PR05]. unified [HN17]. University [Ano05]. Unsigned [ADGT17, ADGT16]. use [Dau87, DP88]. used [Dau16, Nov02a, Nov02b]. Using [BBN^{+}10b, Dau95b, Dau16, MMN^{+}11, P15, RCP^{+}11, WLW06, WC02, WPS^{+}14, YLB^{+}15, AD78b, Ano08, ABMD90, ABMD92, BD03, BS94, CDSY97, CGB^{+}15, CHXL06, DD03, DA15, EBJ^{+}14, GEV12, JPB^{+}09, KAB11, KKJ^{+}10, KLC05, LG08, LQLC10a, LNDD06, LNDD07, MMC^{+}13, MDSW92, NY15, NM13, PMK16, PKG13, PM13b, PSB^{+}16, PJB^{+}09, RGMD15a, RGMD15b, RL97, RSD07a, RSD08, SHN10, SD23, ST15, Sud16, TDK15, VP13, WWFW98, WPS^{+}13, XZCM06]. Utah [SC00, SC02]. utilizando [VP13].

value [XZCM06]. Valued [ADGT17, ADGT16, DKLR14]. Values [Ma16]. variable [PM13a]. variables [DKP87]. Variants [HPH09]. Variation [GYD^{+}18]. Variational [BS94, DT05, RRD05, RRD12, DT04]. Variations [Dau93c]. vector [ABMD90, MDSW92]. vectors [CDP97]. vejvletam [Dau01]. Ventilator [WHB^{+}14]. verrast [Gre11]. versus [UD97]. very [DD03]. via [DWtW15, DWtW16, DMC^{+}16c, DMC^{+}16a, HŚ02b, HŚ02a, WHB^{+}14]. view [Dau96]. VII.3 [Dau08]. village [Gre11]. Vincent [JHB^{+}08]. Virtual [ABD^{+}11, RCP^{+}11, CRG^{+}13]. Visual [BBN^{+}10b, RM95]. vitesse [Dur96]. VLSI [LK91, MMC^{+}13]. vol [DL01a]. Volume [Lun92]. vs [SYSP11, SYSP12]. VUB [Ano05].

W [Ano11a]. Wajah [P15]. Wasserstein [LD11a, LD11b, LPD11, LPD13]. Watermarking [PMK16]. Wave [OWW^{+}16, WWD^{+}15, WC02, WLD^{+}16, PKG13, PR05]. Wavefield [CWC04]. wavefunctions [NG05, NG06]. Wavelet [ADGT17, Ano08, Ano11d, BBN^{+}10a, BBN^{+}10b, CDSY98, CHT98, CD93c, CKE17, Dau93d, Dau98e, DT04, DLW09, DC11, HW06, Kai10, KAB11, Lun92, MDSW92, MMN^{+}11, P15, RB98a, RRD05, SHN10, Sin13, Sud16, SG13, Van08, WLW06, Wel99b, YKIK04, YD16, YD17, AAI13, Abo94, AIK10a, ADGT16, Ano11c, Ano16a, Anoxx, ABMD90, ABMD92, Ara13, CDSY97, CGB^{+}15, CHXL06, CT15, Cob01, CD92, CD93a, CD93b, CDV93, CDGO02, CD97b, Dau90b, DMW92, Dau94b, DM96, DS98, Dau98c, Dau98d, DS00, DH02, DHRS03, DH04, Dau06c, DLW11, DMV09, DAR13, DA15, EGL11, EGL13, Gao14, HŚ02b, HŚ02a, JZL98, Kar03, Kar04, KTJ09, KKJ^{+}10, LK91, LP01, LW09, LXDS11, Lin97a, LLC08, LQLC10b, LQLC10a, LLC11, LD16, LNDD06, LNDD07, MMC^{+}13, NG05, NG06, NY15, OMOE14a]. wavelet [OMOE14b, PR05, RB98b, ST15, SLBD11a, SLBD11b, Tas00, TB94, VBU07, AIK10b, WP05, Wan07, XTZ10, YD18, Zha00, ZCY16]. Wavelet-based [DT04, CDGO02, HŚ02b, HŚ02a, LQLC10b, LLC11]. wavelet-like [SLBD11a, SLBD11b]. Wavelets [ACV01, Ano11b, BB07, CDV93, CK96, DP87, Dau88a, Dau89b, Dau90a, Dau92, Dau93b, Dau93f, Dau93e, Dau95d, Dau95e, Dau95c, DGSS99, Dau01, Dau08, DD11, GLG94, HPH09, LD96, Ma16, MRB^{+}14a, Nie99a, Nie99b, Nie12, NM13, P15, Qix12, RGMD15a, RV09a, RBC^{+}92, SP17, SLBD11a, SLBD11b, Tay08, VBU05a, Wah11, Wel99a, Won11, WPS^{+}14, Ano99, Ano16a, AC14, Bat93b, BS94, Boe01, BSP98, ČF04, CXS04, CLG04, CDF92, CDJV93, CD93d, Daa93, Dau89a, Dau93c, Dau94a, Dau94b, Dau96, DG99, DT05, Dau06b, DT14, Dor94, DKB99, DL96, Dur96, DKLR12, DKLR14, EBJ^{+}14, Fin04a, Fin04b, GL94, GNG^{+}08a, GNG^{+}08b, GVO^{+}11, GB95b, GB95a, GEV12, JPB^{+}09, Jam96, Kar10, KM12, KK10, KKT10, KD96, Koz06, Lai95, LG08, LL00, LM93, LM94]. wavelets [LM95, Lin97b, Lin98, Lu97, MD06, MYN07, MRB^{+}14b, Nov95, Nov98, Nov02a, Nov02b, Nov02c, Old92, PMK16, PKG13, PM15, PSB^{+}16, PM96, RGMD15b, Red15, RLS96, RL97, RF09a, RF09b, RA95, SS98, SLN^{+}11a, SLN^{+}11b, SYSP11, SYSP12, Str92, Tas99, Tem96, Tem97, VBU05b, WWFW98, WPS^{+}13, XZCM06, XWL07, Yan12, lZqJmTjZ08, Dau93a, Grü92, Lun92, Bat93a, C.93, Gri95, Grü92, Hei92, Hei93, Dau93a, Lun92]. waves [Dau93e]. way [BBJ^{+}09]. Weaning [WHB^{+}14]. Weave [YLB^{+}15]. weighted [CD97b, DDFG08a, DDFG08b, DDFG08c]. weights [PM15]. Wexler [DLL95]. Weyl [BD03, Dau80b, Dau83a, Du01]. whale [LG08]. Where [Dau96]. Which [DL92a, BBJ^{+}09, DL01b, DL01a]. White [DC11, XTZ10]. white-noise [XTZ10]. Who [BBJ^{+}09]. Wide [PR05]. Wiener [DK85, DKP87, Dau91, KD82b, KD84].

Wilson [DJJ91a, DJJ91b, NKM12]. windowed [BDV00]. Winner [Duk14]. wire [lZqJmTjZ08]. Wiskundige [Ano05, Gre11, Huy08]. without [LK91]. Women [CL05, Ano14, Dau95d]. Work [ACD^{+}13, Dau10, Dau16]. Workshop [JDB^{+}14, DKRS12, DKRS15]. world [Sta15]. writing [Bon16].

X [ABD^{+}11, ABD^{+}13, DMC^{+}16c,

DMC^{+}16a, FCY^{+}17, YDC^{+}14, YCF^{+}16]. X-Ray [YCF^{+}16, ABD^{+}11, ABD^{+}13, DMC^{+}16c, DMC^{+}16a, FCY^{+}17, YDC^{+}14]. xf [Pol92b, Pol92c].

year [Dau16]. Yves

[Dau93a, Lun92, Dau93a, Dau10].

Zel’dovich [PAHD05, PAHD04]. Zernike [Sud16]. zeros [Boe01, Kar10, Nov02c, Tem96, Tem97]. Zwanzig [PS95].

References

Al-Aifari:2013:CPD

[AADL13] Reema Al-Aifari, Ingrid Daubechies, and Yaron Lipman. Continuous Procrustes distance between two surfaces. Communications on Pure and Applied Mathematics (New York), 66 (6):934–964, June 2013. CO-DEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Ababneh:2013:HDW

[AAI13] Faisal Ababneh, S. Al Wadi, and Mohd Tahir Ismail. Haar and Daubechies wavelet methods in modeling banking sector. International Mathematical Forum, 8(9–12):551–566, 2013. ISSN 1312-7594 (print), 1314-7536 (electronic).

Anitha:2011:VUR

[ABD^{+}11] A. Anitha, A. Brasoveanu, M. F. Duarte, Shannon M. Hughes, Ingrid Daubechies, J. Dik, K. Janssens, and M. Alfeld. Virtual underpainting reconstruction from Xray fluorescence imaging data. In 2011 19th European Signal Processing Conference, volume ????, pages 1239–1243. ????, ????, August 2011. ISSN 2076-1465.

Anitha:2013:RXR

[ABD^{+}13] Anila Anitha, Andrei Brasoveanu, Marco Duarte, Shannon Hughes, Ingrid Daubechies, Joris Dik, Koen Janssens, and Matthias Alfeld. Restoration of X-ray fluorescence images of hidden paintings. Signal Processing, 93(3, SI):592–604, March 2013. CODEN SPRODR. ISSN 0165-1684 (print), 1872-7557 (electronic).

Antonini:1990:ICU

[ABMD90] Marc Antonini, Michel Barlaud, Pierre Mathieu, and Ingrid Daubechies. Image coding using vector quantization in the wavelet transform domain. In International Conference on Acoustics, Speech, and Signal Processing, volume 4, pages 2297–2300. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 1990. ISSN 1520-6149 (print), 2379-190X (electronic).

Antonini:1992:ICU

[ABMD92] Marc Antonini, Michel Barlaud, Pierre Mathieu, and Ingrid Daubechies. Image coding using wavelet transform. IEEE Transactions on Image Processing, 1(2): 205–220, April 1992. CO-DEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

URL http://ieeexplore.
ieee.org/abstract/document/
136597/.

Aboulnasr:1994:MBD

[Abo94] T. Aboulnasr. A mixed Butterworth/Daubechies wavelet analysis/synthesis system. In Mary Wallace, editor, 1994 IEEE International Symposium on Circuits and Systems: ISCAS ’94: the New Connaught Rooms, London, England, May 30–June 2, 1994, volume 6, pages 453–456. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1915-X (paperback), 0-7803-1916-8 (hardcover), 0-7803-1917-6 (microfiche). LCCN TK454.2 .In8 1994.

Arvinti:2014:PDM

[AC14] B. Arvinti and M. Costache. The performance of the Daubechies mother wavelets on ECG compression. In 2014 11th International Symposium on Electronics and Telecommunications (ISETC). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, November 2014.

Abry:2013:SII

[ACD^{+}13] Patrice Abry, Jim Coddington, Ingrid Daubechies, Ella Hendriks, Shannon Hughes, Don H. Johnson, and Eric Postma. Special issue: Image processing for digital art work foreword. Signal Processing, 93(3, SI):525–526, March 2013. CO-DEN SPRODR. ISSN 0165-1684 (print), 1872-7557 (electronic).

Aboufadel:2001:BDS

[ACV01] Edward Aboufadel, Amanda Cox, and Amy Vander Zee. Bivariate Daubechies scaling functions (wavelets). arXiv.org, ??(??):1–10, March 28, 2001. CODEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/math/0103198.

Aerts:1978:ASP

[AD78a] Dirk Aerts and Ingrid Daubechies.

About the structure-preserving maps of a quantum mechanical propositional system. Helvetica Physica Acta, 51(5–6): 637–660 (1979), 1978. CO-DEN HPACAK. ISSN 0018-0238 (print), 2297-1971 (electronic).

Aerts:1978:PJU

[AD78b] Dirk Aerts and Ingrid Daubechies.

Physical justification for using the tensor product to describe two quantum systems as one joint system. Helvetica Physica Acta, 51(5–6): 661–675 (1979), 1978. CO-DEN HPACAK. ISSN 0018-0238 (print), 2297-1971 (elec-

                                        tronic). URL https://www.e-
                                        periodica.ch/digbib/view?
                                        pid=hpa-001:1978:51#679.

AERTS:1979:SPM

[AD79a] Dirk Aerts and Ingrid Daubechies. About the structure-preserving maps of a quantum mechanical propositional system. Helvetica Physica Acta, 51 (5–6):637–660, 1979. CO-DEN HPACAK. ISSN 0018-

                                        0238. URL https://www.e-
                                        periodica.ch/digbib/view?
                                        pid=hpa-001:1978:51#655.

Aerts:1979:CSP

[AD79b] Dirk Aerts and Ingrid Daubechies. A characterization of subsystems in physics. Letters in Mathematical Physics, 3(1):11– 17, 1979. CODEN LMPHDY. ISSN 0377-9017 (print), 1573-0530 (electronic).

Aerts:1979:CBP

[AD79c] Dirk Aerts and Ingrid Daubechies. A connection between propositional systems in Hilbert spaces and von Neumann algebras. Helvetica Physica Acta, 52 (2):184–199, 1979. CODEN HPACAK. ISSN 0018-0238 (print), 2297-1971 (electronic).

Aerts:1979:MCS

[AD79d] Dirk Aerts and Ingrid Daubechies.

A mathematical condition for a sublattice of a propositional system to represent a physical subsystem, with a physical interpretation. Letters in Mathematical Physics, 3(1):19– 27, 1979. CODEN LMPHDY. ISSN 0377-9017 (print), 1573-0530 (electronic).

Aerts:1983:SPS

[AD83] Dirk Aerts and Ingrid Daubechies.

Simple proof that the structure preserving maps between quantum mechanical propositional systems conserve the angles. Helvetica Physica Acta, 56(6):1187–1190, 1983. CO-DEN HPACAK. ISSN 0018-0238 (print), 2297-1971 (electronic).

Aizenman:2024:MMB

[ADG^{+}24] Michael Aizenman, Ingrid Daubechies, Mary Gray, Cathy Kessel, Harriet Pollatsek, Graeme Smith, Elisabeth Werner, and Bei Zeng. In memory of Mary Beth Ruskai. Notices of the American Mathematical Society, 71(10):1362– 1371, October 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Alaifari:2016:RRV

[ADGT16] Rima Alaifari, Ingrid Daubechies, Philipp Grohs, and Gaurav Thakur. Reconstructing realvalued functions from unsigned coefficients with respect to wavelet and other frames. arXiv.org, ??(??):1–14, June 27, 2016. CODEN ???? ISSN 2331-8422. URL https://

                                        arxiv.org/abs/1601.07579.

Alaifari:2017:RRV

[ADGT17] Rima Alaifari, Ingrid Daubechies, Philipp Grohs, and Gaurav Thakur. Reconstructing realvalued functions from unsigned coefficients with respect to wavelet and other frames. Journal of Fourier Analysis and Applications, 23(6):1480– 1494, December 2017. ISSN 1069-5869 (print), 1531-5851 (electronic).

Alaifari:2016:SPR

[ADGY16] Rima Alaifari, Ingrid Daubechies, Philipp Grohs, and Rujie Yin. Stable phase retrieval in infinite dimensions. arXiv.org, ??(??):1–27, August 31, 2016. CODEN ???? ISSN 2331-

                                        8422. URL https://arxiv.
                                        org/abs/1609.00034.

Alaifari:2019:SPR

[ADGY19] Rima Alaifari, Ingrid Daubechies, Philipp Grohs, and Rujie Yin. Stable phase retrieval in infinite dimensions. Foundations of Computational Mathematics. The Journal of the Society for the Foundations of Computational Mathematics, 19(4): 869–900, 2019. ISSN 1615-3375,1615-3383.

Asadzadeh:2013:ECD

[AHK13] M. Asadzadeh, E. Hashemi, and A. Kozakevicius. Efficiency of combined Daubechies and statistical parameters applied in mammography. Applied and Computational Mathematics, 12(3):289–305, 2013. ISSN 1683-3511.

AlWadi:2010:CBD

[AIK10a] S. Al Wadi, Mohd Tahir Ismail, and Samsul Ariffin Abdul Karim. A comparison between the Daubechies wavelet transformation and the Fast Fourier Transformation in analyzing insurance time series data. Far East Journal of Applied Mathematics, 45(1):53– 63, 2010. ISSN 0972-0960.

Wadi:2010:CBD

[AIK10b] S. Al Wadi, Mohd Tahir Ismail, and Samsul Ariffin Abdul Karim. A comparison between the Daubechies wavelet transformation and the fast Fourier transformation in analyzing insurance time series data. Far East Journal of Applied Mathematics, 45(1):53– 63, 2010. ISSN 0972-0960.

Amat:2008:SLM

[AM08] Sergio Amat and María Moncayo. l^{∞}-stability for linear multiresolution algorithms: a new explicit approach. I. The basic rules and the Daubechies case. Applied Mathematics and Computation, 206(1):74– 91, 2008. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:1993:DSA

[Ano93]  Anonymous.  Daubechies

and Schneider among 1992 MacArthur Fellows. Physics Today, 46(1):83–84, January 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1998:ASA

[Ano98] Anonymous. Academy of Sciences adds to its ranks. Physics Today, 51(7):80, July 1998. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https:

                                        //physicstoday.scitation.
                                        org/doi/abs/10.1063/1.2805878.

Anonymous:1999:DWa

[Ano99]  Anonymous.  Daubechies

wavelets. In A Primer on Wavelets and Their Scientific Applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, March 1999.

Anonymous:2000:CMP

[Ano00]  Anonymous.  Contributions

to medical physics honored. Physics Today, 53(11):71, November 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

                                        URL https://physicstoday.
                                        scitation.org/doi/abs/10.
                                        1063/1.2405515. News of Ed-

uard Rhein Foundation Basic Research Award (2000) to Ingrid Daubechies “for the invention, the mathematical development, and the applications of wavelets, which have found widespread applications as orthogonal functions in signal processing, radar, and image processing”.

Anonymous:2005:WID

[Ano05] Anonymous. Wiskundige Ingrid Daubechies komt naar VUB. (Dutch) [Mathematician Ingrid Daubechies is coming to the Free University of Brussels]. Het Nieuwsblad, ??(??): ??, June 14, 2005. URL https:

//www.nieuwsblad.be/cnt/
gnifk3me.

Anonymous:2008:ABF

[Ano08]  Anonymous.  Analysis of

brain function and classification of sleep stage EEG using Daubechies wavelet. Sensors and Materials, page 1, 2008.

Anonymous:2011:SPH

[Ano11a]  Anonymous.  2011 Steele

Prizes [Henryk Iwaniec, Ingrid Daubechies, John W. Milnor]. Notices of the American Mathematical Society, 58(4): 593–596, April 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/201104/rtx110400593p.
pdf.

Anonymous:2011:DSF

[Ano11b] Anonymous. Daubechies scaling functions and wavelets. In Wavelet Theory, pages 233– 276. John Wiley & Sons, Inc., September 2011.

Anonymous:2011:DW

[Ano11c]  Anonymous.  Daubechies

wavelet. Wikipedia article, 2011. URL https:

                                        //en.wikipedia.org/wiki/
                                        Daubechies_wavelet.

Anonymous:2011:DWT

[Ano11d]  Anonymous.  Daubechies

wavelet transformations. In Discrete Wavelet Transformations, pages 223–280. John Wiley & Sons, Inc., March 2011.

Anonymous:2011:DDT

[Ano11e]  Anonymous.  The discrete

Daubechies transformation and applications. In Wavelet Theory, pages 277–323. John Wiley & Sons, Inc., September 2011.

Anonymous:2012:DAN

[Ano12]  Anonymous.  Daubechies

awarded Nemmers Prize. Notices of the American Mathematical Society, 59(7):962, 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2014:MW

[Ano14] Anonymous. Maths is (also) for women. Web interview., July 29, 2014. URL https:

                                        //twas.org/article/maths-
                                        also-women.

Anonymous:2015:PL

[Ano15]  Anonymous.  Pauli Lec-

tures 2015. Web story,

2015.  URL http://www.
pauli-lectures.ethz.ch/archive/
lectures15.html.

Anonymous:2016:DWT

[Ano16a] Anonymous. Discrete wavelet transform: Daubechies wavelets. Wikipedia article, 2016. URL

https://en.wikipedia.org/
wiki/Discrete_wavelet_transform.

Anonymous:2016:ICI

[Ano16b]  Anonymous.  Interview at

CIRM: Ingrid Daubechies. Video interview (16m24s)., November 29, 2016. URL

https://www.youtube.com/
watch?v=lhR7K6xp2Cg.

Anonymous:2017:ID

[Ano17] Anonymous. Ingrid Daubechies.

Physics Today, 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:2018:ID

[Ano18] Anonymous. Ingrid Daubechies.

Notices of the American Mathematical Society, 65(3):260– 261, March 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
journals/notices/201803/
201803FullIssue.pdf. Spe-

cial issue in honor of Women’s History Month, with profiles of 27 women mathematicians.

Anonymous:20xx:IDM

[Anoxx] Anonymous. Ingrid Daubechies, mother of the wavelet. Web story, 20xx. URL http:/

/focusonbelgium.be/en/Do%
20you%20know%20these%20Belgians/
ingrid-daubechies-mother-
wavelet.

Anthoine:2006:DMD

[APD06] S. Anthoine, E. Pierpaoli, and Ingrid Daubechies. Deux méthodes de déconvolution et séparation simultanées: application à la reconstruction des amas de galaxies. (French) [Two methods of simultaneous deconvolution and separation: application to the reconstruction of clusters of galaxies]. Traitement du signal, 23(5–6): 439–447, 2006. ISSN 0765-0019 (print), 1958-5608 (electronic).

Arai:2013:MDH

[Ara13] Kohei Arai. Method for data hiding based on LeGall 5/3 (Cohen–Daubechies–Feauveau: CDF 5/3) wavelet with data compression and random scanning of secret imagery data. International Journal of Wavelets, Multiresolution and Information Processing, 11(4):1360006, 18, 2013. CODEN IJWMIP. ISSN 0219-6913 (print), 1793-690X (electronic).

Bettaibi:2011:ADL

[BAE11] Néji Bettaibi, Jilani Alaya, and Moufida Elguénichi. On a qanalogue of the Daubechies– Lagarias two-scale equation. Mathematical Sciences Research Journal, 15(4):91–103, 2011. ISSN 1537-5978.

Bates:1993:BRT

[Bat93a] Bruce J. Bates. Book review:

Ten Lectures on Wavelets, by Ingrid Daubechies. Journal of the Acoustical Society of America, 93(3):1671, March 1993. CODEN JASMAN. ISSN 0001-

4966. URL https://asa.
scitation.org/doi/abs/10.
1121/1.406784.

Battle:1993:KGP

[Bat93b] Guy Battle. Klein–Gordon propagation of Daubechies wavelets. Journal of Mathematical Physics, 34(3):1095–1109, March 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic).

Bagct:2007:CDG

[BB07] Ulas Bagct and Li Bai. A comparison of Daubechies and Gabor wavelets for classification of MR images. In 2007 IEEE International Conference on Signal Processing and Communications. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2007.

Bunn:2009:WCW

[BBJ^{+}09] Jonathan Bunn, Doug Boyer, Jukka Jernvall, Yaron Lipman, and Ingrid Daubechies. Who cares which way is up? An orientation and landmark-free quantification of tooth shape. Journal of Vertebrate Paleontology, 29(S):72A–73A, 2009. ISSN 0272-4634.

Bunn:2011:CDN

[BBL^{+}11] Jonathan M. Bunn, Doug M. Boyer, Yaron Lipman, Elizabeth M. St Clair, Jukka Jernvall, and Ingrid Daubechies. Comparing Dirichlet normal surface energy of tooth crowns, a new technique of molar shape quantification for dietary inference, with previous methods in isolation and in combination. American Journal of Physical Anthropology, 145(2):247–261, June 2011. CODEN AJPNA9. ISSN 0002-9483 (print), 1096-8644 (electronic).

Bhowmik:2010:FDW

[BBN^{+}10a] Mrinal Kanti Bhowmik, Debotosh Bhattacharjee, Mita Nasipuri, Dipak Kumar Basu, and Mahantapas Kundu. Fusion of Daubechies wavelet coefficients for human face recognition. arXiv.org, ??(??):1–7, July 5, 2010. CODEN ???? ISSN 2331-8422. URL https:

//arxiv.org/abs/1007.0621.

Bhowmik:2010:QBM

[BBN^{+}10b] Mrinal Kanti Bhowmik, Debotosh Bhattacharjee, Mita Nasipuri, Dipak Kumar Basu, and Mahantapas Kundu. Quotient based multiresolution image fusion of thermal and visual images using Daubechies wavelet transform for human face recognition. arXiv.org, ?? (??):1–10, July 5, 2010. CO-DEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/1007.0620.

Balan:2003:OSE

[BD03] Radu Balan and Ingrid Daubechies.

Optimal stochastic encoding and approximation schemes using Weyl–Heisenberg sets. In Advances in Gabor Analysis, Appl. Numer. Harmon. Anal., pages 259–320. Birkhäuser Boston Inc., Cambridge, MA, USA, 2003. ISBN 0-8176-4239-0 (hardcover).

Brodie:2007:SSM

[BDD^{+}07] Joshua Brodie, Ingrid Daubechies, Christine De Mol, Domenico Giannone, and Ignace Loris. Sparse and stable Markowitz portfolios. arXiv.org, ??(??):1– 17, July 31, 2007. CODEN ???? ISSN 2331-8422. URL https:

//arxiv.org/abs/0708.0046.

Brodie:2009:SSM

[BDD^{+}09] Joshua Brodie, Ingrid Daubechies, Christine De Mol, Domenico Giannone, and Ignace Loris. Sparse and stable Markowitz portfolios. Proceedings of the National Academy of Sciences of the United States of America, 106(30):12267– 12272, July 28, 2009. CO-DEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (elec-

tronic). URL http://www.
jstor.org/stable/40484125.

Bourguignon:2014:WSS

[BDKP14] Jean-Pierre Bourguignon, Ingrid Daubechies, Myung-Hwan Kim, and Youngah Park. Why STEM (science, technology, engineering and mathematics)?

In Proceedings of the International Congress of Mathematicians (ICM 2014), Seoul, Korea, August 13–21, 2014. Vol. I: Plenary lectures and ceremonies, pages 787–797. Kyung Moon Sa, Seoul, South Korea, 2014. ISBN 89-6105-804-5 (hardcover); 89-6105-803-7 (set).

Bi:1998:ABB

[BDS98] N. Bi, L. Debnath, and Q. Sun. Asymptotic behavior of M-band scaling functions of Daubechies type. Zeitschrift für Analysis und ihre Anwendungen, 17(4):813–830, 1998. ISSN 0232-2064 (print), 1661-4534 (electronic).

Balan:2000:ADW

[BDV00] Radu Balan, Ingrid Daubechies, and Vinay Vaishampayan. The analysis and design of windowed Fourier frame based multiple description source coding schemes. IEEE Transactions on Information Theory, 46(7):2491–2536, November 2000. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Balachandran:2014:DAH

[BGS14] Athira Balachandran, M. Ganesan, and E. P. Sumesh. Daubechies algorithm for highly accurate ECG feature extraction. In 2014 International Conference on Green Computing Communication and Electrical Engineering (ICGC-CEE). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, March 2014.

Boyer:2011:AAQa

[BLS^{+}11a] Doug M. Boyer, Yaron Lipman, Elizabeth St. Clair, Jesus Puente, Thomas Funkhouser, Biren A. Patel, Jukka Jernvall, and Ingrid Daubechies. Algorithms to automatically quantify the geometric similarity of anatomical surfaces. arXiv.org, ??(??):1–6, October 17, 2011. CODEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/1110.3649.  Pub-

lished in PNAS 2011 108 (45) 18221–18226.

Boyer:2011:AAQb

[BLS^{+}11b] Doug M. Boyer, Yaron Lipman, Elizabeth St. Clair, Jesus Puente, Biren A. Patel, Thomas Funkhouser, Jukka Jernvall, and Ingrid Daubechies. Algorithms to automatically quantify the geometric similarity of anatomical surfaces. Proceedings of the National Academy of Sciences of the United States of America, 108(45):18221–18226, November 8, 2011. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

URL http://www.jstor.org/
stable/41352695.

Boese:2001:DZP

[Boe01] F. G. Boese. On the distribution of the zeros of polynomials related to the Daubechies wavelets. Zeitschrift für Angewandte Mathematik und Mechanik, 81:709–710, 2001. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Bonato:2016:IMR

[Bon16] Anthony Bonato. The intrepid mathematician reading, writing and arithmetic: Interview with a mathematician: Ingrid Daubechies. Web site., December 7, 2016. URL https:

                                        //anthonybonato.com/2016/
                                        12/07/interview-with-a-mathematician-
                                        ingrid-daubechies/.

Bownik:2003:PD

[Bow03] M. Bownik. On a problem of Daubechies. Constructive Approximation, 19(2):179–190, March 2003. ISSN 0176-4276 (print), 1432-0940 (electronic).

Bellanca-Penel:2023:EAI

[BPC23] Pascal Bellanca-Penel and Hugues Chabot. Un entretien avec Ingrid Daubechies: partie

2. (French) [An interview with Ingrid Daubechies: part 2]. Gazette des Mathématiciens, (178):59–67, 2023. ISSN 0224-8999.

Boyer:2015:NFA

[BPG^{+}15] Doug M. Boyer, Jesus Puente, Justin T. Gladman, Chris Glynn, Sayan Mukherjee, Gabriel S. Yapuncich, and Ingrid Daubechies. A new fully automated approach for aligning and comparing shapes.

The Anatomical Record — Advances in Integrative Anatomy and Evolutionary Biology, 298 (1(SI)):249–276, January 2015. ISSN 1932-8486 (print), 1932-8494 (electronic).

Best:1994:VDS

[BS94] Christoph Best and Andreas Schaefer. Variational description of statistical field theories using Daubechies’ wavelets. arXiv.org, ??(??):1–21, February 16, 1994. CODEN ???? ISSN 2331-8422. URL https:

//arxiv.org/abs/hep-lat/
9402012.

Brito:1998:DWQ

[BSP98] N. S. D. Brito, B. A. Souza, and F. A. C. Pires. Daubechies wavelets in quality of electrical power. In 8th International Conference on Harmonics and Quality of Power. Proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. IEEE Catalog Number 98EX227.

C:1993:BRT

[C.93] E. W. C. Book review: Ten Lectures on Wavelets, by Ingrid Daubechies. Mathematics of Computation, 61(204):941– 942, October 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.jstor.org/
stable/2153268.

Cahill:2016:PRIa

[CCD16a] Jameson Cahill, Peter G. Casazza, and Ingrid Daubechies. Phase retrieval in infinitedimensional Hilbert spaces. arXiv.org, ??(??):1–16, June 24, 2016. CODEN ???? ISSN 2331-8422. URL https://

                                        arxiv.org/abs/1601.06411.

Cahill:2016:PRIb

[CCD16b] Jameson Cahill, Peter G. Casazza, and Ingrid Daubechies. Phase retrieval in infinitedimensional Hilbert spaces. Transactions of the American Mathematical Society. Series B, 3:63–76, 2016. ISSN 2330-0000 (electronic).

Cohen:1992:SCB

[CD92] Albert Cohen and Ingrid Daubechies. A stability criterion for biorthogonal wavelet bases and their related subband coding scheme. Duke Mathematical Journal, 68(2): 313–335, November 1992. CO-DEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic).

Cohen:1993:NBW

[CD93a] Albert Cohen and Ingrid Daubechies. Non-separable bidimensional wavelet bases. Revista Matemática Iberoamericana, 9(1):51–137, 1993. ISSN 0213-2230 (print), 2235-0616 (electronic).

Cohen:1993:NSB

[CD93b] Albert Cohen and Ingrid Daubechies. Non-separable bidimensional wavelet bases. Revista Matemática Iberoamericana, 9(1):51–137, 1993. ISSN 0213-2230 (print), 2235-0616 (electronic).

Cohen:1993:IAB

[CD93c] Albert Cohen and Ingrid Daubechies. On the instability of arbitrary biorthogonal wavelet packets. SIAM Journal on Mathematical Analysis, 24(5):1340–1354, September 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1993:OBC

[CD93d] Albert Cohen and Ingrid Daubechies. Orthonormal bases of compactly supported wavelets. III. Better frequency resolution. SIAM Journal on Mathematical Analysis, 24(2): 520–527, March 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1996:NTE

[CD96] Albert Cohen and Ingrid Daubechies. A new technique to estimate the regularity of refinable functions. Revista Matemática Iberoamericana, 12(2):527–591, 1996. ISSN 0213-2230 (print), 2235-0616 (electronic). See corrigenda [CD97a].

Cohen:1997:CNT

[CD97a] Albert Cohen and Ingrid Daubechies. Corrigenda: “A new technique to estimate the regularity of refinable functions” [Rev. Mat. Iberoamericana 12 (1996), no. 2, 527– 591; MR1402677 (97g:42025)]. Revista Matemática Iberoamericana, 13(2):471, 1997. ISSN 0213-2230 (print), 2235-0616 (electronic). See [CD96].

Culik:1997:IDW

[CD97b] Karel Culik II and Simant Dube. Implementing Daubechies wavelet transform with weighted finite automata. Acta Informatica, 34(5):347–366, May 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cvetkovic:2000:SBO

[CD00] Z. Cvetkovic and Ingrid Daubechies. Single-bit oversampled A/D conversion with exponential accuracy in the bit-rate. In Storer and Cohn [SC00], pages 343–352. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000.

URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=838174. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Calderbank:2002:PCD

[CD02] A. Robert Calderbank and Ingrid Daubechies. The pros and cons of democracy. IEEE Transactions on Information Theory, 48(6):1721–1725, June 2002. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic). Special issue on Shannon theory: perspective, trends, and applications.

Cohen:2012:CRF

[CDD^{+}12] Albert Cohen, Ingrid Daubechies, Ronald DeVore, Gerard Kerkyacharian, and Dominique Picard. Capturing ridge functions in high dimensions from point queries. Constructive Approximation, 35(2):225–243, April 2012. ISSN 0176-4276 (print), 1432-0940 (electronic).

Cohen:2001:TAO

[CDDD01] Albert Cohen, Wolfgang Dahmen, Ingrid Daubechies, and Ronald DeVore. Tree approximation and optimal encoding. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 11(2):192–226, September 2001. ISSN 1063-5203 (print), 1096-603x (electronic).

Cohen:2003:HAS

[CDDD03] Albert Cohen, Wolfgang Dahmen, Ingrid Daubechies, and Ronald DeVore. Harmonic analysis of the space BV. Revista Matemática Iberoamericana, 19(1):235–263, 2003. ISSN 0213-2230 (print), 2235-0616 (electronic).

Cheng:2021:SPR

[CDDL21] Cheng Cheng, Ingrid Daubechies, Nadav Dym, and Jianfeng Lu. Stable phase retrieval from locally stable and conditionally connected measurements. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 55:440–465, 2021. ISSN 1063-5203 (print), 1096-603X (electronic).

Cohen:1992:BBC

[CDF92] Albert Cohen, Ingrid Daubechies, and J.-C. C. Feauveau. Biorthogonal bases of compactly supported wavelets. Communications on Pure and Applied Mathematics (New York), 45 (5):485–560, June 1992. CO-DEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Cohen:2002:ICW

[CDGO02] Albert Cohen, Ingrid Daubechies, Onur G. Guleryuz, and Michael T. Orchard. On the importance of combining wavelet-based nonlinear approximation with coding strategies. IEEE Transactions on Information Theory, 48(7):1895–1921, July 2002. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Cohen:1993:MAW

[CDJV93] Albert Cohen, Ingrid Daubechies, Björn Jawerth, and Pierre Vial. Multiresolution analysis, wavelets and fast algorithms on an interval. Comptes rendus de l’Académie des sciences. Série I, Mathématique, 316(5): 417–421, March 4, 1993. CO-DEN CASMEI. ISSN 0764-4442 (print), 1778-3577 (electronic).

Cvetkovic:2002:IBF

[CDL02] Z. Cvetkovic, Ingrid Daubechies, and B. F. Logan. Interpolation of bandlimited functions from quantized irregular samples. In Storer and Cohn [SC02], pages 412–421. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002.

URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=999981. IEEE

Computer Society Order Number PR01477.

Cvetkovic:2007:SBO

[CDL07] Zoran Cvetković, Ingrid Daubechies, and Benjamin F. Logan, Jr. Single-bit oversampled A/D conversion with exponential accuracy in the bit rate. IEEE Transactions on Information Theory, 53(11):3979– 3989, November 2007. CO-DEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic). Data Compression Conference (DCC 2002), Snowbird, UT, April 02–04, 2002.

Cohen:1997:RRF

[CDP97] Albert Cohen, Ingrid Daubechies, and Gerlind Plonka. Regularity of refinable function vectors. Journal of Fourier Analysis and Applications, 3(3):295– 324, 1997. ISSN 1069-5869 (print), 1531-5851 (electronic).

Cohen:1996:HSS

[CDR96] Albert Cohen, Ingrid Daubechies, and Amos Ron. How smooth is the smoothest function in a given refinable space? Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 3(1):87–89, January 1996. ISSN 1063-5203 (print), 1096-603x (electronic).

Calderbank:1997:LIC

[CDSY97] A. Robert Calderbank, Ingrid Daubechies, Wim Sweldens, and Boon-Lock Yeo. Lossless image compression using integer to integer wavelet transforms. In IEEE, editor, Proceedings of International Conference on Image Processing, 26–29 October 1997, pages 596–599. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-8184-5 (hardcover), 0-8186-8183-7, 0-8186-8185-3 (microfiche). LCCN TA1637 I61 1997.

URL http://ieeexplore.
ieee.org/servlet/opac?punumber=
4998.

Calderbank:1998:WTM

[CDSY98] A. Robert Calderbank, Ingrid Daubechies, Wim Sweldens, and Boon-Lock Yeo. Wavelet transforms that map integers to integers. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 5(3):332–369, July 1998. ISSN 1063-5203 (print), 1096-603x (electronic).

Cohen:1993:WIF

[CDV93] Albert Cohen, Ingrid Daubechies, and Pierre Vial. Wavelets on the interval and fast wavelet transforms. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 1(1):54–81, 1993. ISSN 1063-5203 (print), 1096-603x (electronic).

Clemens:2014:NIN

[CDW14] Herbert Clemens, Ingrid Daubechies, and Carol Wood. The new IMU needs you! Notices of the American Mathematical Society, 61(6):632–534, June/July 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cerna:2004:CSC

[ČF04] Dana Černá and Václav Finěk. On the computation of scaling coefficients of Daubechies’ wavelets. Central European Journal of Mathematics, 2(3): 399–419, 2004. ISSN 1644-3616 (print), 1895-1074 (electronic).

Chatterjee:2015:MIF

[CGB^{+}15] Pubali Chatterjee, Somoballi Ghoshal, Biswajit Biswas, Amlan Chakrabarti, and Kashi Nath Dey. Medical image fusion using Daubechies complex wavelet and near set. In Transactions on Computational Science XXV, volume 9030 of Lecture Notes in Comput. Sci., pages 90–111. Springer, Heidelberg, 2015.

Carmona:1998:PTF

[CHT98] René Carmona, Wen-Liang Hwang, and Bruno Torrésani. Practical Time-frequency Analysis. Gabor and Wavelet Transforms with an Implementation in S. Academic Press, New York, NY, USA, 1998. ISBN 0-12-160170-6. 490 pp. With a preface by Ingrid Daubechies.

Chen:2006:DML

[CHXL06] Xuefeng Chen, Zhengjia He, Jiawei Xiang, and Bing Li. A dynamic multiscale lifting computation method using Daubechies wavelet. Journal of Computational and Applied Mathematics, 188(2):228– 245, 2006. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic).

Cohen:1996:WMB

[CK96] Albert Cohen and Jelena Kovačević. Wavelets: the mathematical background. Proceedings of the IEEE, 84(4): 514–522, April 1996. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Coskun:2017:MFI

[CKE17] Adem Coskun, Izzet Kale, and Yaprak Eminaga. Multiplier free implementation of 8-tap Daubechies wavelet filters for biomedical applications. In 2017 New Generation of CAS (NGCAS). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 2017.

Case:2005:CWM

[CL05] Bettye Anne Case and Anne M.

Leggett, editors. Complexities: Women in Mathematics. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-11462-5, 0-691-17109-2, 1-4008-8016-5 (e-book). xix + 412 + 22 pp. LCCN QA27.5 .C66

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0668/2004048843-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0668/2004048843-
d.html;   http://www.loc.
gov/catdir/toc/fy0612/2004048843.
html.

Clonda:2004:CDW

[CLG04] Diego Clonda, J.-M. Lina, and Bernard Goulard. Complex Daubechies wavelets: Properties and statistical image modelling. Signal Processing, 84 (1):1–23, January 2004. CO-DEN SPRODR. ISSN 0165-1684 (print), 1872-7557 (electronic).

Chassande-Mottin:1997:DR

[CMDAF97] E. Chassande-Mottin, Ingrid Daubechies, F. Auger, and P. Flandrin. Differential reassignment. IEEE Signal Processing Letters, 4(10):293– 294, October 1997. CODEN ISPLEM. ISSN 1070-9908 (print), 1558-2361 (electronic).

Coburn:2001:BIG

[Cob01] L. A. Coburn. The Bargmann isometry and Gabor–Daubechies wavelet localization operators. In Systems, approximation, singular integral operators, and related topics. Proceedings of the 11th international workshop on operator theory and applications, IWOTA 2000, Bordeaux, France, June 13–16, 2000, volume 129 of Oper. Theory Adv. Appl., pages 169–178. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2001. ISBN 3-7643-6645-1 (hardcover).

Cook:2011:ICD

[Coo11] Mariana Cook. Ingrid Chantal Daubechies. Mitteilungen der Deutschen Mathematiker-Vereinigung, 19(1):34–35, 2011. ISSN 0947-4471.

Cornelis:2013:CDI

[CRG^{+}13] B. Cornelis, T. Ružić, E. Gezels, A. Dooms, A. Pizurica, L. Platiša, J. Cornelis, M. Martens, M. De Mey, and Ingrid Daubechies. Crack detection and inpainting for virtual restoration of paintings: The case of the ghent altarpiece. Signal Processing, 93 (3, SI):605–619, March 2013. CODEN SPRODR. ISSN 0165-1684 (print), 1872-7557 (electronic).

Cai:1999:BRT

[CS99] Tianxi Cai and Jianhong Shen. Boundedness is redundant in a theorem of Daubechies. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 6(3):400–404, 1999. ISSN 1063-5203 (print), 1096-603x (electronic).

Chen:2015:SMA

[CT15] Jian Ping Chen and Wen Yong Tang. Study of a meshless analysis method for ship structures based on Daubechies wavelet basis theory. J. Dalian Univ. Technol., 55(5):498–503, 2015. ISSN 1000-8608.

Charlety:2013:GST

[CVN^{+}13] Jean Charlety, Sergey Voronin, Guust Nolet, Ignace Loris, Frederik J. Simons, Karin Sigloch, and Ingrid C. Daubechies. Global seismic tomography with sparsity constraints: Comparison with smoothing and damping regularization. Journal of Geophysical Research. Solid Earth, 118(9):4887–4899, September 2013. ISSN 2169-9313 (print), 2169-9356 (electronic).

Chen:2004:CAL

[CWC04] Ling Chen, Ru-Shan Wu, and Yong Chen. Construction and application of local scattering matrix based on wavefield extrapolation in Gabor– Daubechies beamlet domain. Chinese Journal of Geophysics, 47(2):322–332, 2004. ISSN 2326-0440.

Chen:2004:DSA

[CXS04] Duan Chen, Chang Fa Xu, and Yang Guang Sun. Deltasequence approach to solving PDE based on Daubechies wavelets. J. Huazhong Univ. Sci. Technol. Nat. Sci., 32(10): 114–116, 2004. ISSN 1671-4512.

Cornelis:2017:RCP

[CYG^{+}17] Bruno Cornelis, Haizhao Yang, Alex Goodfriend, Noelle Ocon, Jianfeng Lu, and Ingrid Daubechies. Removal of canvas patterns in digital acquisitions of paintings. IEEE Transactions on Image Processing, 26(1):160–171, January 2017. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Cornelis:2013:BCDa

[CYV^{+}13a] Bruno Cornelis, Yun Yang, Joshua T. Vogelstein, Ann Dooms, Ingrid Daubechies, and David Dunson. Bayesian crack detection in ultra high resolution multimodal images of paintings. arXiv.org, ??(??):1– 8, April 22, 2013. CODEN ???? ISSN 2331-8422. URL https:

//arxiv.org/abs/1304.5894.

Cornelis:2013:BCDb

[CYV^{+}13b] Bruno Cornelis, Yun Yang, Joshua T. Vogelstein, Ann Dooms, Ingrid Daubechies, and David Dunson. Bayesian crack detection in ultra high resolution multimodal images of paintings. In 2013 18th International Conference on Digital Signal Processing (DSP), volume ????, pages 1–8. ????, ????, July 2013. ISSN 1546-1874.

Dishabi:2015:DPP

[DA15] Mohammad Reza Ebrahimi Dishabi and Mohammad Abdollahi Azgomi. Differential privacy preserving clustering using Daubechies-2 wavelet transform. International Journal of Wavelets, Multiresolution and Information Processing, 13(4):1550028, 35, 2015. CODEN IJWMIP. ISSN 0219-6913 (print), 1793-690X (electronic).

Daalhuis:1993:CDW

[Daa93] Adri B. Olde Daalhuis. Computing with Daubechies’ wavelets. In Wavelets: an Elementary Treatment of Theory and Applications, volume 1 of Ser. Approx. Decompos., pages 93–105. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993.

Dishabi:2013:ITP

[DAR13] Mohammad Reza Ebrahimi Dishabi, Mohammad Abdollahi Azgomi, and Amir Masoud Rahmani. An improved technique for privacy preserving clustering based on Daubechies-2 wavelet transform. International Journal of Wavelets, Multiresolution and Information Processing, 11 (5):1350039, 42, 2013. CO-DEN IJWMIP. ISSN 0219-6913 (print), 1793-690X (electronic).

Daubechies:1977:AHO

[Dau78a] Ingrid Daubechies. An application of hyperdifferential operators to holomorphic quantization. Letters in Mathematical Physics, 2(6):459–469, 1977– 1978. CODEN LMPHDY. ISSN 0377-9017 (print), 1573-0530 (electronic).

Daubechies:1978:AHO

[Dau78b] Ingrid Daubechies. An application of hyperdifferential operators to holomorphic quantization. Letters in Mathematical Physics, 2(6):459–469, November 1978. CODEN LMPHDY.

ISSN 0377-9017 (print), 1573-0530 (electronic).

Daubechies:1978:ITR

[Dau78c] Ingrid Daubechies. An integral transform related to quantization. Technical report, CNRS, Paris, France, 1978.

Daubechies:1980:CSP

[Dau80a] Ingrid Daubechies. Coherent states and projective representation of the linear canonical transformations. Journal of Mathematical Physics, 21(6): 1377–1389, June 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/
jmapaq/v21/i6/p1377_s1.

Daubechies:1980:DCB

[Dau80b] Ingrid Daubechies. On the distributions corresponding to bounded operators in the Weyl quantization. Communications in Mathematical Physics, 75 (3):229–238, 1980. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916 (electronic).

URL http://projecteuclid.
org/euclid.cmp/1103908147.

Daubechies:1980:RQM

[Dau80c] Ingrid C. Daubechies. Representation of Quantum Mechanical Operators by Kernels on Hilbert Spaces of Analytic Functions. Ph.D. thesis, Fakulteit der Wetenschappen, Vrije Universiteit Brussel, Brussels, Belgium, 1980. ???? pp.

Daubechies:1983:CSC

[Dau83a] Ingrid Daubechies. Continuity statements and counterintuitive examples in connection with Weyl quantization. Journal of Mathematical Physics, 24(6):1453–1461, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

                                        //jmp.aip.org/resource/1/
                                        jmapaq/v24/i6/p1453_s1.

Daubechies:1983:UPF

[Dau83b] Ingrid Daubechies. An uncertainty principle for fermions with generalized kinetic energy. Communications in Mathematical Physics, 90(4): 511–520, 1983. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916 (electronic).

                                        URL http://projecteuclid.
                                        org/euclid.cmp/1103940414.

Daubechies:1984:OEM

[Dau84]  Ingrid Daubechies.  One-

electron molecules with relativistic kinetic energy: properties of the discrete spectrum. Communications in Mathematical Physics, 94(4): 523–535, 1984. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916 (electronic).

                                        URL http://projecteuclid.
                                        org/euclid.cmp/1103941407.

Daubechies:1987:DSC

[Dau87] Ingrid Daubechies. Discrete sets of coherent states and their use in signal analysis.

Lecture Notes in Mathematics, 1285:73–82, 1987. CO-DEN LNMAA2. ISBN 3-540-18479-1 (print), 3-540-47983-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/BFb0080582/.

Daubechies:1988:OBC

[Dau88a]  Ingrid Daubechies.  Or-

thonormal bases of compactly supported wavelets. Communications on Pure and Applied Mathematics (New York), 41(7):909–996, October 1988. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic). Reprinted in [Dau06b].

Daubechies:1988:TFLa

[Dau88b]  Ingrid Daubechies.  Time–

frequency localization operators: a geometric phase space approach. IEEE Transactions on Information Theory, 34(4): 605–612, July 1988. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Daubechies:1989:OBW

[Dau89a] Ingrid Daubechies. Orthogonal bases of wavelets with finite support — connection with discrete filters. In Wavelets. Time-frequency methods and phase space. Proceedings of the international conference, Marseille, France, December 14–18, 1987, Inverse Probl. Theoret. Imaging, pages 38–66. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51159-8 (hardcover).

Daubechies:1989:WTT

[Dau89b] Ingrid Daubechies. Wavelets: a tool for time-frequency analysis. In Sixth Multidimensional Signal Processing Workshop, page 98. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 1989. ISSN ????

Daubechies:1990:OBWa

[Dau90a] Ingrid Daubechies. Orthonormal bases of wavelets with finite support — connection with discrete filters. In Inverse Problems and Theoretical Imaging, pages 38–66. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990.

Daubechies:1990:WTT

[Dau90b] Ingrid Daubechies. The wavelet transform, time frequency localization and signal analysis. IEEE Transactions on Information Theory, 36(5):961– 1005, September 1990. CO-DEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic). Reprinted in [Dau06c].

Daubechies:1991:PSP

[Dau91] Ingrid Daubechies. Phase space path integrals, coherent states and Wiener measure. Rendiconti del Circolo Matematico di Palermo. Serie II. Supplemento, 25:157–176, 1991. ISSN 1592-9531. Proceedings of the Joint Concordia Sherbrooke Seminar Series on Functional Integration Methods in Stochastic Quantum Mechanics (Sherbrooke, PQ and Montréal, PQ, 1987).

Daubechies:1992:TLW

[Dau92] Ingrid Daubechies. Ten Lectures on Wavelets, volume 61 of CBMS-NSF Regional Conference Series in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-274-2 (paperback). xix + 357

pp. URL http://bookstore.
siam.org/cb61/.

Daubechies:1993:BRW

[Dau93a] Ingrid Daubechies. Book reviews: Wavelets and Operators — Meyer, Yves; Wavelets — Algorithms and Applications — Meyer, Yves. Science, 262(5139):1589–1591, December 3, 1993. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.jstor.org/
stable/2882490.

Daubechies:1993:DPW

[Dau93b] Ingrid Daubechies, editor. Different Perspectives on Wavelets: American Mathematical Society short course, January 11– 12, 1993 San Antonio, Texas., number 47 in Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1993. ISBN 0-8218-5503-4 (hardcover). ISSN 0160-7634. LCCN QA403.3 .D54 1993. Lecture notes prepared for the American Mathematical Society Short Course Wavelets and Applications.

Daubechies:1993:OBC

[Dau93c] Ingrid Daubechies. Orthonormal bases of compactly supported wavelets. II. Variations on a theme. SIAM Journal on Mathematical Analysis, 24 (2):499–519, March 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Daubechies:1993:WTO

[Dau93d] Ingrid Daubechies. Wavelet transforms and orthonormal wavelet bases. In Different Perspectives on Wavelets. American Mathematical Society Short Course on Wavelets and Applications, Held in San Antonio, TX (USA), January 11– 12, 1993 [Dau93b], pages 1– 33. ISBN 0-8218-5503-4 (hardcover). ISSN 0160-7634. LCCN QA403.3 .D54 1993. Lecture notes prepared for the American Mathematical Society Short Course Wavelets and Applications.

Daubechies:1993:WMW

[Dau93e] Ingrid Daubechies. Wavelets making waves in mathematics and engineering. American Mathematical Society, Providence, RI, USA, 1993. ISBN 0-8218-8082-9. 1 videocassette (NTSC; 1/2 inch; VHS) (60 min.); sd., col pp. A joint AMS-MAA invited address presented in Baltimore, Maryland, January 1992.

Daubechies:1993:WI

[Dau93f] Ingrid Daubechies. Wavelets on the interval. In Progress in wavelet analysis and applications. Proceedings of the 3rd international conference on wavelets and applications, Toulouse, France, June 8–13, 1992, pages 95–107. Frontières, Gif-sur-Yvette, France, 1993. ISBN 2-86332-130-7 (hardcover).

Daubechies:1994:ACS

[Dau94a]  Ingrid Daubechies.  Affine

coherent states and wavelets. In On Klauder’s Path: a Field Trip. Essays in Honor of John R. Klauder, pages 35–42. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1994. ISBN 981-02-1687-4.

Daubechies:1994:TRR

[Dau94b] Ingrid Daubechies. Two recent results on wavelets: wavelet bases for the interval, and biorthogonal wavelets diagonalizing the derivative operator. In Recent Advances in Wavelet Analysis, volume 3 of Wavelet Anal. Appl., pages 237–257. Academic Press, New York, NY, USA, 1994. ISBN 0-12-632370-4 (hardcover).

Daubechies:1995:BDF

[Dau95a]  Ingrid Daubechies.  Better

dual functions for Gabor timefrequency lattices. In Approximation theory VIII. Vol. 2. Wavelets and multilevel approximation. Papers from the 8th Texas international conference, College Station, TX, USA, January 8–12, 1995, volume 6 of Ser. Approx. Decompos., pages 113–116. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1995. ISBN 981-02-2972-0 (hardcover).

Daubechies:1995:UFD

[Dau95b] Ingrid Daubechies. Using Fredholm determinants to estimate the smoothness of refinable functions. In Approximation theory VIII. Vol. 2. Wavelets and multilevel approximation. Papers from the 8th Texas international conference, College Station, TX, USA, January 8– 12, 1995, volume 6 of Ser. Approx. Decompos., pages 89– 112. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1995. ISBN 981-02-2972-0 (hardcover).

Daubechies:1995:WOR

[Dau95c] Ingrid Daubechies. Wavelets: an overview, with recent applications. In Proceedings of 1995 IEEE International Symposium on Information Theory, pages 5–. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 1995. ISSN ????

Daubechies:1995:WAC

[Dau95d] Ingrid Daubechies. Wavelets and applications: a celebration of women in mathematics. Notices of the American Mathematical Society, 42 (1):34–35, January 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.
ams.org/journals/notices/
199501/wim.pdf.

Daubechies:1995:WOP

[Dau95e] Ingrid Daubechies. Wavelets and other phase space localization methods. In Proceedings of the International Congress of Mathematicians, ICM ’94, August 3–11, 1994, Zürich, Switzerland. Vol. I, pages 57– 74. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1995. ISBN 3-7643-5153-5 (hardcover).

Daubechies:1996:WDW

[Dau96] Ingrid Daubechies. Where do wavelets come from? A personal point of view. Proceedings of the IEEE, 84(4):510–513, April 1996. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Daubechies:1997:OFP

[Dau97a] Ingrid Daubechies. From the original framer to present-day time–frequency and time-scale frames. Journal of Fourier Analysis and Applications, 3 (5):485–486, 1997. ISSN 1069-5869 (print), 1531-5851 (electronic). Dedicated to the memory of Richard J. Duffin.

Daubechies:1997:TTF

[Dau97b] Ingrid Daubechies. Tools for time-frequency analysis. Report, Program in Applied and Computational Mathematics, Princeton University, Princeton, NJ, USA, September 30, 1997. 5

                                        pp. URL http://www.dtic.
                                        mil/dtic/tr/fulltext/u2/
                                        a337887.pdf.

Daubechies:1998:P

[Dau98a] Ingrid Daubechies. Preface. In Practical Time-frequency Analysis. Gabor and Wavelet Transforms with an Implementation in S [CHT98], page 5. ISBN 0-12-160170-6. With a preface by Ingrid Daubechies.

Daubechies:1998:Pa

[Dau98b] Ingrid Daubechies. Preface. In Practical Time-Frequency Analysis — Gabor and Wavelet Transforms with an Implementation in S, page 5. Elsevier, Amsterdam, The Netherlands,

Daubechies:1998:RRWa

[Dau98c] Ingrid Daubechies. Recent results in wavelet applications. In Harold H. Szu, editor, Wavelet Applications V, page ?? SPIE Optical Engineering Press, Bellingham, WA, USA, March 26, 1998.

Daubechies:1998:RRWb

[Dau98d] Ingrid Daubechies. Recent results in wavelet applications. Journal of Electronic Imaging, 7(4):719–724, October 1998. CODEN JEIME5. ISSN 1017-9909 (print), 1560-229x (electronic).

Daubechies:1998:WTT

[Dau98e] Ingrid Daubechies. Wavelet transform, time-frequency localization and signal analysis. ????, 25(13):42–??, ???? 1998.

Daubechies:2001:DLP

[Dau01] Ingrid Daubechies. Desyat’ lektsij po vejvletam. (Russian) [Ten Lectures on Wavelets]. NITS, Regulyarnaya i Khaoticheskaya Dinamika, Moskva, Russia, 2001. ISBN 5-93972-044-7. 464 pp.

Daubechies:2005:TP

[Dau05] Ingrid Daubechies. Thought problems. In Case and Leggett [CL05], pages 358– 360. ISBN 0-691-11462-5, 0-691-17109-2, 1-4008-8016-5 (ebook). LCCN QA27.5 .C66

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0668/2004048843-b.html;

 http://www.loc.gov/catdir/
enhancements/fy0668/2004048843-
d.html;   http://www.loc.
gov/catdir/toc/fy0612/2004048843.
html.

Daubechies:2006:F

[Dau06a] Ingrid Daubechies. Foreword. In Heil and Walnut [HW06], pages xv–xvi. ISBN 0-691-11453-6 (hardcover), 0-691-12705-0 (paperback), 1-4008-2726-4. With a foreword by Ingrid Daubechies. Introduction by John J. Benedetto.

Daubechies:2006:OBC

[Dau06b] Ingrid Daubechies. Orthonormal bases of compactly supported wavelets. In Heil and Walnut [HW06], pages 564–654. ISBN 0-691-11453-6 (hardcover), 0-691-12705-0 (paperback), 1-4008-2726-4. Reprint of [Dau88a].

Daubechies:2006:WTT

[Dau06c] Ingrid Daubechies. The wavelet transform, time-frequency localization and signal analysis. In Heil and Walnut [HW06], pages 442–486. ISBN 0-691-11453-6 (hardcover), 0-691-12705-0 (paperback), 1-4008-2726-4. Reprint of [Dau90b].

Daubechies:2008:VWA

[Dau08]  Ingrid Daubechies.  VII.3

wavelets and applications. In Gowers et al. [GBGL08], pages 848–862. ISBN 0-691-11880-9 (hardcover). LCCN QA11.2 .P745 2008. URL http:

//www.loc.gov/catdir/toc/
ecip0818/2008020450.html.

Daubechies:2010:WYM

[Dau10] Ingrid Daubechies. The work of Yves Meyer. In Proceedings of the International Congress of Mathematicians. Volume I, pages 115–124. Hindustan Book Agency, New Delhi, India, 2010. ISBN 981-4324-30-2 (set); 81-85931-08-9 (hardcover); 981-4324-31-0 (hardcover); 981-4324-35-3 (ebook).

Daubechies:2011:CPD

[Dau11] Yaron Lipman Reema Al-Aifari Ingrid Daubechies. The continuous Procrustes distance between two surfaces. arXiv.org, ??(??):1–21, June 22, 2011. CODEN ???? ISSN 2331-8422. URL https://

arxiv.org/abs/1106.4588.

Daubechies:2015:OAPa

[Dau15] Ingrid Daubechies. Opening address: President of IMU. In The Proceedings of the 12th International Congress on Mathematical Education, pages 3–4. Springer International Publishing, ????, 2015.

Daubechies:2016:UMR

[Dau16]  Ingrid Daubechies.  Us-

ing mathematics to repair a masterpiece: the author shows how new mathematical techniques can be used to revitalize a 650-year-old work of art. Quanta Magazine, ??(??):??, September 29,

2016. URL https://www.
quantamagazine.org/using-
mathematics-to-repair-a-
masterpiece-20160929.

Ding:2011:ADW

[DC11] Fang Ding and Tianjie Cao. Application of Daubechies wavelet transform in the estimation of standard deviation of white noise. In IEEE, editor, Proceedings of the 2011 Second International Conference on Digital Manufacturing & Automation (ICDMA 2011): 5–7 August 2011, Zhangjiajie, Hunan, China, pages 212–

215. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, August 2011. ISBN 1-4577-0755-1 (print), 0-7695-4455-X. LCCN TA345; TS183.

URL http://ieeexplore.
ieee.org/servlet/opac?punumber=

6050153. IEEE Computer Society Order Number E4455.

Daubechies:2003:ABF

[DD03] Ingrid Daubechies and Ronald DeVore. Approximating a bandlimited function using very coarsely quantized data: a family of stable sigma–delta modulators of arbitrary order. Annals of Mathematics (2), 158(2):679–710, September 2003. CODEN AN-MAAH. ISSN 0003-486x (print), 1939-8980 (electronic).

URL http://www.jstor.org/
stable/3597299.

Dooms:2011:W

[DD11] Ann Dooms and Ingrid Daubechies.

Wavelets, pages 135–154. Wiley-VCH Verlag GmbH & Co. KGaA, ????, 2011. ISBN 3-527-63524-6.

Daubechies:2003:ITA

[DDD03] Ingrid Daubechies, Michel Defrise, and Christine De Mol. An iterative thresholding algorithm for linear inverse problems with a sparsity constraint. arXiv.org, ??(??):1– 30, July 10, 2003. CO-DEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/math/0307152.

Daubechies:2004:ITA

[DDD04] Ingrid Daubechies, Michel Defrise, and Christine De Mol. An iterative thresholding algorithm for linear inverse problems with a sparsity constraint. Communications on Pure and Applied Mathematics (New York), 57(11):1413–1457, November 2004. CODEN CPA-MAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Daubechies:2016:SER

[DDD16] Ingrid Daubechies, Michel Defrise, and Christine De Mol. Sparsity-enforcing regularisation and ISTA revisited. Inverse Problems, 32(10):104001, 15, October 2016. CODEN IN-PEEY. ISSN 0266-5611 (print), 1361-6420 (electronic).

Daubechies:2023:NNA

[DDD^{+}23] Ingrid Daubechies, Ronald DeVore, Nadav Dym, Shira Faigenbaum-Golovin, Shahar Z. Kovalsky, Kung-Chin Lin, Josiah Park, Guergana Petrova, and Barak Sober. Neural network approximation of refinable functions. IEEE Transactions on Information Theory, 69(1):482–495, 2023. CO-DEN IETTAW. ISSN 0018-9448,1557-9654.

Daubechies:2022:NAD

[DDF^{+}22] I. Daubechies, R. DeVore, S. Foucart, B. Hanin, and G. Petrova. Nonlinear approximation and (deep) ReLU networks. Constructive Approximation, 55(1):127–172, 2022. ISSN 0176-4276,1432-0940.

Daubechies:2008:IRWa

[DDFG08a] Ingrid Daubechies, Ronald De-Vore, Massimo Fornasier, and C. Sinan Güntürk. Iteratively re-weighted least squares minimization for sparse recovery. Report, Program in Applied and Computational Mathematics, Princeton University, Princeton, NJ, USA, June 14, 2008. 36

                                        pp. URL http://www.dtic.
                                        mil/dtic/tr/fulltext/u2/
                                        a528510.pdf.

Daubechies:2008:IRWb

[DDFG08b] Ingrid Daubechies, Ronald De-Vore, Massimo Fornasier, and C. Sinan Güntürk. Iteratively re-weighted least squares minimization for sparse recovery. arXiv.org, ??(??):1–35, July 3, 2008. CODEN ???? ISSN 2331-8422. URL https://

arxiv.org/abs/0807.0575.

Daubechies:2008:IRWc

[DDFG08c] Ingrid Daubechies, Ronald De-Vore, Massimo Fornasier, and C. Sinan Güntürk. Iteratively re-weighted least squares minimization: Proof of faster than linear rate for sparse recovery. In 2008 42nd Annual Conference on Information Sciences and Systems (CISS 2008), volume ????, pages 26–29. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, March 2008. ISSN ????

Daubechies:2010:IRL

[DDFG10] Ingrid Daubechies, Ronald De-Vore, Massimo Fornasier, and C. Sinan Güntürk. Iteratively reweighted least squares minimization for sparse recovery. Communications on Pure and Applied Mathematics (New York), 63(1):1–38, January 2010. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Daubechies:2002:BEN

[DDGV02] Ingrid Daubechies, Ronald De-Vore, C. Sinan Güntürk, and Vinay A. Vaishampayan. Beta expansions: a new approach to digitally corrected A/D conversion. In 2002 IEEE International Symposium on Circuits and Systems. Proceedings, volume 2, pages II–784–II–787. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISSN ???? IEEE Catalog Number 02CH37353.

Daubechies:2006:DCI

[DDGV06] Ingrid Daubechies, Ronald A. DeVore, C. Sinan Güntürk, and Vinay A. Vaishampayan. A/D conversion with imperfect quantizers. IEEE Transactions on Information Theory, 52(3): 874–885, March 2006. CO-DEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic). IEEE International Symposium on Circuits and Systems, Phoenix, AZ, May 26–29, 2002.

Daubechies:2005:DSA

[DDK05] Ingrid Daubechies, Konstantinos Drakakis, and Tanya Khovanova. A detailed study of the attachment strategies of new autonomous systems in the AS connectivity graph. Internet Mathematics, 2(2):185– 246, 2005. ISSN 1542-7951 (print), 1944-9488 (electronic).

URL http://projecteuclid.
org/euclid.im/1137446621.

Daubechies:2007:APG

[DFL07] Ingrid Daubechies, Massimo Fornasier, and Ignace Loris.

Accelerated projected gradient method for linear inverse problems with sparsity constraints. arXiv.org, ??(??):1–24, June 28, 2007. CODEN ???? ISSN 2331-8422. URL https://

arxiv.org/abs/0706.4297.

Daubechies:2008:APG

[DFL08] Ingrid Daubechies, Massimo Fornasier, and Ignace Loris. Accelerated projected gradient method for linear inverse problems with sparsity constraints. Journal of Fourier Analysis and Applications, 14(5–6):764– 792, December 2008. ISSN 1069-5869 (print), 1531-5851 (electronic).

Daubechies:1980:ITR

[DG80] Ingrid Daubechies and Alex Grossmann. An integral transform related to quantization. Journal of Mathematical Physics, 21(8):2080–2090, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/
resource/1/jmapaq/v21/i8/
p2080_s1.

Daubechies:1988:FBS

[DG88] Ingrid Daubechies and Alex Grossmann. Frames in the Bargmann space of entire functions. Communications on Pure and Applied Mathematics (New York), 41(2):151–164, March 1988. CODEN CPA-MAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Daubechies:1999:HAW

[DG99] Ingrid C. Daubechies and Anna C. Gilbert. Harmonic analysis, wavelets and applications. In Hyperbolic Equations and Frequency Interactions (Park City, UT, 1995), volume 5 of IAS/Park City Math. Ser., pages 159–226. American Mathematical Society, Providence, RI, USA, 1999. ISBN 0-8218-0592-4 (hardcover).

Daubechies:1986:PNE

[DGM86] Ingrid Daubechies, Alex Grossmann, and Yves Meyer. Painless nonorthogonal expansions. Journal of Mathematical Physics, 27(5):1271–1283, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

                                        URL http://jmp.aip.org/
                                        resource/1/jmapaq/v27/i5/

p1271_s1. Reprinted in [DGM06].

Daubechies:2006:PNE

[DGM06] Ingrid Daubechies, Alex Grossmann, and Yves Meyer. Painless nonorthogonal expansions. In Heil and Walnut [HW06], pages 372–384. ISBN 0-691-11453-6 (hardcover), 0-691-12705-0 (paperback), 1-4008-2726-4. Reprint of [DGM86].

Daubechies:1983:ITR

[DGR83] Ingrid Daubechies, Alex Grossmann, and Jean Reignier. An integral transform related to quantization. II. Some mathematical properties. Journal of Mathematical Physics, 24(2): 239–254, February 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/
jmapaq/v24/i2/p239_s1.

Daubechies:1999:RIS

[DGS99] Ingrid Daubechies, I. Guskov, and W. Sweldens. Regularity of irregular subdivision. Constructive Approximation, 15(3): 381–426, 1999. ISSN 0176-4276 (print), 1432-0940 (electronic).

Daubechies:2001:CIS

[DGS01] Ingrid Daubechies, Igor Guskov, and Wim Sweldens. Commutation for irregular subdivision. Constructive Approximation, 17(4):479–514, 2001. ISSN 0176-4276 (print), 1432-0940 (electronic).

Daubechies:1999:WIP

[DGSS99] Ingrid Daubechies, Igor Guskov, Peter Schröder, and Wim Sweldens. Wavelets on irregular point sets. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 357 (1760):2397–2413, September 15, 1999. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

                                        URL http://www.jstor.org/
                                        stable/55170.

Daubechies:2008:GRE

[DGWY08] Ingrid Daubechies, C. S. Güntürk, Y. Wang, and Ö. Yılmaz. The golden ratio encoder. arXiv.org, ?? (??):1–24, September 7, 2008. CODEN ???? ISSN 2331-

                                        8422. URL https://arxiv.
                                        org/abs/0809.1257.

Daubechies:2010:GRE

[DGWY10] Ingrid Daubechies, Sinan Güntürk, Yang Wang, and Özgür Yılmaz. The golden ratio encoder. IEEE Transactions on Information Theory, 56(10):5097–5110, October 2010. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Daubechies:1994:DTR

[DH94] Ingrid Daubechies and Y. Huang. A decay theorem for refinable functions. Applied Mathematics Letters, 7(4):1–4, July 1994. CODEN AMLEEL. ISSN 0893-9659 (print), 1873-5452 (electronic).

Daubechies:1995:HDT

[DH95] Ingrid Daubechies and Ying Huang. How does truncation of the mask affect a refinable function? Constructive Approximation, 11(3):365–380, 1995. ISSN 0176-4276 (print), 1432-0940 (electronic).

Daubechies:2002:CDF

[DH02] Ingrid Daubechies and Bin Han. The canonical dual frame of a wavelet frame. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 12(3):269–285, May 2002. ISSN 1063-5203 (print), 1096-603x (electronic).

Daubechies:2004:PDW

[DH04] Ingrid Daubechies and Bin Han. Pairs of dual wavelet frames from any two refinable functions. Constructive Approximation, 20(3):325–352, 2004. ISSN 0176-4276 (print), 1432-0940 (electronic).

Daubechies:2003:FMB

[DHRS03] Ingrid Daubechies, Bin Han, Amos Ron, and Zuowei W. Shen. Framelets: MRAbased constructions of wavelet frames. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 14(1):1–46, January 2003. ISSN 1063-5203 (print), 1096-603x (electronic).

Daubechies:1993:TTL

[DJ93] Ingrid Daubechies and Ajem J. E. M. Janssen. Two theorems on lattice expansions. IEEE Transactions on Information Theory, 39(1):3–6, January 1993. CODEN IETTAW.

ISSN 0018-9448 (print), 1557-9654 (electronic).

Dominguez-Jimenez:2011:SEP

[DJF11] María Elena Domínguez-Jiménez and Paulo J. S. G. Ferreira. Some extremal properties of Daubechies filters and other orthonormal filters. Signal Processing, 91(1):85–89, 2011. CO-DEN SPRODR. ISSN 0165-1684 (print), 1872-7557 (electronic).

Daubechies:1991:ESW

[DJJ91a] Ingrid Daubechies, Stéphane Jaffard, and Jean-Lin Journé. Erratum: “A simple Wilson orthonormal basis with exponential decay”. SIAM Journal on Mathematical Analysis, 22(3):878, May 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [DJJ91b].

Daubechies:1991:SWO

[DJJ91b] Ingrid Daubechies, Stéphane Jaffard, and Jean-Lin Journé. A simple Wilson orthonormal basis with exponential decay. SIAM Journal on Mathematical Analysis, 22(2):554– 573, March 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [DJJ91a].

Daubechies:1982:CMP

[DK82] Ingrid Daubechies and John R. Klauder. Constructing measures for path integrals. Journal of Mathematical Physics, 23(10):1806–1822, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/
resource/1/jmapaq/v23/i10/
p1806_s1.

Daubechies:1983:MMQ

[DK83] Ingrid Daubechies and John R.

Klauder. Measures for more quadratic path integrals. Letters in Mathematical Physics, 7(3):229–234, 1983. CODEN LMPHDY. ISSN 0377-9017 (print), 1573-0530 (electronic).

Daubechies:1985:QMP

[DK85] Ingrid Daubechies and John R.

Klauder. Quantum-mechanical path integrals with Wiener measure for all polynomial Hamiltonians. II. Journal of Mathematical Physics, 26(9): 2239–2256, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/
jmapaq/v26/i9/p2239_s1.

Daubechies:1986:TMR

[DK86] Ingrid Daubechies and John R.

Klauder. True measures for real time path integrals. In Path Integrals from meV to MeV (Bielefeld, 1985), Bielefeld Encount. Phys. Math., VII, pages 425–432. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1986.

Drouiche:1999:DMW

[DKB99] Karim Drouiche, Djalil Kateb, and Daniel Boichu. Daubechies– Matzinger wavelets and Lorentz polynomials. Advances in Computational Mathematics, 10(3–4):239–260, 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

Dyn:2012:RGD

[DKLR12] N. Dyn, O. Kounchev, D. Levin, and H. Render. Regularity of generalized Daubechies wavelets reproducing exponential polynomials. arXiv.org, ?? (??):1–27, October 30, 2012. CODEN ???? ISSN 2331-

                                        8422. URL https://arxiv.
                                        org/abs/1210.7938.

Dyn:2014:RGD

[DKLR14] Nira Dyn, Ognyan Kounchev, David Levin, and Hermann Render. Regularity of generalized Daubechies wavelets reproducing exponential polynomials with real-valued parameters. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 37(2):288–306, 2014. ISSN 1063-5203 (print), 1096-603x (electronic).

Daubechies:1987:WMP

[DKP87] Ingrid Daubechies, John R. Klauder, and Thierry Paul. Wiener measures for path integrals with affine kinematic variables. Journal of Mathematical Physics, 28(1):85–102, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/
resource/1/jmapaq/v28/i1/
p85_s1.

Daubechies:2012:AHA

[DKRS12] Ingrid Daubechies, Gitta Kutyniok, Holger Rauhut, and Thomas Strohmer. Applied harmonic analysis and sparse approximation. Abstracts from the workshop held June 10– 16, 2012. Oberwolfach Reports, 9(2):1759–1843, 2012. ISSN 1660-8933 (print), 1660-8941 (electronic).

Daubechies:2015:AHA

[DKRS15] Ingrid Daubechies, Gitta Kutyniok, Holger Rauhut, and Thomas Strohmer. Applied harmonic analysis and sparse approximation. Abstracts from the workshop held August 16– 22, 2015. Oberwolfach Reports, 12(3):2189–2263, 2015. ISSN 1660-8933 (print), 1660-8941 (electronic).

Daubechies:2018:AHA

[DKRS18] Ingrid Daubechies, Gitta Kutyniok, Holger Rauhut, and Thomas Strohmer. Applied harmonic analysis and data processing. Oberwolfach Reports, 15(1):723–792, 2018. ISSN 1660-8933,1660-8941. Abstracts from the workshop held March 25–31, 2018, Organized by Ingrid Daubechies, Gitta Kutyniok, Holger Rauhut and Thomas Strohmer.

Daubechies:2021:AHA

[DKRS21] Ingrid Daubechies, Gitta Kutyniok, Holger Rauhut, and Thomas Strohmer. Applied harmonic analysis and data science (hybrid meeting). Oberwolfach Reports, 18 (4):3007–3066, 2021. ISSN 1660-8933,1660-8941. Abstracts from the workshop held November 28–December 4, 2021, Organized by Ingrid Daubechies, Gitta Kutyniok, Holger Rauhut and Thomas Strohmer.

Daubechies:1983:OER

[DL83] Ingrid Daubechies and Elliott H. Lieb. One electron relativistic molecules with Coulomb interaction. Communications in Mathematical Physics, 90(4):497–510, 1983. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916 (electronic). URL http://

                                        projecteuclid.org/euclid.
                                        cmp/1103940413.

Daubechies:1984:RMC

[DL84] Ingrid Daubechies and Elliott H. Lieb. Relativistic molecules with Coulomb interaction. In Differential Equations (Birmingham, Ala., 1983), volume 92 of North-Holland Math. Stud., pages 143–148. North-Holland, Amsterdam, The Netherlands, 1984.

Daubechies:1991:TSD

[DL91] Ingrid Daubechies and Jeffrey C. Lagarias. Two-scale difference equations. I. Existence and global regularity of solutions. SIAM Journal on Mathematical Analysis, 22(5):1388– 1410, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Daubechies:1992:SMA

[DL92a] Ingrid Daubechies and Jeffrey C. Lagarias. Sets of matrices all infinite products of which converge. Linear Algebra and its Applications, 161 (1):227–263, January 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See corrigendum [DL01b].

Daubechies:1992:TSD

[DL92b] Ingrid Daubechies and Jeffrey C. Lagarias. Two-scale difference equations. II. Local regularity, infinite products of matrices and fractals. SIAM Journal on Mathematical Analysis, 23(4):1031–1079, July 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Daubechies:1994:TFMa

[DL94a] Ingrid Daubechies and Jeffrey C. Lagarias. On the thermodynamic formalism for multifractal functions. Reviews in Mathematical Physics, 6(5A): 1033–1070, 1994. CODEN RMPHEX. ISSN 0129-055X (print), 1793-6659 (electronic). Special issue dedicated to Elliott H. Lieb.

Daubechies:1994:TFMb

[DL94b] Ingrid Daubechies and Jeffrey C. Lagarias. On the thermodynamic formalism for multifractal functions. In The State of Matter (Copenhagen, 1992), volume 20 of Adv. Ser. Math. Phys., pages 213–264. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1994.

Dumont:1996:RPE

[DL96] S. Dumont and F. Lebon. Representation of plane elastostatics operators in Daubechies wavelets. Computers and Structures, 60(4):561–569, June 1996. CODEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Daubechies:2001:SMA

[DL01a] Ingrid Daubechies and J. C. Lagarias. Sets of matrices all infinite products of which converge (vol 161, pg 227, 1992). Linear Algebra and its Applications, 327(1–3):69–83, April 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Daubechies:2001:CAS

[DL01b] Ingrid Daubechies and Jeffrey C. Lagarias. Corrigendum/addendum to: “Sets of matrices all infinite products of which converge” [Linear Algebra Appl. 161 (1992), 227– 263; MR1142737 (93f:15006)]. Linear Algebra and its Applications, 327(1–3):69–83, April 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/30/
19/152/25/31/abstract.html;
 http://www.elsevier.nl/
gej-ng/10/30/19/152/25/31/
article.pdf.

Daubechies:2005:OER

[DL05] Ingrid Daubechies and Elliott H. Lieb. One-electron relativistic molecules with Coulomb interaction. In Lieb and Thirring [LT05], pages 471–484. ISBN 3-540-22212-X (hardcover), 3-540-27056-6. LCCN QC173.4.T48 L54 2005.

URL http://link.springer.
com/10.1007/b138553.

Daubechies:1995:GTF

[DLL95] Ingrid Daubechies, H. J. Landau, and Zeph Landau. Gabor time-frequency lattices and the Wexler–Raz identity. Journal of Fourier Analysis and Applications, 1(4):437–478, 1995. ISSN 1069-5869 (print), 1531-5851 (electronic).

Daubechies:2007:LER

[DLM^{+}07] Ingrid Daubechies, Robert Lazarsfeld, John Morgan, Andrei Okounkov, and Terence Tao. Letter to the Editor: Reply to Davey, Henriksen, Markovic and Pratt. Notices of the American Mathematical Society, 54(6):694, June/July 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL http://www.ams.org/
                                        notices/200706/200706-full-
                                        issue.pdf.

Daubechies:2009:SWT

[DLW09] Ingrid Daubechies, Jianfeng Lu, and Hau-Tieng Wu. Synchrosqueezed wavelet transforms: a tool for empirical mode decomposition. arXiv.org, ??(??):1–23, December 12, 2009. CODEN ???? ISSN 2331-8422. URL https:

                                        //arxiv.org/abs/0912.2437.

Daubechies:2011:SWT

[DLW11] Ingrid Daubechies, Jianfeng Lu, and Hau-Tieng Wu. Synchrosqueezed wavelet transforms: an empirical mode decomposition-like tool. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 30(2):243–261, March 2011. ISSN 1063-5203 (print), 1096-603x (electronic).

Daubechies:1996:NSC

[DM96] Ingrid Daubechies and Stéphane Maes. A nonlinear squeezing of the continuous wavelet transform based on auditory nerve models. In Wavelets in Medicine and Biology, pages 527–546. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1996. ISBN 0-8493-9483-X (hardcover).

Daubechies:1999:AFA

[DM99] Ingrid Daubechies and Stéphane Maes. An application of a formula of Alberto Calderón to speaker identification. In Harmonic Analysis and Partial Differential Equations. Essays in Honor of Alberto P. Calderón’s 75th Birthday. Proceedings of a Conference, University of Chicago, IL, USA, February 1996, Chicago Lectures in Math., pages 163–181. University of Chicago Press, Chicago, IL, USA and London, UK, 1999. ISBN 0-226-10456-7 (hardcover).

Deligiannis:2016:XRIb

[DMC^{+}16a] N. Deligiannis, J. F. C. Mota, B. Cornelis, M. R. D. Rodrigues, and Ingrid Daubechies. X-ray image separation via coupled dictionary learning. In 2016 IEEE International Conference on Image Processing (ICIP), pages 3533–3537. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 2016. ISBN 1-4673-9961-2. ISSN 1522-4880 (print), 2381-8549 (electronic).

Deligiannis:2016:MMD

[DMC^{+}16b] Nikos Deligiannis, Joao F. C. Mota, Bruno Cornelis, Miguel R. D. Rodrigues, and Ingrid Daubechies. Multi-modal dictionary learning for image separation with application in art investigation. arXiv.org, ?? (??):1–13, July 14, 2016. CO-DEN ???? ISSN 2331-

                                        8422. URL https://arxiv.
                                        org/abs/1607.04147.

Deligiannis:2016:XRIa

[DMC^{+}16c] Nikos Deligiannis, João F. C. Mota, Bruno Cornelis, Miguel R. D. Rodrigues, and Ingrid Daubechies. X-ray image separation via coupled dictionary learning. arXiv.org, ??(??):1– 6, May 20, 2016. CODEN ???? ISSN 2331-8422. URL https:/

                                        /arxiv.org/abs/1605.06474.

Deligiannis:2017:MMD

[DMC^{+}17] Nikos Deligiannis, João F. C. Mota, Bruno Cornelis, Miguel R. D. Rodrigues, and Ingrid Daubechies. Multi-modal dictionary learning for image separation with application in art investigation. IEEE Transactions on Image Processing, 26(2):751–764, February 2017. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic). 23rd IEEE International Conference on Image Processing (ICIP), Phoenix, AZ, SEP 25-28, 2016.

Diaz:2009:DWB

[DMV09] Lilliam Alvarez Díaz, María T.

Martín, and Victoria Vampa. Daubechies wavelet beam and plate finite elements. Finite Elements in Analysis and Design, 45(3):200–209, 2009. CO-DEN FEADEU. ISSN 0168-874X (print), 1872-6925 (electronic).

Daubechies:1992:SIW

[DMW92] Ingrid Daubechies, S. Mallat, and A. S. Willsky. Special issue on wavelet transforms and multiresolution signal analysis — introduction. IEEE Transactions on Information Theory, 38(2, 2):529–531, March 1992. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Doroslovacki:1994:DAD

[Dor94] M. Doroslovacki. Diminishing the asymmetry of Daubechies’ wavelets. In Proceedings of IEEE–SP International Symposium on Time-Frequency and Time-Scale Analysis. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994.

Daubechies:1987:WA

[DP87] Ingrid Daubechies and Thierry Paul. Wavelets and applications. In VIIIth International Congress on Mathematical Physics (Marseille, 1986), pages 675–686. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1987.

Daubechies:1988:TFLb

[DP88] Ingrid Daubechies and Thierry Paul. Time–frequency localisation operators — a geometric phase space approach. II. The use of dilations. Inverse Problems, 4(3):661–680, August 1988. CODEN INPEEY. ISSN 0266-5611 (print), 1361-6420 (electronic). URL http://

                                        stacks.iop.org/0266-5611/
                                        4/661.

Daubechies:2002:AGT

[DP02] Ingrid Daubechies and Fabrice Planchon. Adaptive Gabor transforms. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 13(1):1–21, July 2002. ISSN 1063-5203 (print), 1096-603x (electronic).

Daubechies:2004:NMA

[DRS04] Ingrid Daubechies, Olof Runborg, and Wim Sweldens. Normal multiresolution approximation of curves. Constructive Approximation, 20(3):399– 463, May 2004. ISSN 0176-4276 (print), 1432-0940 (electronic).

Daubechies:2009:ICA

[DRT^{+}09] Ingrid Daubechies, E. Roussos, S. Takerkart, M. Benharrosh, C. Golden, K. D’Ardenne, W. Richter, J. D. Cohen, and J. Haxby. Independent component analysis for brain fMRI does not select for independence. Proceedings of the National Academy of Sciences of the United States of America, 106(26):10415–10422, June 30, 2009. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

URL http://www.jstor.org/
stable/40483565.

Daubechies:2007:SSM

[DRZ07] Ingrid Daubechies, Olof Runborg, and Jing Zou. A sparse spectral method for homogenization multiscale problems. Multiscale Modeling and Simulation, 6(3):711–740, ???? 2007. CODEN MMSUBT. ISSN 1540-3459 (print), 1540-3467 (electronic).

Daubechies:1998:FWT

[DS98] Ingrid Daubechies and Wim Sweldens. Factoring wavelet transforms into lifting steps. Journal of Fourier Analysis and Applications, 4(3):247– 269, 1998. ISSN 1069-5869 (print), 1531-5851 (electronic).

Daubechies:2000:FWT

[DS00] Ingrid Daubechies and Wim Sweldens. Factoring wavelet transforms into lifting steps. In Roland Klees and R. H. N. Haagmans, editors, Wavelets in the Geosciences. Collection of the Lecture Notes of the School of Wavelets in the Geosciences, Delft, Netherlands, October 4– 9, 1998, volume 90 of Lecture Notes in Earth Sciences, pages 131–157. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-66951-5 (paperback), 3-540-46590-1 (electronic). ISSN 0930-0317. LCCN QE33.2.W38 W38 2000.

Daubechies:2015:DADa

[DS15a] Ingrid Daubechies and Rayan Saab. A deterministic analysis of decimation for sigma– delta quantization of bandlimited functions. arXiv.org, ?? (??):1–9, May 30, 2015. CO-DEN ???? ISSN 2331-

                                        8422. URL https://arxiv.
                                        org/abs/1506.00179.

Daubechies:2015:DADb

[DS15b] Ingrid Daubechies and Rayan Saab. A deterministic analysis of decimation for sigma– delta quantization of bandlimited functions. IEEE Signal Processing Letters, 22(11): 2093–2096, November 2015. CODEN ISPLEM. ISSN 1070-9908 (print), 1558-2361 (electronic).

Daubechies:2004:WBI

[DT04] Ingrid Daubechies and Gerd Teschke. Wavelet-based image decomposition by variational functionals. Proceedings of the SPIE — The International Society for Optical Engineering, 5266:94–105, February 27, 2004. CODEN PSISDG.

ISSN 0277-786X (print), 1996-756X (electronic).

Daubechies:2005:VIR

[DT05] Ingrid Daubechies and G. Teschke.

Variational image restoration by means of wavelets: Simultaneous decomposition, deblurring, and denoising. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 19(1):1–16, July 2005. ISSN 1063-5203 (print), 1096-603x (electronic).

Davari:2014:ADW

[DT14] A. Davari and M. Torabi.

Application of Daubechies wavelets for solving Kuramoto– Sivashinsky type equations. Caspian Journal of Mathematical Sciences, 3(1):57–66, 2014. ISSN 1735-0611.

Daubechies:2007:ISL

[DTV07] Ingrid Daubechies, Gerd Teschke, and Luminita Vese. Iteratively solving linear inverse problems under general convex constraints. Inverse Problems and Imaging, 1(1):29–46, February 2007. ISSN 1930-8337 (print), 1930-8345 (electronic).

Daubechies:2008:SIC

[DTV08] Ingrid Daubechies, Gerd Teschke, and Luminita Vese. On some iterative concepts for image restoration. Advances in Imaging and Electron Physics, 150:

1–51, 2008. ISBN 0-12-374217-X. ISSN 1076-5670.

Du:2001:WTD

[Du01] Jingde Du. Weyl transforms and Daubechies operators: Products, traces, eigenvalues and eigenfunctions. ProQuest LLC, Ann Arbor, MI, USA, 2001. ISBN 0-612-70568-4. 87

                                        pp. URL http://gateway.
                                        proquest.com/openurl?url_
                                        ver=Z39.88-2004&rft_val_
                                        fmt=info:ofi/fmt:kev:mtx:
                                        dissertation&res_dat=xri:
                                        pqdiss&rft_dat=xri:pqdiss:

NQ70568. Thesis (Ph.D.)–York University (Canada).

Staff:2014:DFF

[Duk14] Duke Today Staff. Daubechies on the first female winner of the Fields Math Medal. Web story., August 14, 2014.

Staff:2016:SFA

[Duk16] Duke Today Staff. Simons Foundation awards Duke’s Ingrid Daubechies $1.5M grant. Web story., August 1, 2016.

                                        URL https://today.duke.
                                        edu/2016/08/simonsdaubechies.

Durand:1993:CCA

[Dur93] Sylvain Durand. Convergence of cascade algorithms introduced by I. Daubechies. Numerical Algorithms, 4(3):307– 322, October 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Durand:1996:EVC

[Dur96] Sylvain Durand. Étude de la vitesse de convergence de l’algorithme en cascade dans la construction des ondelettes d’Ingrid Daubechies. (French) [Study of the speed of convergence of the cascading algorithm in the construction of the wavelets of Ingrid Daubechies]. Revista Matemática Iberoamericana, 12(2):277–297, 1996. ISSN 0213-2230 (print), 2235-0616 (electronic).

Donoho:1998:DCH

[DVDD98] David L. Donoho, Martin Vetterli, R. A. DeVore, and Ingrid Daubechies. Data compression and harmonic analysis. IEEE Transactions on Information Theory, 44(6):2435– 2476, October 1998. CO-DEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic). Information theory: 1948–1998.

Du:2000:GFD

[DW00a] Jingde Du and M. W. Wong.

Gaussian functions and Daubechies operators. Integral Equations and Operator Theory, 38(1): 1–8, 2000. ISSN 0378-620X (print), 1420-8989 (electronic).

Du:2000:GSD

[DW00b] Jingde Du and M. W. Wong.

Gaussian series and Daubechies operators. Applicable Analysis, 76(1–2):83–91, October 2000. CODEN APANCC. ISSN 0003-6811 (print), 1563-504X (electronic).

Du:2001:EED

[DW01] Jingde Du and M. W. Wong. Eigenvalues and eigenfunctions of Daubechies operators with radial symbols. Far East Journal of Mathematical Sciences (FJMS), Spec. Vol.(Special Volume, Part III):257–269, 2001. ISSN 0972-0871.

Daubechies:2015:CCF

[DWtW15] Ingrid Daubechies, Yi Wang, and Hau tieng Wu. ConceFT: Concentration of frequency and time via a multitapered synchrosqueezed transform. arXiv.org, ??(??):1–44, July 20, 2015. CODEN ???? ISSN 2331-8422. URL https:/

                                        /arxiv.org/abs/1507.05366.

Daubechies:2016:CCF

[DWtW16] Ingrid Daubechies, Yi (Grace) Wang, and Hau tieng Wu. ConceFT: concentration of frequency and time via a multitapered synchrosqueezed transform. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 374(2065):20150193, April 13, 2016. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Daubechies:2006:RPA

[DY06] Ingrid Daubechies and Özgür Yılmaz. Robust and practical analog-to-digital conversion with exponential precision. IEEE Transactions on Information Theory, 52(8):3533– 3545, August 2006. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Elgendi:2014:CAB

[EBJ^{+}14] Mohamed Elgendi, Prashant Bobhate, Shreepal Jain, Jennifer Rutledge, Yashu Coe, Roger Zemp, Dale Schuurmans, and Ian Adatia. Detection of the first and second heart sounds in children using Daubechies wavelets. Pulmonary Circulation, 4(3):527, September 2014. ISSN 2045-8932 (print), 2045-8940 (elec-

tronic). URL http://www.
jstor.org/stable/10.1086/

677368. Abstracts of the 7th International Conference on Neonatal and Childhood Pulmonary Vascular Disease, editor Jeffrey Fineman.

Ephremidze:2011:ADW

[EGL11] Lasha Ephremidze, Aleksander Gamkrelidze, and Edem Lagvilava. An approximation of Daubechies wavelet matrices by perfect reconstruction filter banks with rational coefficients. arXiv.org, ??(??):1–10, June 27, 2011. CODEN ???? ISSN 2331-8422. URL https:

//arxiv.org/abs/1106.5439.

Ephremidze:2013:ADW

[EGL13] L. Ephremidze, A. Gamkrelidze, and E. Lagvilava. An approximation of Daubechies wavelet matrices by perfect reconstruction filter banks with rational coefficients. Advances in Computational Mathematics, 38(1):147–158, 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

Fodor:2017:CRX

[FCY^{+}17] Gábor Fodor, Bruno Cornelis, Rujie Yin, Ann Dooms, and Ingrid Daubechies. Cradle removal in X-ray images of panel paintings. IPOL Journal. Image Processing Online, 7:23–42, 2017. ISSN 2105-1232.

Fulwood:2017:ASM

[FGDB17] Ethan Lucas. Fulwood, Tingran Gao, Ingrid Daubechies, and Doug M. Boyer. Automatic segmentation of morphological structure into biologically corresponding features: implications for systematics and ecomorphology. American Journal of Physical Anthropology, 162(64):188, April 2017. CO-DEN AJPNA9. ISSN 0002-9483 (print), 1096-8644 (electronic). 86th Annual Meeting of the American Association of Physical Anthropologists (AAPA), New Orleans, LA, April 19–22, 2017.

Finek:2004:DWI

[Fin04a]  Václav Finěk.  Daubechies

wavelets on intervals with application to BVPs. Applications of Mathematics, 49(5): 465–481, 2004. CODEN APM-TEO. ISSN 0862-7940 (print), 1572-9109 (electronic).

Finvek:2004:DWI

[Fin04b]  Václav Finěk.  Daubechies

wavelets on intervals with application to BVPs. Applications of Mathematics, Praha, 49(5):465–481, 2004. ISSN 0862-7940 (print), 1572-9109 (electronic).

Fornasier:2021:RRE

[FVD21] Massimo Fornasier, Jan Vybíral, and Ingrid Daubechies. Robust and resource efficient identification of shallow neural networks by fewest samples. Inf. Inference, 10(2):625–695, 2021. ISSN 2049-8764,2049-8772.

Gao:2014:CAB

[Gao14] Jing Gao. Comparison analysis based on the cubic spline wavelet and Daubechies wavelet of harmonic balance method. Abstract and Applied Analysis, pages Art. ID 634974, 6, 2014. ISSN 1085-3375 (print), 1687-0409 (electronic).

Gimbutas:1995:DCSb

[GB95a] Ž. Gimbutas and A. Bastys.

Daubechies compactly supported wavelets with minimal Heisenberg boxes. Lithuanian Mathematical Journal, 35(4): 343–362, 1995. ISSN 0363-1672 (print), 1573-8825 (electronic).

Gimbutas:1995:DCSa

[GB95b] Ž. Gimbutas and A. Bastys. I. Daubechies compactly supported wavelets with minimal Heisenberg boxes. Lietuvos Matematikos Rinkinys, 35(4): 432–455, 1995. ISSN 0132-2818.

Gowers:2008:PCM

[GBGL08] Timothy Gowers, June Barrow-Green, and Imre Leader, editors. The Princeton Companion to Mathematics. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-11880-9 (hardcover). xx + 1034 pp. LCCN QA11.2 .P745 2008. URL http:

//www.loc.gov/catdir/toc/
ecip0818/2008020450.html.

Govaerts:2009:KDCa

[GBM09a] Jan Govaerts, Calvin Matondo Bwayi, and Olivier Mattelaer. The Klauder–Daubechies construction of the phase space path integral and the harmonic oscillator. arXiv.org, ??(??): 1–18, August 6, 2009. CO-DEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/0908.0805.

Govaerts:2009:KDCb

[GBM09b] Jan Govaerts, Calvin Matondo Bwayi, and Olivier Mattelaer. The Klauder–Daubechies construction of the phase space path integral and the harmonic oscillator. Journal of Physics A: Mathematical and Theoretical, 42(44):445304, October 2009. CODEN JPAMB5. ISSN 1751-8113 (print), 1751-8121 (electronic).

Gupta:2012:MUI

[GEV12] Rishu Gupta, I. Elamvazuthi, and P. Vasant. Musculoskeletal ultrasound image denoising using Daubechies wavelets. AIP Conference Proceedings, 1499:263–??, 2012. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Gao:2018:GPL

[GKBD18] Tingran Gao, Shahar Z. Kovalsky, Doug M. Boyer, and Ingrid Daubechies. Gaussian process landmarking on manifolds. arXiv.org, ??(??):1–41, February 9, 2018. CODEN ???? ISSN 2331-8422. URL https:/

/arxiv.org/abs/1802.03479.

Gao:2019:GPLb

[GKBD19] Tingran Gao, Shahar Z. Kovalsky, Doug M. Boyer, and Ingrid Daubechies. Gaussian process landmarking for threedimensional geometric morphometrics. SIAM Journal on Mathematics of Data Science, 1 (1):237–267, ???? 2019. CO-DEN SJMDAQ. ISSN 2577-0187.

Gao:2019:GPLa

[GKD19] Tingran Gao, Shahar Z. Kovalsky, and Ingrid Daubechies. Gaussian process landmarking on manifolds. SIAM Journal on Mathematics of Data Science, 1(1):208–236, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Gagnon:1994:SDW

[GL94] L. Gagnon and J.-M. M. Lina. Symmetric Daubechies’ wavelets and numerical solution of NLS equations. Journal of Physics A (Mathematical and General), 27 (24):8207–8230, 1994. CO-DEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (elec-

tronic). URL http://stacks.
iop.org/0305-4470/27/8207.

Gagnon:1994:ACD

[GLG94] Langis Gagnon, J. M. Lina, and B. Goulard. Application of complex Daubechies’ wavelets to numerical simulation of a nonlinear signal propagation model. arXiv.org, ?? (??):1–4, March 31, 1994. CO-DEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/comp-gas/9403006.

Genovese:2008:DWBa

[GNG^{+}08a] Luigi Genovese, Alexey Neelov, Stefan Goedecker, Thierry Deutsch, Seyed Alireza Ghasemi, Alexander Willand, Damien Caliste, Oded Zilberberg, Mark Rayson, Anders Bergman, and Reinhold Schneider. Daubechies wavelets as a basis set for density functional pseudopotential calculations. arXiv.org, ??(??):1–15, April 16, 2008. CODEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/0804.2583.  Pub-

lished in J. Chem. Phys. 129, 014109 (2008).

Genovese:2008:DWBb

[GNG^{+}08b] Luigi Genovese, Alexey Neelov, Stefan Goedecker, Thierry Deutsch, Seyed Alireza Ghasemi, Alexander Willand, Damien Caliste, Oded Zilberberg, Mark Rayson, Anders Bergman, and Reinhold Schneider. Daubechies wavelets as a basis set for density functional pseudopotential calculations. Journal of Chemical Physics, 129(1):014109, July 2008. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Gregoor:2011:BVI

[Gre11] Ralph Gregoor. Burgemeester verrast internationaal bekende wiskundige Ingrid Daubechies krijgt straat in geboortedorp. (Dutch) [Mayor surprises internationally known mathematician: Ingrid Daubechies gets a street in birth village]. Het Nieuwsblad, ??(??): ??, July 7, 2011. URL https:

//www.nieuwsblad.be/cnt/
ef3ce3u4.

Griffel:1995:BRT

[Gri95] D. H. Griffel. Book review:

Ten Lectures on Wavelets, by Ingrid Daubechies. Mathematical Gazette, 79(484):224– 227, March 1995. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

URL http://www.jstor.org/
stable/3620105.

Grunbaum:1992:BRF

[Grü92] F. Alberto Grünbaum. Book review: Functional analysis: Ten Lectures on Wavelets, by Ingrid Daubechies; An Introduction to Wavelets, by Charles K. Chui. Science, 257(5071):821–822, August 7, 1992. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.jstor.org/
stable/2877684.

Genovese:2011:DWH

[GVO^{+}11] Luigi Genovese, Brice Videau, Matthieu Ospici, Thierry Deutsch, Stefan Goedecker, and Jean-François Méhaut. Daubechies wavelets for high performance electronic structure calculations: the BigDFT project. Comptes Rendus. Mécanique. Académie des Sciences, Paris, 339(2–3):149– 164, 2011. ISSN 1631-0721 (print), 1873-7234 (electronic).

Gao:2018:DAF

[GYD^{+}18] Tingran Gao, Gabriel S. Yapuncich, Ingrid Daubechies, Sayan Mukherjee, and Doug M. Boyer. Development and assessment of fully automated and globally transitive geometric morphometric methods, with application to a biological comparative dataset with high interspecific variation. The Anatomical Record — Advances in Integrative Anatomy and Evolutionary Biology, 301 (4):656–658, April 2018. ISSN 1932-8486 (print), 1932-8494 (electronic).

Hughes:2006:SAI

[HD06] Shannon M. Hughes and Ingrid Daubechies. Simpler alternatives to information theoretic similarity metrics for multimodal image alignment. In 2006 International Conference on Image Processing, pages 365–368. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 2006. ISSN 1522-4880 (print), 2381-8549 (electronic).

Huang:2016:ADA

[HDH16] Norden E. Huang, Ingrid Daubechies, and Thomas Y. Hou. Adaptive data analysis: theory and applications. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 374(2065):20150207, April 13, 2016. CODEN PTR-MAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Heil:1992:BRI

[Hei92] Christopher Heil. Book reviews: Ingrid Daubechies, Ten Lectures on Wavelets. Computers in Physics, 6(6): 697, November 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

                                        aip.scitation.org/doi/10.
                                        1063/1.4823127.

Heil:1993:BRT

[Hei93] Christopher Heil. Book review: Ten Lectures on Wavelets, by Ingrid Daubechies. SIAM Review, 35(4):666–669, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

                                        URL http://www.jstor.org/
                                        stable/2132406.

He:2015:NDA

[HN15] Tian-Xiao He and Tung Nguyen. A note on the Daubechies approach in the construction of spline type orthogonal scaling functions. arXiv.org, ??(??):1–14, July 11, 2015. CODEN ???? ISSN 2331-8422. URL https://

                                        arxiv.org/abs/1507.03057.

He:2017:UAC

[HN17] Tianxiao He and Tung Nguyen. A unified approach to construct a class of Daubechies orthogonal scaling functions. Journal of Mathematical Research with Applications, 37(1):29–39, 2017. ISSN 2095-2651.

Hope:2018:FFI

[Hop17] Alan Hope. Face of Flanders: Ingrid Daubechies. FlandersToday, ??(??):??, September

                                        25, 2017. URL http://www.
                                        flanderstoday.eu/current-
                                        affairs/face-flanders-ingrid-
                                        daubechies.

Hofer:2009:DWI

[HPH09] Vera Hofer, Jürgen Pilz, and Thorgeir S. Helgason. Daubechies wavelets for identification of rock variants from IR spectra. In Pilz [Pil09], pages 79–88. ISBN 3-540-33235-9 (hardcover), 1-281-92749-X, 3-540-33236-7 (ebook). LCCN QE33.2.S82

I58 2009. URL http://www.
springerlink.com/openurl.
asp?genre=book%26isbn=978-
3-540-33236-7.

Hasiewicz:2002:INL

[HŚ02a] Z. Hasiewicz and P. Śliwiński.

Identification of non-linear characteristics of a class of block-oriented non-linear systems via Daubechies waveletbased models. International Journal of Systems Science, 33(14):1121–1144, 2002. CO-DEN IJSYA9. ISSN 0020-7721 (print), 1464-5319 (electronic).

Hasiewicz:2002:INC

[HŚ02b] Z. Hasiewicz and P. Śliwiński.

Identification of nonlinear characteristics of a class of blockoriented nonlinear systems via Daubechies wavelet-based models. International Journal of Systems Science, 33 (14):1121–1144, 2002. CO-DEN IJSYA9. ISSN 0020-7721 (print), 1464-5319 (electronic).

Huang:1997:IRB

[HSZ97] Daren Huang, Qiyu Sun, and Zeyin Zhang. Integral representation of M-band filters of Daubechies type. Chinese Science Bulletin, 42(10):803–807, 1997. ISSN 1001-6538 (print), 1861-9541 (electronic).

Huylebrouck:2008:IMI

[Huy08]  Dirk Huylebrouck.  Inter-

view met Ingrid Daubechies: De typische reflex van de wiskundige. (Dutch) [Interview with Ingrid Daubechies: The typical reflex of the mathematician]. Nieuw Archief voor Wiskunde. Vijfde Serie, 9 (3):198–203, 2008. CODEN NAWIA7. ISSN 0028-9825.

Heil:2006:FPW

[HW06] Christopher Heil and David F. Walnut, editors. Fundamental Papers in Wavelet Theory. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-11453-6 (hardcover), 0-691-12705-0 (paperback), 1-4008-2726-4. xviii + 878 pp. With a foreword by Ingrid Daubechies. Introduction by John J. Benedetto.

Jameson:1996:DMD

[Jam96] Leland Jameson. The differentiation matrix for Daubechiesbased wavelets on an interval. SIAM Journal on Scientific Computing, 17(2):498– 516, 1996. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Jacques:2014:PSI

[JDB^{+}14] L. Jacques, C. De Vleeschouwer, Y. Boursier, P. Sudhakar, C. De Mol, A. Pizurica, S. Anthoine, P. Vandergheynst, P. Frossard, C. Bilen, S. Kitic, N. Bertin, R. Gribonval, N. Boumal, B. Mishra, P.-A. Absil, R. Sepulchre, S. Bundervoet, C. Schretter, A. Dooms, P. Schelkens, O. Chabiron, F. Malgouyres, J.-Y. Tourneret, N. Dobigeon, P. Chainais, C. Richard, B. Cornelis, I. Daubechies, D. Dunson, M. Dankova, P. Rajmic, K. Degraux, V. Cambareri, B. Geelen, G. Lafruit, G. Setti, J.-F. Determe, J. Louveaux, F. Horlin, A. Drémeau, P. Heas, C. Herzet, V. Duval, G. Peyré, A. Fawzi, M. Davies, N. Gillis, S. A. Vavasis, C. Soussen, L. Le Magoarou, J. Liang, J. Fadili, A. Liutkus, D. Martina, S. Gigan, L. Daudet, M. Maggioni, S. Minsker, N. Strawn, C. Mory, F. Ngole, J.-L. Starck, Ignace Loris, S. Vaiter, et al. Proceedings of the second international traveling workshop on interactions between sparse models and technology (iTWIST’14). arXiv.org, ?? (??):1–69, October 2, 2014. CODEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/1410.0719.

Johnson:2008:IPA

[JHB^{+}08] C. Richard Johnson, Jr., Ella Hendriks, Igor J. Berezhnoy, Eugene Brevdo, Shannon M. Hughes, Ingrid Daubechies, Jia Li, Eric Postma, and James Z. Wang. Image processing for artist identification: Computerized analysis of Vincent van Gogh’s painting brushstrokes. IEEE Signal Processing Magazine, 25(4):37–48, July 2008. CODEN ISPRE6. ISSN 1053-5888 (print), 1558-0792 (electronic).

Jafarpour:2009:SAP

[JPB^{+}09] S. Jafarpour, G. Polatkan, E. Brevdo, Shannon Hughes, A. Brasoveanu, and Ingrid Daubechies. Stylistic analysis of paintings using wavelets and machine learning. In 2009 17th European Signal Processing Conference, volume ????, pages 1220–1224. ????, ????, August 2009. ISSN ????

Jiankang:1998:TSS

[JZL98] Zhang Jiankang, Bao Zheng, and Jiao Licheng. Transformation of Shannon’s sampling points into Daubechies’ wavelet sampling points. In ICSP ’98. 1998 Fourth International Conference on Signal Processing, pages 305–308. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. IEEE catalog number 98TH8344.

Khalid:2011:WAH

[KAB11] Azra Khalid, Uzma Afsheen, and Saad Umer Baig. Wavelet analysis and high quality JPEG2000 compression using Daubechies wavelet. In International Conference on Graphic and Image Processing (ICGIP 2011). SPIE Optical Engineering Press, Bellingham, WA, USA, October 2011.

Kaiser:2010:DOW

[Kai10] Gerald Kaiser. Daubechies’ orthonormal wavelet bases. In A Friendly Guide to Wavelets, pages 176–199. Birkhäuser Boston Inc., Cambridge, MA, USA, October 2010.

Karam:2003:CDW

[Kar03] Jalal R. Karam. Connecting Daubechies wavelet db8 with the Kakeya–Eneström theorem. International Journal of Applied Mathematics, 14(2): 175–186, 2003. ISSN 1311-1728 (print), 1314-8060 (electronic).

Karam:2004:CDW

[Kar04] Jalal R. Karam. Connecting Daubechies wavelet db8 with the Kakeya–Enestrom theorem. International Journal of Applied Mathematics, 14(2): 175–186, 2004. ISSN 1311-1728 (print), 1314-8060 (electronic).

Karam:2007:RDP

[Kar07] Jalal Karam. On the roots of Daubechies polynomials. International Journal of Applied Mathematics, 20(8):1067–1074, 2007. ISSN 1311-1728 (print), 1314-8060 (electronic).

Karam:2010:DZD

[Kar10] Jalal Karam. On the distribution of zeros for Daubechies orthogonal wavelets and associated polynomials. In Recent Researches in Applied Mathematics. 15th WSEAS International Conference on Applied Mathematics (MATH ’10), Vouliagmeni, Athens, Greece, December 29–31, 2010, pages 101–104. WSEAS Press, Athens, Greece, 2010. ISBN 960-474-263-9 (hardcover), 960-474-266-3 (CD-ROM).

Klauder:1982:MPI

[KD82a] John R. Klauder and Ingrid Daubechies. Measures for path integrals. Physical Review Letters, 48(3):117–120, January 18, 1982. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Klauder:1982:WMQ

[KD82b] John R. Klauder and Ingrid Daubechies. Wiener measures for quantum-mechanical path-integrals. Lecture Notes in Physics, 173:245–247, 1982. CODEN LNPHA4. ISSN 0075-8450 (print), 1616-6361 (electronic). Stochastic processes in quantum theory and statistical physics, Proc. int. Workshop, Marseille 1981.

Klauder:1984:QMP

[KD84] John R. Klauder and Ingrid Daubechies. Quantum mechanical path integrals with Wiener measures for all polynomial Hamiltonians. Physical Review Letters, 52(14):1161–1164, April 2, 1984. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Kovacevic:1996:SIW

[KD96] Jelena Kovačević and Ingrid Daubechies. Special issue on wavelets. Proceedings of the IEEE, 84(4):507–509, April 1996. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Kehoe:2013:DMR

[Keh13] Elaine Kehoe. Daubechies and Mumford receive BBVA Foundation Award. Notices of the American Mathematical Society, 60(5):603–604, May

2013. URL http://www.ams.
org/notices/201305/rnoti-
p603.pdf.

Kounchev:2010:PDTa

[KK10] Ognyan Kounchev and Damyan Kalaglarsky. Polyharmonic Daubechies type wavelets in image processing and astronomy, I. arXiv.org, ??(??):1–6, June 5, 2010. CODEN ???? ISSN 2331-8422. URL https:

//arxiv.org/abs/1006.1043.

Khare:2010:DMU

[KKJ^{+}10] Ashish Khare, Manish Khare, Yongyeon Jeong, Hongkook Kim, and Moongu Jeon. Despeckling of medical ultrasound images using Daubechies complex wavelet transform. Signal Processing, 90(2):428–439, 2010. CODEN SPRODR. ISSN 0165-1684 (print), 1872-7557 (electronic).

Kounchev:2010:PDTb

[KKT10] Ognyan Kounchev, Damyan Kalaglarsky, and Milcho Tsvetkov. Polyharmonic Daubechies type wavelets in image processing and astronomy, II. arXiv.org, ??(??):1–9, June 29, 2010. CODEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/1006.5739.

Klappenecker:1997:APS

[Kla97] Andreas Klappenecker. On algebraic properties of selfreciprocal polynomials and of Daubechies filters of low order. In Proceedings of IEEE International Symposium on Information Theory: ISlT 1997, Ulm, Germany, June 29–July 4, page 80. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-3956-8.

Kovvali:2005:OAA

[KLC05] Narayan Kovvali, Wenbin Lin, and Lawrence Carin. Order of accuracy analysis for multiresolution time-domain using Daubechies bases. Microwave and Optical Technology Letters, 45(4):290–293, 2005. CODEN MOTLEO. ISSN 1098-2760.

Kateb:1995:ABD

[KLR95a] Djalil Kateb and Pierre Gilles Lemarié-Rieusset. Asymptotic behavior of the Daubechies filters. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 2(4):398–399, 1995. ISSN 1063-5203 (print), 1096-603x (electronic).

Kateb:1995:PFD

[KLR95b] Djalil Kateb and Pierre-Gilles Lemarié-Rieusset. Sur la phase des filtres de Daubechies. Comptes rendus de l’Académie des sciences. Série I, Mathématique, 320(1):5–8, 1995. CODEN CASMEI. ISSN 0764-4442 (print), 1778-3577 (electronic).

Kateb:1997:PDF

[KLR97] Djalil Kateb and Pierre Gilles Lemarié-Rieusset. The phase of the Daubechies filters. Revista Matemática Iberoamericana, 13 (2):245–305, 1997. ISSN 0213-2230 (print), 2235-0616 (electronic).

Kobiler:2010:HCB

[KLT^{+}10] Oren Kobiler, Yaron Lipman, Kate Therkelsen, Ingrid Daubechies, and Lynn W. Enquist. Herpesviruses carrying a Brainbow cassette reveal replication and expression of limited numbers of incoming genomes. Nature Communications, 1(9):146–153, December 21, 2010. CODEN NCAOBW. ISSN 2041-1723 (electronic).

URL http://www.nature.
com/ncomms/journal/v1/n9/
full/ncomms1145.html.

Karam:2012:RDP

[KM12] Jalal Karam and Samer E. Mansour. On the roots of Daubechies polynomials for biorthogonal wavelets. In 2012 International Conference on Communications and Information Technology (ICCIT). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 2012.

Kozbial:2006:ADW

[Koz06] Tomasz Kozbial. Application of Daubechies wavelets approximation to plate bending. Proceedings in Applied Mathematics and Mechanics, 6(1):231– 232, December 2006. ISSN 1617-7061.

Khare:2009:DCW

[KTJ09] Ashish Khare, Uma Shanker Tiwary, and Moongu Jeon. Daubechies complex wavelet transform based multilevel shrinkage for deblurring of medical images in presence of noise. International Journal of Wavelets, Multiresolution and Information Processing, 7(5): 587–604, 2009. ISSN 0219-6913 (print), 1793-690X (electronic).

Lai:1995:DFA

[Lai95] Ming-Jun Lai. On the digital filter associated with Daubechies’ wavelets. IEEE Transactions on Signal Processing, 43(9):2203–2205, 1995. CODEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic).

Lipman:2010:SFE

[LCDF10] Yaron Lipman, Xiaobai Chen, Ingrid Daubechies, and Thomas Funkhouser. Symmetry factored embedding and distance. ACM Transactions on Graphics, 29(4):103:1–103:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lina:1996:IPS

[LD96] J.-M. Lina and P. Drouilly. The importance of the phase of the symmetric Daubechies wavelets representation of signals. In Proceedings IWISP ’96, pages 61–64. Elsevier, 1996.

Lipman:2009:SCM

[LD09] Yaron Lipman and Ingrid Daubechies. Surface comparison with mass transportation. arXiv.org, ??(??):1–36, December 17, 2009. CODEN ???? ISSN 2331-8422. URL https:

//arxiv.org/abs/0912.3488.

Lipman:2011:CWDa

[LD11a] Yaron Lipman and Ingrid Daubechies. Conformal Wasserstein distances: comparing surfaces in polynomial time. arXiv.org, ??(??):1–23, March 22, 2011. CODEN ???? ISSN 2331-8422. URL https:/

/arxiv.org/abs/1103.4408.

Published in Advances in Mathematics, vol. 227 (2010) pp. 1047–1077.

Lipman:2011:CWDb

[LD11b] Yaron Lipman and Ingrid Daubechies. Conformal Wasserstein distances: comparing surfaces in polynomial time. Advances in Mathematics, 227(3): 1047–1077, June 20, 2011. CO-DEN ADMTA4. ISSN 0001-8708 (print), 1090-2082 (electronic).

Liu:2016:NMB

[LD16] Yanan Liu and Keqin Din. A numerical method based on Daubechies wavelet basis and B-spline patches for elasticity problems. Mathematical Problems in Engineering, pages Art. ID 2549213, 11, 2016. ISSN 1024-123X.

Loris:2008:NRT

[LDN^{+}08] Ignace Loris, H. Douma, G. Nolet, Ingrid Daubechies, and C. Regone. Nonlinear regularization techniques for seismic tomography. arXiv.org, ??(??):1–23, August 26, 2008. CODEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/0808.3472.

Loris:2010:NRT

[LDN^{+}10] Ignace Loris, H. Douma, G. Nolet, Ingrid Daubechies, and C. Regone. Nonlinear regularization techniques for seismic tomography. Journal of Computational Physics, 229(3):890–905, February 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0021999109005725.

Lelandais:2008:MMP

[LG08] Fabien Lelandais and Herve Glotin. Mallat’s matching pursuit of sperm whale clicks in real-time using Daubechies 15 wavelets. In 2008 New Trends for Environmental Monitoring Using Passive Systems. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 2008.

Lin:1997:BFA

[Lin97a] Chien-Hsin Lin. Basis function analysis in Daubechies wavelet transformation. Thesis (Ph.D.), Department of Electrical Engineering and Computer Science, Tulane University, New Orleans, LA 70118, USA, December 16, 1997. xi + 181 pp. URL

https://search.proquest.
com/docview/304456212.

Lina:1997:IPC

[Lin97b] Jean-Marc Lina. Image processing with complex Daubechies wavelets. Journal of Mathematical Imaging and Vision: JMIV, 7(3):211–223, 1997. CO-DEN JMIVEK. ISSN 0924-9907 (print), 1573-7683 (electronic).

Lina:1998:CDW

[Lin98] J.-M. Lina. Complex Daubechies wavelets: Filters design and applications. In Inverse problems, tomography, and image processing. Proceedings of sessions from the 1st international congress of ISAAC (International Society for Analysis, Applications and Computing), University of Delaware, Newark, DE, USA, June 3–7, 1997, pages 95–112. Plenum Press, New York, NY, USA; London, UK, 1998. ISBN 0-306-45828-4.

Lewis:1991:VAD

[LK91] A. S. Lewis and G. Knowles. VLSI architecture for 2-D Daubechies wavelet transform without multipliers. Electronics Letters, 27(2):171, 1991. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Lin:2005:DAC

[LKC05] Wenbin Lin, Narayan Kovvali, and Lawrence Carin. Direct algorithm for computation of derivatives of the Daubechies basis functions. Applied Mathematics and Computation, 170 (2):1006–1013, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lin:2000:AIC

[LL00] W. Lin and Q. Li. The algorithm implementation of Cauchy singular integral in Daubechies wavelets on the interval. In Proceedings of the second ISAAC congress. Vol.

1. Proceedings of the International Society for Analysis, its Applications and Computation congress, Fukuoka, Japan, August 16–21, 1999, volume 7 of Int. Soc. Anal. Appl. Comput., pages 153–161. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6597-6 (hardcover).

Liu:2008:DWM

[LLC08] Ya’nan Liu, Yinghua Liu, and Zhangzhi Cen. Daubechies wavelet meshless method for 2-D elastic problems. Tsinghua Sci. Technol., 13(5):605–608, 2008. ISSN 1007-0214.

Liu:2011:MSD

[LLC11] Yanan Liu, Yinghua Liu, and Zhangzhi Cen. Multiscale Daubechies wavelet-based method for 2-D elastic problems. Finite Elements in Analysis and Design, 47(4):334–341, 2011. CODEN FEADEU. ISSN 0168-874X (print), 1872-6925 (electronic).

Lina:1993:PDW

[LM93] Jean-Marc Lina and Michel Mayrand. Parametrizations for Daubechies wavelets. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 48(6): R4160–R4163, December 1, 1993. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic).

Lina:1994:PCD

[LM94] Jean-Marc Lina and Michel Mayrand. Parameterizations for complex Daubechies wavelets. In Harold H. Szu, editor, Wavelet Applications. SPIE Optical Engineering Press, Bellingham, WA, USA, March 1994.

Lina:1995:CDW

[LM95] Jean-Marc Lina and Michel Mayrand. Complex Daubechies wavelets. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 2(3):219–229, 1995. ISSN 1063-5203 (print), 1096-603x (electronic).

Loris:2006:TIU

[LNDD06] Ignace Loris, Guust Nolet, Ingrid Daubechies, and F. A. Dahlen. Tomographic inversion using ℓ_{1}-norm regularization of wavelet coefficients. arXiv.org, ??(??):1–19, August 8, 2006. CODEN ???? ISSN 2331-8422. URL https:

//arxiv.org/abs/physics/

0608094. Published in Geophysical Journal International 170 (2007) 359–370.

Loris:2007:TIU

[LNDD07] Ignace Loris, Guust Nolet, Ingrid Daubechies, and F. A. Dahlen. Tomographic inversion using l(1)-norm regularization of wavelet coefficients. Geophysical journal international, 170(1):359–370, July 2007. CODEN GJINEA. ISSN 0956-540x (print), 1365-246x (electronic).

Li:2001:CDD

[LP01] Dengfeng Li and Silong Peng.

A characterization of ndimensional Daubechies type tensor product wavelet. Acta Mathematicae Applicatae Sinica. English Series. Yingyong Shuxue Xuebao, 17(3):382–392, 2001. ISSN 0168-9673 (print), 1618-3932 (electronic).

Lipman:2011:CWDc

[LPD11] Yaron Lipman, Jesus Puente, and Ingrid Daubechies. Conformal Wasserstein distance: II. Computational aspects and extensions. arXiv.org, ??(??): 1–42, March 24, 2011. CO-DEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/1103.4681.  Pub-

lished in Math. Comp. 82 (2013), 331–381.

Lipman:2013:CWD

[LPD13] Yaron Lipman, Jesus Puente, and Ingrid Daubechies. Conformal Wasserstein distance: II. Computational aspects and extensions. Mathematics of Computation, 82(281):331– 381, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-281/
S0025-5718-2012-02569-5;
 http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-

                                        2012-02569-5/S0025-5718-
                                        2012-02569-5.pdf.

Liu:2010:LDA

[LQLC10a] Yanan Liu, Fei Qin, Yinghua Liu, and Zhangzhi Cen. The 2D large deformation analysis using Daubechies wavelet. Computational mechanics, 45 (2–3):179–187, 2010. CODEN CMMEEE. ISSN 0178-7675 (print), 1432-0924 (electronic).

Liu:2010:DWB

[LQLC10b] Yanan Liu, Fei Qin, Yinghua Liu, and Zhangzhi Cen. A Daubechies wavelet-based method for elastic problems. Engineering Analysis with Boundary Elements, 34(2): 114–121, 2010. ISSN 0955-7997 (print), 1873-197x (electronic).

Lau:2000:ARD

[LS00] Ka-Sing Lau and Qiyu Sun. Asymptotic regularity of Daubechies’ scaling functions. Proceedings of the American Mathematical Society, 128(4):1087– 1095, April 2000. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.2307/119783.

Lieb:2005:OER

[LT05] Elliott H. Lieb and Walter Thirring, editors. One-Electron Relativistic Molecules with Coulomb Interaction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2005. ISBN 3-540-22212-X (hardcover), 3-540-27056-6. xv + 932 pp. LCCN QC173.4.T48 L54 2005. URL

http://link.springer.com/
10.1007/b138553.

Lu:1997:SRD

[Lu97] X. G. Lu. Series representation of Daubechies’ wavelets. Journal of Computational Mathematics, 15(1):81–96, 1997. CO-DEN JCMMEB. ISSN 0254-9409 (print), 1991-7139 (electronic).

Lundberg:1992:BRW

[Lun92]  Matthew Lundberg.  Book

reviews: Wavelets: a Tutorial in Theory and Applications (Volume 2 in series ‘Wavelet Analysis and its Applications’) Charles K. Chui Academic Press, Harcourt Brace Jovanovich, San Diego, CA, 1992; ISBN 0-12-174590-2, 736 pp., hardcover, $69.95. Wavelets and Their Applications Mary Beth Ruskai, Gregory Beylkin, Ronald Coifman, Ingrid Daubechies, Stephane Mallat, Yves Meyer and Louise Raphael, editors Jones and Bartlett, Boston, MA, 1992; ISBN 0-86720-225-4, 480 pp., hardcover, $59.95. Computers in Physics, 6(6):698– ??, November 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.4823128.

Li:2009:CCD

[LW09] Dengfeng Li and Guochang Wu. Construction of a class of Daubechies type wavelet bases. Chaos, Solitons & Fractals, 42(1):620–625, October 2009. CODEN CSFOEH. ISSN 0960-0779 (print), 1873-2887 (electronic).

Li:2011:DWM

[LXDS11] Ying Li, Maohui Xia, Kai Dong, and Hongling Shen. Daubechies wavelet meshless method of numerical calculation in the electromagnetic field. In 2011 International Conference on Multimedia Technology. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, July 2011.

Zhao:2008:DWA

[lZqJmTjZ08] Yu lei Zhao, Zhi qin Ju, Chuang ming Tong, and Wei jun Zhong. Daubechies wavelets in analyzing grooved wire. In 2008 International Conference on Microwave and Millimeter Wave Technology. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 2008.

Ma:2016:EVD

[Ma16] Hajji Ma. On the exact values of Daubechies wavelets. Journal of Physical Mathematics, 07(01), 2016. ISSN 2090-0899 (print), 2090-0902 (electronic).

Maleknejad:2006:CMH

[MD06] K. Maleknejad and H. Derili. The collocation method for Hammerstein equations by Daubechies wavelets. Applied Mathematics and Computation, 172(2):846–864, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Moayeri:1992:WTI

[MDSW92] N. Moayeri, Ingrid Daubechies, Q. Song, and H. S. Wang. Wavelet transform image coding using trellis coded vector quantization. In [Proceedings] ICASSP-92: 1992 IEEE International Conference on Acoustics, Speech, and Signal Processing, volume 4, pages 405– 408. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, March 1992. ISBN 0-7803-0532-9. ISSN 1520-6149 (print), 2379-190X (electronic).

Madishetty:2013:VAT

[MMC^{+}13] Shiva Kumar Madishetty, Arjuna Madanayake, Renato J. Cintra, Vassil S. Dimitrov, and Dale H. Mugler. VLSI architectures for the 4-tap and 6tap 2-D Daubechies wavelet filters using algebraic integers. IEEE Transactions on Circuits and Systems I: Regular Papers, 60(6):1455–1468, 2013. ISSN 1549-8328 (print), 1558-0806 (electronic).

Moraru:2011:NUI

[MMN^{+}11] Luminita Moraru, Simona Moldovanu, Mariana Carmen Nicolae, Madalin Bunoiu, and Iosif Malaescu. De-noising ultrasound images of colon tumors using Daubechies wavelet transform. AIP Conference Proceedings, 1387:294–??, 2011. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Morel:2023:EAI

[Mor23] Jean-Michel Morel. Un entretien avec Ingrid Daubechies: (partie 1). Gazette des Mathématiciens, (177):56–59, 2023. ISSN 0224-8999.

Morel:2024:EAI

[Mor24] Jean-Michel Morel. Un entretien avec Ingrid Daubechies: partie 3. (French) [An interview with Ingrid Daubechies: part 3]. Gazette des Mathématiciens, (179):40–44, 2024. ISSN 0224-8999.

Mohr:2014:DWLa

[MRB^{+}14a] Stephan Mohr, Laura E. Ratcliff, Paul Boulanger, Luigi Genovese, Damien Caliste, Thierry Deutsch, and Stefan Goedecker. Daubechies wavelets for linear scaling density functional theory. arXiv.org, ??(??):1–18, January 29, 2014. CODEN ???? ISSN 2331-8422. URL https:

                                        //arxiv.org/abs/1401.7441.

Mohr:2014:DWLb

[MRB^{+}14b] Stephan Mohr, Laura E. Ratcliff, Paul Boulanger, Luigi Genovese, Damien Caliste, Thierry Deutsch, and Stefan Goedecker. Daubechies wavelets for linear scaling density functional theory. Journal of Chemical Physics, 140(20): 204110:1–204110:16, May 2014. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Maleknejad:2007:CMI

[MYN07] K. Maleknejad, M. Yousefi, and K. Nouri. Computational methods for integrals involving functions and Daubechies wavelets. Applied Mathematics and Computation, 189(2): 1828–1840, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Neelov:2005:ENQ

[NG05] A. I. Neelov and S. Goedecker.

An efficient numerical quadrature for the calculation of the potential energy of wavefunctions expressed in the Daubechies wavelet basis. arXiv.org, ??(??):1–35, June 15, 2005. CODEN ???? ISSN 2331-8422. URL https:

//arxiv.org/abs/physics/
0506139.

Neelov:2006:ENQ

[NG06] A. I. Neelov and S. Goedecker.

An efficient numerical quadrature for the calculation of the potential energy of wavefunctions expressed in the Daubechies wavelet basis. Journal of Computational Physics, 217(2):312–339, September 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Nievergelt:1999:ADW

[Nie99a] Yves Nievergelt. Algorithms for Daubechies wavelets. In Wavelets Made Easy, pages 73– 113. Birkhäuser Boston Inc., Cambridge, MA, USA, 1999.

Nievergelt:1999:DWD

[Nie99b] Yves Nievergelt. Daubechies wavelets design. In Wavelets Made Easy, pages 238–261. Birkhäuser Boston Inc., Cambridge, MA, USA, 1999.

Nievergelt:2012:ADW

[Nie12] Yves Nievergelt. Algorithms for Daubechies wavelets. In Wavelets Made Easy, pages 73– 113. Springer New York, October 2012.

Necula:2012:TAF

[NKM12] V. Necula, S. Klimenko, and G. Mitselmakher. Transient analysis with fast Wilson– Daubechies time-frequency transform. Journal of Physics: Conference Series, 363:012032, 2012. CODEN JPCSDZ. ISSN 1742-6588 (print), 1742-6596 (electronic).

NiranjanaMurthy:2013:ESD

[NM13] H. S. NiranjanaMurthy and M. Meenakshi. ECG signal denoising and ischemic event feature extraction using Daubechies wavelets. International Journal of Computer Applications, 67(2):29– 33, April 2013. ISSN 0975-8887.

Novikov:1995:MDW

[Nov95]  I. Ya. Novikov.  Modified

Daubechies wavelets preserving localization with growth of smoothness. East Journal on Approximations, 1(3):341–348, 1995. ISSN 1310-6236.

Novikov:1998:UCM

[Nov98]  I. Ya. Novikov.  Uncer-

tainty constants for modified Daubechies wavelets. Izv. Tul. Gos. Univ. Ser. Mat. Mekh. Inform., 4(1):107–111, 165, 1998. Proceedings of the International Conference “Approximation Theory and Harmonic Analysis” (Russian) (Tula, 1998).

Novikov:2002:ARBa

[Nov02a] I. Ya. Novikov. Asymptotics of the roots of Bernstein polynomials used in the construction of modified Daubechies wavelets. Matematicheskie Zametki, 71(2):239–253, 2002. ISSN 0025-567X.

Novikov:2002:ARBb

[Nov02b] I. Ya. Novikov. Asymptotics of the roots of Bernstein polynomials used in the construction of modified Daubechies wavelets. Mathematical Notes of the Academy of Sciences of the USSR = Matematicheskie Zametki, 71(2):217–229, 2002. CODEN MTHNB2. ISSN 0001-4346 (print), 1573-8876 (electronic).

Novikov:2002:AZB

[Nov02c]  I. Ya. Novikov.  Asymp-

totics of zeros of Bernstein polynomials that are related to modified Daubechies wavelets. In Wavelet analysis and applications. Proceedings of an international conference, Guangzhou, China, November 15–20, 1999, volume 25 of AMS/IP Stud. Adv. Math., pages 225–234. American Mathematical Society, Providence, RI, USA, 2002. ISBN 0-8218-2991-2 (paperback).

Nikam:2015:CIE

[NY15] Swati D. Nikam and Rajesh U. Yawale. Color image enhancement using Daubechies wavelet transform and HIS color model. In 2015 International Conference on Industrial Instrumentation and Control (ICIC). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 2015.

OldeDaalhuis:1992:CDW

[Old92] A. B. Olde Daalhuis. Computing with Daubechies’ wavelets. CWI Quarterly, 5(1):63–72, 1992. ISSN 0922-5366.

OjedaGonzalez:2014:DWCa

[OMOE14a] Arian Ojeda González, Odim Mendes Junior, Margarete Oliveira Domingues, and Varlei Everton Menconi. Daubechies wavelet coefficients: a tool to study interplanetary magnetic field fluctuations. arXiv.org, ??(??):1–15, April 10, 2014. CODEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/1404.2835.  Pub-

lished in Geofisica Internacional, 53-2: 101-115, ISSN: 0016-7169, 2014.

OjedaGonzalez:2014:DWCb

[OMOE14b] Arian Ojeda González, Odim Mendes Junior, Margarete Oliveira Domingues, and Varlei Everton Menconi. Daubechies wavelet coefficients: a tool to study interplanetary magnetic field fluctuations. Geofísica Internacional, 53(2):101–115, ???? 2014. ISSN 0016-7169.

URL http://www.geofisica.
unam.mx/unid_apoyo/editorial/
publicaciones/investigacion/
geofisica_internacional/
anteriores/2014/02/1_ojeda.
pdf.

ONeal:2016:EJW

[OWW^{+}16] Wesley T. O’Neal, Yi (Grace)

Wang, Hau-Tieng Wu, Zhu-Ming Zhang, Yabing Li, Larisa G. Tereshchenko, E. Harvey Estes, Ingrid Daubechies, and Elsayed Z. Soliman. Electrocardiographic J wave and cardiovascular outcomes in the general population (from the.

Atherosclerosis Risk In Communities Study). American Journal of Cardiology, 118(6): 811–815, September 15, 2016. CODEN AJCDAG. ISSN 0002-9149 (print), 1879-1913 (electronic).

P:2015:SKE

[P15] Riskyana Dewi Intan P. Studi komparasi ekstraksi fitur pada pengenalan wajah menggunakan principal component analysis (PCA) dan wavelet Daubechies. (Indonesian) [Comparative study of feature extraction on facial recognition using Principal Component Analysis (PCA) and Daubechies wavelets]. Jurnal Masyarakat Informatika, 6(12), October 2015.

Pierpaoli:2004:RSZ

[PAHD04] E. Pierpaoli, S. Anthoine, K. Huffenberger, and Ing Daubechies. Reconstructing Sunyaev–Zeldovich clusters in future CMB experiments. arXiv.org, ??(??):1–12, December 9, 2004. CODEN ???? ISSN 2331-8422. URL https:

                                        //arxiv.org/abs/astro-ph/

0412197. Published in Mon. Not. Roy. Astron. Soc. 359 (2005) 261–271.

Pierpaoli:2005:RSZ

[PAHD05] Elena Pierpaoli, S. Anthoine, Kevin M. Huffenberger, and Ingrid Daubechies. Reconstructing Sunyaev–Zel’dovich clusters in future cosmic microwave background experiments. Monthly Notices of the Royal Astronomical Society, 359(1):261–271, May 1, 2005. CODEN MNRAA4. ISSN 0035-8711 (print), 1365-2966 (electronic).

Puente:2013:AAG

[PBGD13] Jesus Puente, Douglas M.

Boyer, Justin T. Gladman, and Ingrid C. Daubechies. Automated approaches to geometric morphometrics. American Journal of Physical Anthropology, 150(56, SI):226, 2013. CO-DEN AJPNA9. ISSN 0002-9483 (print), 1096-8644 (electronic). 82nd Annual Meeting of the American-Association-of-Physical-Anthropologists, Knoxville, TN, APR 09–13, 2013.

Platisa:2011:SFC

[PCR^{+}11] L. Platiša, B. Cornells, T. Ružić, A. Pižurica, A. Dooms, M. Martens, M. De Mey, and Ingrid Daubechies. Spatiogram features to characterize pearls in paintings. In 2011 18th IEEE International Conference on Image Processing, pages 801–

804. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 2011. ISSN 1522-4880 (print), 2381-8549 (electronic).

Pilz:2009:IGG

[Pil09] Jürgen Pilz, editor. Interfacing Geostatistics and GIS. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-33235-9 (hardcover), 1-281-92749-X, 3-540-33236-7 (e-book). xv + 282 pp. LCCN QE33.2.S82 I58

2009.  URL http://www.
springerlink.com/openurl.
asp?genre=book%26isbn=978-
3-540-33236-7.

Polatkan:2009:DFP

[PJB^{+}09] Gungor Polatkan, Sina Jafarpour, Andrei Brasoveanu, Shannon Hughes, and Ingrid Daubechies. Detection of forgery in paintings using supervised learning. In 2009 16th IEEE International Conference on Image Processing (ICIP), volume ????, pages 2921–2924. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, November 2009. ISSN 1522-4880 (print), 2381-8549 (electronic).

Pahlavan:2013:SFF

[PKG13] Lotfollah Pahlavan, Christos Kassapoglou, and Zafer Gürdal. Spectral formulation of finite element methods using Daubechies compactlysupported wavelets for elastic wave propagation simulation. Wave Motion, 50(3):558– 578, 2013. CODEN WAMOD9. ISSN 0165-2125 (print), 1878-433x (electronic).

Patton:1996:ODF

[PM96] Richard D. Patton and Patrick C. Marks. One-dimensional finite elements based on the Daubechies family of wavelets. American Institute of Aeronautics and Astronautics Journal, 34(8):1696–1698, 1996. CO-DEN AIAJAH. ISSN 0001-1452 (print), 1533-385X (electronic).

Panja:2011:NOP

[PM11] M. M. Panja and B. N. Mandal.

A note on one-point quadrature formula for Daubechies scale function with partial support. Applied Mathematics and Computation, 218(8): 4147–4151, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Panja:2013:DSF

[PM13a] M. M. Panja and B. N. Mandal. Daubechies scale function based quadrature rules for singular and hypersingular integrals with variable singularities. Investigations in Mathematical Sciences, 3(1):155–176, 2013. ISSN 2250-1436.

Panja:2013:SSK

[PM13b] M. M. Panja and B. N. Mandal. Solution of second kind integral equation with Cauchy type kernel using Daubechies scale function. Journal of Computational and Applied Mathematics, 241:130–142, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Panja:2015:GTQ

[PM15] M. M. Panja and B. N. Mandal. Gauss-type quadrature rule with complex nodes and weights for integrals involving Daubechies scale functions and wavelets. Journal of Computational and Applied Mathematics, 290:609–632, 2015. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

P:2016:DIW

[PMK16] Saravanan P, Sreekara M, and Manikantan K. Digital image watermarking using Daubechies wavelets. In 2016 3rd International Conference on Signal Processing and Integrated Networks (SPIN). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, February 2016.

Pollen:1992:DSF

[Pol92a] David Pollen. Daubechies’ scaling function on [0, 3]. In Wavelets: a tutorial in theory and applications, volume 2 of Wavelet Anal. Appl., pages 3– 13. Academic Press, New York, NY, USA, 1992. ISBN 0-12-174590-2 (hardcover).

Pollen:1992:DXF

[Pol92b] David Pollen. Daubechies’ xf scaling function on [0, 3]. In Wavelets, pages 3–13. Elsevier,

Pollen:1992:DXS

[Pol92c] David Pollen. Daubechies’ xf scaling function on [0, 3]. In Wavelets, pages 3–13. Elsevier, 1992.

Pizurica:2015:DIP

[PPR^{+}15] Aleksandra Pizurica, Ljiljana Platiša, Tijana Ružić, Bruno Cornelis, Ann Dooms, Maximiliaan Martens, Helene Dubois, Bart Devolder, Marc De Mey, and Ingrid Daubechies. Digital image processing of the ghent altarpiece: Supporting the painting’s study and conservation treatment. IEEE Signal Processing Magazine, 32 (4):112–122, July 2015. CO-DEN ISPRE6. ISSN 1053-5888 (print), 1558-0792 (electronic).

Paskyabi:2005:WAP

[PR05] Mostafa Bakhoday Paskyabi and Farzan Rashidi. Wide angle parabolic equation based on one-periodic Daubechies wavelet for modelling underwater wave propagation. WSEAS Transactions on Mathematics, 4(3):204–211, 2005. ISSN 1109-2769 (print), 2224-2880 (electronic).

Phillies:1995:MZD

[PS95] George D. J. Phillies and Jonathan Stott. Mori– Zwanzig–Daubechies decomposition of Ising-model Monte Carlo dynamics. Computers in Physics, 9(2):225–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

                                        aip.scitation.org/doi/10.
                                        1063/1.168527.

Panja:2016:CES

[PSB^{+}16] M. M. Panja, M. K. Saha, U. Basu, D. Datta, and B. N. Mandal. Computing eigenelements of Sturm–Liouville problems by using Daubechies wavelets. Indian Journal of Pure and Applied Mathematics, 47(3):553–579, 2016. ISSN 0019-5588 (print), 0975-7465 (electronic).

Polatkan:2012:BNA

[PZC^{+}12] Gungor Polatkan, Mingyuan Zhou, Lawrence Carin, David Blei, and Ingrid Daubechies. A Bayesian nonparametric approach to image superresolution. arXiv.org, ??(??): 1–30, September 22, 2012. CODEN ???? ISSN 2331-

                                        8422. URL https://arxiv.
                                        org/abs/1209.5019.

Polatkan:2015:BNA

[PZC^{+}15] Guengoer Polatkan, Mingyuan Zhou, Lawrence Carin, David Blei, and Ingrid Daubechies. A Bayesian nonparametric approach to image superresolution. IEEE Transactions on Pattern Analysis and Machine Intelligence, 37(2):346– 358, February 2015. CODEN ITPIDJ. ISSN 0162-8828.

Qixiang:2012:CMS

[Qix12] Yang Qixiang. Characterization of multiplier spaces with Daubechies wavelets. Acta Mathematica Scientia, 32(6): 2315–2321, November 2012. ISSN 0252-9602.

Rowe:1995:DWM

[RA95] Alistair C. H. Rowe and Paul C. Abbott. Daubechies wavelets and Mathematica. Computers in Physics, 9(6): 635–??, November 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.168556.

Rakowski:1998:DFWb

[RB98a] Waldemar Rakowski and Zbigniew Bartosiewicz. Daubechies filters for 2D wavelet transforms. In Rough Sets and Current Trends in Computing, pages 369–372. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998.

Rakowski:1998:DFWa

[RB98b] Waldemar Rakowski and Zbigniew Bartosiewicz. Daubechies filters in wavelet image compression. Fundamenta Informaticae, 34(4):455–467, 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ruskai:1992:WTA

[RBC^{+}92] Mary Beth Ruskai, Gregory Beylkin, Ronald Coifman, Ingrid Daubechies, Stéphane Mallat, Yves Meyer, and Louise Raphael, editors. Wavelets and Their Applications. Jones and Bartlett, Boston, MA, USA, 1992. ISBN 0-86720-225-4 (hardcover). xiii + 474 pp. LCCN QA403.3 .W38 1992.

Ruzic:2011:VRG

[RCP^{+}11] Tijana Ružić, Bruno Cornelis, Ljiljana Platiša, Aleksandra Pižurica, Ann Dooms, Wilfried Philips, Maximiliaan Martens, Marc De Mey, and Ingrid Daubechies. Virtual restoration of the Ghent altarpiece using crack detection and inpainting. In Jacques Blanc-Talon, editor, Advanced concepts for intelligent vision systems: 13th International Conference, ACIVS 2011, Ghent, Belgium, August 22-25, 2011, proceedings, volume 6915 of Lecture Notes in Computer Science, pages 417–428. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-23687-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1634 .A25 2011.

Rudin:2003:DAC

[RDS04a] Cynthia Rudin, Ingrid Daubechies, and Robert E. Schapire. The dynamics of AdaBoost: cyclic behavior and convergence of margins. Journal of Machine Learning Research (JMLR), 5: 1557–1595, 2003–2004. CO-DEN JMLRAJ. ISSN 1532-4435 (print), 1533-7928 (electronic).

Rudin:2004:DAC

[RDS04b] Cynthia Rudin, Ingrid Daubechies, and Robert E. Schapire. The dynamics of AdaBoost: Cyclic behavior and convergence of margins. Journal of Machine Learning Research (JMLR), 5:1557–1595, December 2004. CODEN JMLRAJ. ISSN 1532-4435 (print), 1533-7928 (electronic).

Rudin:2003:DB

[RDS04c] Cynthia Rudin, Ingrid Daubechies, and Robert E. Schapire. On the dynamics of boosting. In Thrun et al. [TSS04], pages 1101–1108. ISBN 0-262-20152-6 (hardcover). LCCN QA76.87.

URL http://papers.nips.
cc/paper/2535-on-the-dynamics-
of-boosting.pdf.

Redko:2015:FCR

[Red15] I. V. Redko. On Fourier coefficients with respect to the system of Daubechies wavelets {ψ_{jk}^{(2)}}_{j,k∈Z}. Vestn. Voronezh. Gos. Univ., Ser. Fiz. Mat., 2015(1):168–176, 2015. ISSN 1609-0705.

Rodionov:2009:ESDa

[RF09a] E. A. Rodionov and Yu. A.

Farkov. Estimates for the smoothness of dyadic orthogonal wavelets of Daubechies type. Matematicheskie Zametki, 86(3):429–444, 2009. ISSN 0025-567X.

Rodionov:2009:ESDb

[RF09b] E. A. Rodionov and Yu.A. Farkov. Estimates of the smoothness of dyadic orthogonal wavelets of Daubechies type. Mathematical Notes of the Academy of Sciences of the USSR = Matematicheskie Zametki, 86(3):407–421, 2009. CODEN MTHNB2. ISSN 0001-4346 (print), 1573-8876 (electronic).

Ratcliff:2015:FACa

[RGMD15a] Laura E. Ratcliff, Luigi Genovese, Stephan Mohr, and Thierry Deutsch. Fragment approach to constrained density functional theory calculations using Daubechies wavelets. arXiv.org, ??(??):1–14, March 20, 2015. CODEN ???? ISSN 2331-8422. URL https://

arxiv.org/abs/1503.06229.

Ratcliff:2015:FACb

[RGMD15b] Laura E. Ratcliff, Luigi Genovese, Stephan Mohr, and Thierry Deutsch. Fragment approach to constrained density functional theory calculations using Daubechies wavelets. Journal of Chemical Physics, 142(23):234105:1– 234105:15, June 2015. CO-DEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Restrepo:1997:IPC

[RL97] Juan Mario Restrepo and Gary K. Leaf. Inner product computations using periodized Daubechies wavelets. International Journal for Numerical Methods in Engineering, 40(19):3557–3578, 1997. CO-DEN IJNMBH. ISSN 0029-5981 (print), 1097-0207 (electronic).

Restrepo:1996:PDW

[RLS96] J. M. Restrepo, G. K. Leaf, and G. Schlossnagle. Periodized Daubechies wavelets. Technical Report MCS-P423-0394, DE96007647, Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, IL, USA, March 1, 1996.

Rosiek:1995:DBV

[RM95] Mark R. Rosiek and Mark L.

McLeod. Daubechies-based visual table of contents. In Harold H. Szu, editor, Wavelet Applications II. SPIE Optical Engineering Press, Bellingham, WA, USA, April 1995.

Roussos:2005:VBL

[RRD05] E. Roussos, S. Roberts, and Ingrid Daubechies. Variational Bayesian learning for wavelet independent component analysis. In AIP Conference Proceedings, volume 803, pages 274–281. American Institute of Physics, Woodbury, NY, USA, 2005. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Roussos:2012:VBL

[RRD12] E. Roussos, S. Roberts, and Ingrid Daubechies. Variational Bayesian learning of sparse representations and its application in functional neuroimaging. Lecture Notes in Computer Science, 7263:218–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rafiee:2009:ADM

[RRPT09] J. Rafiee, M. A. Rafiee, N. Prause, and P. W. Tse. Application of Daubechies 44 in machine fault diagnostics. In Louise Marston, editor, 2009 2nd International Conference on Computer, Control and Communication (IC4 2009): 17–18 February 2009, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, February 2009. ISBN 1-4244-3313-4 (print), 1-4244-3314-2 (CD-ROM), 1-4244-3312-6 (electronic). IEEE Catalog Number CFP0963.

Rudin:2004:BBS

[RSD04] Cynthia Rudin, Robert E. Schapire, and Ingrid Daubechies. Boosting based on a smooth margin. In John Shawe-Taylor and Yoram Singer, editors, Learning Theory. 17th Annual Conference on Learning Theory, COLT 2004, Banff, Canada, July 1–4, 2004. Proceedings, volume 3120 of Lecture Notes in Artificial Intelligence, pages 502–517. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-22282-0, 3-540-27819-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334-342.

Rudin:2007:ABA

[RSD07a] Cynthia Rudin, Robert E.

Schapire, and Ingrid Daubechies. Analysis of boosting algorithms using the smooth margin function. Annals of Statistics, 35 (6):2723–2768, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.
aos/1201012978;   http://
www.jstor.org/stable/25464607.

Rudin:2007:PSC

[RSD07b] Cynthia Rudin, Robert E.

Schapire, and Ingrid Daubechies. Precise statements of convergence for AdaBoost and arc-gv. In Prediction and Discovery. AMS-IMS-SIAM Joint Summer Research Conference on Machine and Statistical Learning: Prediction and Discovery, Snowbird, UT, USA, June 25–29, 2006, volume 443 of Contemp. Math., pages 131– 145. American Mathematical Society, Providence, RI, USA, 2007. ISBN 0-8218-4195-5 (paperback).

Rudin:2008:ABA

[RSD08] Cynthia Rudin, Robert E.

Schapire, and Ingrid Daubechies.

Analysis of boosting algorithms using the smooth margin function. arXiv.org, ??(??):1–49, March 28, 2008. CODEN ???? ISSN 2331-8422. URL https:

                                        //arxiv.org/abs/0803.4092.

Published in Annals of Statistics 2007, Vol. 35, No. 6, 2723– 2768.

Ruch:2009:DSF

[RV09a] David K. Ruch and Patrick J. Van Fleet. Daubechies Scaling Functions and Wavelets, pages 233–276. Wiley, New York, NY, USA, 2009. ISBN 1-118-16565-9.

Ruch:2009:DDT

[RV09b] David K. Ruch and Patrick J. Van Fleet. The Discrete Daubechies Transformation and Applications, chapter 7, pages 277–323. Wiley, New York, NY, USA, 2009. ISBN 1-118-16565-9.

Storer:2000:DPD

[SC00] James A. (James Andrew) Storer and Martin Cohn, editors. DCC 2000: Proceedings, Data Compression Conference, March 28–30, 2000, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000.

                                        URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=838139. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Storer:2002:DPD

[SC02] James A. (James Andrew)

Storer and Martin Cohn, editors. DCC 2002: proceedings: Data Compression Conference: April 2–4, 2002, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002.

URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=999937. IEEE

Computer Society Order Number PR01477.

Shan:2023:DMU

[SD23] Shan Shan and Ingrid Daubechies.

Diffusion maps: using the semigroup property for parameter tuning. In Theoretical physics, wavelets, analysis, genomics— an indisciplinary tribute to Alex Grossmann, Appl. Numer. Harmon. Anal., pages 409–424. Birkhäuser/Springer, Cham, 2023. ISBN 3-030-45846-6; 3-030-45847-4.

Sumainna:2013:ITW

[SG13] Suma’inna Suma’inna and Gugun Gumilar. Implementasi transformasi wavelet Daubechies pada kompresi citra digital. Cauchy, 2(4):211, May 2013. ISSN 2086-0382 (print), 2477-3344 (electronic).

Satiyan:2010:CPU

[SHN10] M. Satiyan, M. Hariharan, and R. Nagarajan. Comparison of performance using Daubechies wavelet family for facial expression recognition. In Mohd Nasir Taib, Ramli Adnan, Abd Manan Samad, Nooritawati Md Tahir, and Mohd Hezri Fazalul Rahiman, editors, 2010 6th International Colloquium on Signal Processing & its Applications (CSPA 2010): 21– 23 May 2010, Mahkota Hotel, Melaka, Malaysia, pages 354–358. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 2010. ISBN 1-4244-7122-2, 1-4244-7121-4 (print), 1-4244-7120-6 (online). LCCN TK5102.9.

URL http://ieeexplore.
ieee.org/servlet/opac?punumber=
5504579.

Singh:2013:PAD

[Sin13] Raj Ranjan Singh. Performance analysis of Daubechies wavelet in image deblurring and denoising. IOSR Journal of Engineering (IOSRJEN), 3(1): 1–5, January 2013. ISSN 2278-8719 (print), 2250-3021 (electronic).

Safari:2012:FTD

[SK12] Azadeh Safari and Yinan Kong.

Four tap Daubechies filter banks based on RNS. In 2012 International Symposium on Communications and Information Technologies (ISCIT). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 2012.

Simons:2011:WWL

[SLBD11a] Frederik J. Simons, Ignace Loris, Eugene Brevdo, and Ingrid C. Daubechies. Wavelets and wavelet-like transforms on the sphere and their application to geophysical data inversion. Proceedings of the SPIE — The International Society for Optical Engineering, 8138: ??, 2011. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Simons:2011:WWLa

[SLBD11b] Frederik J. Simons, Ignace Loris, Eugene Brevdo, and Ingrid C. Daubechies. Wavelets and wavelet-like transforms on the sphere and their application to geophysical data inversion. arXiv.org, ??(??):1– 15, September 8, 2011. CO-DEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/1109.1718.

Simons:2011:SRGa

[SLN^{+}11a] Frederik J. Simons, Ignace Loris, Guust Nolet, Ingrid C. Daubechies, S. Voronin, J. S.

Judd, P. A. Vetter, J. Charlety, and C. Vonesch. Solving or resolving global tomographic models with spherical wavelets, and the scale and sparsity of seismic heterogeneity. arXiv.org, ??(??):1–43, April 14, 2011. CODEN ???? ISSN 2331-8422. URL https:

//arxiv.org/abs/1104.3151.

Simons:2011:SRGb

[SLN^{+}11b] Frederik J. Simons, Ignace Loris, Guust Nolet, Ingrid C. Daubechies, S. Voronin, J. S. Judd, P. A. Vetter, J. Charlety, and C. Vonesch. Solving or resolving global tomographic models with spherical wavelets, and the scale and sparsity of seismic heterogeneity. Geophysical journal international, 187(2):969– 988, November 2011. CO-DEN GJINEA. ISSN 0956-540x (print), 1365-246x (electronic).

Sengupta:2012:SAI

[SM12] Madhumita Sengupta and J. K. Mandal. Self authentication of image through Daubechies transform technique (SADT). arXiv.org, ??(??):1–4, December 9, 2012. CODEN ???? ISSN 2331-8422. URL https:

//arxiv.org/abs/1212.1863.

Sanders:2017:MHO

[SP17] Toby Sanders and Rodrigo B. Platte. Multiscale higher order TV operators for ℓ_{1} regularization and their relationship to Daubechies wavelets. arXiv.org, ??(??):1–21, March 5, 2017. CODEN ???? ISSN 2331-8422. URL https://

arxiv.org/abs/1703.02404.

Shen:1996:AAD

[SS96] Jianhong Shen and Gilbert Strang. Asymptotic analysis of Daubechies polynomials. Proceedings of the American Mathematical Society, 124(12):3819– 3833, December 1996. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

URL http://www.jstor.org/
stable/10.2307/2161554.

Shen:1998:ADF

[SS98] Jianhong Shen and Gilbert Strang. Asymptotics of Daubechies filters, scaling functions, and wavelets. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 5(3):312–331, 1998. ISSN 1063-5203 (print), 1096-603x (electronic).

Shrestha:2015:CIS

[ST15] Ajaya Shrestha and Arun Timalsina. Color image steganography technique using Daubechies discrete wavelet transform. In 2015 9th International Conference on Software, Knowledge, Information Management and Applications (SKIMA). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, December 2015.

Starckx:2015:FPI

[Sta15] Senne Starckx. Flemish professor [Ingrid Daubechies] on how maths can change the world. FlandersToday, ??(??): ??, October 14, 2015. URL

                                        http://www.flanderstoday.
                                        eu/innovation/flemish-professor-
                                        how-maths-can-change-world.

Soman:2009:DLA

[STAV09] K. P. Soman, Arathi T, Mridula Sara Augustine, and Arunima S. V. Daubechieslagarias algorithm — a simplified approach. In 2009 International Conference on Advances in Computing, Control, and Telecommunication Technologies. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, December 2009.

Strang:1992:OCD

[Str92] Gilbert Strang. The optimal coefficients in Daubechies wavelets. Physica. D, Nonlinear Phenomena, 60(1–4):239– 244, 1992. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic). Experimental mathematics: computational issues in nonlinear science (Los Alamos, NM, 1991).

Sudarvizhi:2016:FBI

[Sud16] D. Sudarvizhi. Feature based image retrieval system using Zernike moments and Daubechies Wavelet Transform. In 2016 International Conference on Recent Trends in Information Technology (ICRTIT). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 2016.

Sun:1999:SEE

[Sun99] Qiyu Sun. Sobolev exponent estimate and asymptotic regularity of the M-band Daubechies’ scaling functions. Constructive Approximation, 15(3):441–465, 1999. ISSN 0176-4276 (print), 1432-0940 (electronic).

Srivastava:2011:MTW

[SYSP11] Madhur Srivastava, Yashwant Yashu, Satish K. Singh, and Prasanta K. Panigrahi. Multisegmentation through wavelets: Comparing the efficacy of Daubechies vs Coiflets. In ????, editor, Proceedings in Signal Processing and Real Time Operating System (SPRTOS), March 26–27, 2011, page ?? ????, ????, 2011.

Srivastava:2012:MTW

[SYSP12] Madhur Srivastava, Yashwant Yashu, Satish K. Singh, and Prasanta K. Panigrahi. Multisegmentation through wavelets: Comparing the efficacy of Daubechies vs Coiflets. arXiv.org, ??(??): 1–4, July 20, 2012. CO-DEN ???? ISSN 2331-

8422. URL https://arxiv.

                                        org/abs/1207.5007.  Pub-

lished in Proceedings in Signal Processing and Real Time Operating System (SPRTOS), March 26–27, 2011.

Sun:1997:IRA

[SZH97] Qi Yu Sun, Ze Yin Zhang, and Da Ren Huang. An integral representation for nary Daubechies filters. Kexue Tongbao (Chinese), 42(8):807– 810, 1997. ISSN 0023-074X.

Taswell:1999:LMD

[Tas99]  C. Taswell.  Least and

most disjoint root sets for Daubechies wavelets. In 1999 IEEE International Conference on Acoustics, Speech, and Signal Processing. Proceedings. ICASSP99 (Cat. No.99CH36258). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999.

Taswell:2000:CSS

[Tas00]  Carl Taswell.  Constraint-

selected and search-optimized families of Daubechies wavelet filters computable by spectral factorization. Journal of Computational and Applied Mathematics, 121(1–2):179– 195, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Numerical analysis in the 20th century, Vol. I, Approximation theory.

Tay:2008:DWA

[Tay08] D. B. H. Tay. Daubechies wavelets as approximate hilbertpairs? IEEE Signal Processing Letters, 15:57–60, 2008. CO-DEN ISPLEM. ISSN 1070-9908 (print), 1558-2361 (electronic).

Tieng:1994:CDW

[TB94] Quang Minh Tieng and W. W.

Boles. Complex Daubechies wavelet based affine invariant representation for object recognition. In Proceedings of 1st International Conference on Image Processing, pages 198–202. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994.

Tiwari:2015:CMI

[TDK15] Prabhat Kumar Tiwari, Bhanu Devi, and Yogendera Kumar. Compression of MRT images using Daubechies 9/7 and thresholding technique. In International Conference on Computing, Communication & Automation. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 2015.

Temme:1996:ANZ

[Tem96] Nico M. Temme. Asymptotics and numerics of zeros of polynomials that are related to Daubechies wavelets. arXiv.org, ??(??):1–13, October 21, 1996. CODEN ???? ISSN 2331-8422.

URL https://arxiv.org/
abs/math/9610225. Report

number OP-SF 21 October

1996.

Temme:1997:ANZ

[Tem97] Nico M. Temme. Asymptotics and numerics of zeros of polynomials that are related to Daubechies wavelets. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 4(4):414–428, 1997. ISSN 1063-5203 (print), 1096-603x (electronic).

Thrun:2004:ANI

[TSS04] Sebastian Thrun, Lawrence K. Saul, and Bernhard Schölkopf, editors. Advances in neural information processing systems 16: (NIPS 2003): proceedings of the 2003 conference. MIT Press, Cambridge, MA, USA, 2004. ISBN 0-262-20152-6 (hardcover). LCCN QA76.87.

                                        URL http://papers.nips.
                                        cc/book/advances-in-neural-
                                        information-processing-systems-
                                        16-2003.

Unser:1997:APC

[UD97] Michael Unser and Ingrid Daubechies. On the approximation power of convolutionbased least squares versus interpolation. IEEE Transactions on Signal Processing, 45 (7):1697–1711, July 1997. CO-DEN ITPRED. ISSN 1053-587x (print), 1941-0476 (electronic).

VanFleet:2008:DWT

[Van08] Patrick J. Van Fleet. Daubechies Wavelet Transformations, chapter 7, pages 223–280. Wiley, New York, NY, USA, 2008. ISBN 1-118-03240-3.

Vonesch:2005:GDW

[VBU05a] C. Vonesch, Thierry Blu, and Michael Unser. Generalized Daubechies wavelets. In Proceedings. (ICASSP ’05). IEEE International Conference on Acoustics, Speech, and Signal Processing, 2005. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005.

Vonesch:2005:GBD

[VBU05b] Cedric Vonesch, Thierry Blu, and Michael Unser. Generalized biorthogonal Daubechies wavelets. In Manos Papadakis, Andrew F. Laine, and Michael A. Unser, editors, Wavelets XI. SPIE Optical Engineering Press, Bellingham, WA, USA, August 2005.

Vonesch:2007:GDW

[VBU07] Cédric Vonesch, Thierry Blu, and Michael Unser. Generalized Daubechies wavelet families. IEEE Transactions on Signal Processing, 55(9):4415– 4429, 2007. CODEN ITPRED. ISSN 1053-587x (print), 1941-0476 (electronic).

Voronin:2015:IRL

[VD15] Sergey Voronin and Ingrid Daubechies. An iteratively reweighted least squares algorithm for sparse regularization. arXiv.org, ??(??):1–17, November 29, 2015. CODEN ???? ISSN 2331-8422. URL https:/

/arxiv.org/abs/1511.08970.

Voronin:2017:IRL

[VD17] Sergey Voronin and Ingrid Daubechies. An iteratively reweighted least squares algorithm for sparse regularization. In Functional Analysis, Harmonic Analysis, and Image Processing: a Collection of Papers in Honor of Björn Jawerth, volume 693 of Contemp. Math., pages 391–411. American Mathematical Society, Providence, RI, USA, January 2017.

Varanis:2013:EPP

[VP13] Marcus V. M. Varanis and Robson Pederiva. Extração de parâmetros para classificação de falhas em motores elétricos utilizando banco de filtros Daubechies. (Portuguese) [Extraction of parameters for classification of motor failures using the Daubechies filter bank.]. In Proceeding Series of the Brazilian Society of Computational and Applied Mathematics. SBMAC: Sociedade de Matemática Aplicada e Computacional, São Carlos, Brasil, October 2013.

Wahid:2011:LCI

[Wah11] Khan Wahid. Low complexity implementation of Daubechies wavelets for medical imaging applications. In Discrete Wavelet Transforms - Algorithms and Applications. In-Tech, August 2011.

Wang:2001:NDS

[Wan01] Yang Wang. On the number of Daubechies scaling functions and a conjecture of Chyzak et al. Experimental Mathematics, 10(1):87–89, 2001. ISSN 1058-6458 (print), 1944-950x (electronic). URL http://

projecteuclid.org/euclid.
em/999188422.

Wang:2007:ACC

[Wan07] Shuzhong Wang. The algorithm for computing connection coefficients with Daubechies wavelet. J. Nat. Sci. Heilongjiang Univ., 24(3):403–406, 2007. ISSN 1001-7011.

Wu:2002:WPI

[WC02] Ru-Shan Wu and Ling Chen.

Wave propagation and imaging using gabor-daubechies beamlets. In Theoretical and Computational Acoustics 2001, pages 661–670. WORLD SCI-ENTIFIC, July 2002.

Welstead:1999:DW

[Wel99a] Stephen Welstead. Daubechies wavelets. In Fractal and Wavelet Image Compression Techniques [Wel99b], chapter 6, pages 119–130. ISBN 0-8194-3503-1 (print), 0-8194-7859-8 (e-PDF), 1-61583-713-2 (e-book). LCCN TA1637 .W45 1999.

Welstead:1999:FWI

[Wel99b] Stephen Welstead. Fractal and Wavelet Image Compression Techniques. SPIE Optical Engineering Press, Bellingham, WA, USA, 1999. ISBN 0-8194-3503-1 (print), 0-8194-7859-8 (e-PDF), 1-61583-713-2 (e-book). xv + 232 pp. LCCN TA1637 .W45 1999.

Wu:2011:OTF

[WFD11] Hau-Tieng Wu, Patrick Flandrin, and Ingrid Daubechies. One or two frequencies? The synchrosqueezing answers. Advances in Adaptive Data Analysis. Theory and Applications, 3 (1–2):29–39, 2011. ISSN 1793-5369 (print), 1793-7175 (electronic).

Wu:2014:EPD

[WHB^{+}14] Hau-Tieng Wu, Shu-Shua Hseu, Mauo-Ying Bien, Yu Ru Kou, and Ingrid Daubechies. Evaluating physiological dynamics via synchrosqueezing: Prediction of ventilator weaning. IEEE Transactions on Biomedical Engineering, 61(3): 736–744, March 2014. CO-DEN IEBEAX. ISSN 0018-9294 (print), 1558-2531 (electronic).

Wu:2016:OEI

[WLD^{+}16] Hau-Tieng Wu, Gregory F. Lewis, Maria I. Davila, Ingrid Daubechies, and Stephen W. Porges. Optimizing estimates of instantaneous heart rate from pulse wave signals with the synchrosqueezing transform. Methods of Information in Medicine = Methodik der Information in der Medizin, 55 (5):463–472, 2016. CODEN MIMCAI. ISSN 0026-1270.

WANG:2006:DWC

[WLW06] Yan Wang, De-Qun Liang, and Bai-Suo Wang. Daubechies wavelet construction using continuation method. In Wavelet Active Media Technology and Information Processing. World Scientific Publishing Company, August 2006.

Wolff:2011:UES

[WMJ^{+}11] J. Wolff, M. Martens, S. Jafarpour, Ingrid Daubechies, and Robert Calderbank. Uncovering elements of style. In 2011 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), pages 1017–1020. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 2011. ISSN 1520-6149 (print), 2379-190X (electronic).

Wong:2002:PDO

[Won02] M. W. Wong. Products of Daubechies operators. In Wavelet Transforms and Localization Operators, pages 124– 128. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2002.

Wong:2011:DW

[Won11] M. W. Wong. Daubechies wavelets. In Discrete Fourier Analysis, pages 79–86. Springer Basel, 2011.

Wang:2005:NMC

[WP05] Yong Ge Wang and Li Zhong Peng. A new method of constructing compactly supported and symmetric Daubechies’s 3-band wavelet systems. Acta Mathematicae Applicatae Sinica. Yingyong Shuxue Xuebao, 28(4):659–667, 2005. ISSN 0254-3079.

Wang:2011:DFB

[WP11] Le Wang and G. Pan. Daubechies framework based Diffracted Gaussian beam approach on parallel computing platforms. In 2011 IEEE International Symposium on Antennas and Propagation (APSURSI). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, July 2011.

Wu:2013:PAU

[WPS^{+}13] Tong Wu, Gungor Polatkan, David Steel, William Brown, Ingrid Daubechies, and Robert Calderbank. Painting analysis using wavelets and probabilistic topic models. In 2013 IEEE International Conference on Image Processing, volume ????, pages 3264–3268. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 2013. ISSN 1522-4880 (print), 2381-8549 (electronic).

Wu:2014:PAU

[WPS^{+}14] Tong Wu, Gungor Polatkan, David Steel, William Brown, Ingrid Daubechies, and Robert Calderbank. Painting analysis using wavelets and probabilistic topic models. arXiv.org, ??(??):1–5, January 26, 2014. CODEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/1401.6638.

Wang:2015:AWD

[WWD^{+}15] Yi (Grace) Wang, Hau-Tieng Wu, Ingrid Daubechies, Yabing Li, E. Harvey Estes, and Elsayed Z. Soliman. Automated J wave detection from digital 12lead electrocardiogram. Journal of Electrocardiology, 48(1): 21–28, January–February 2015. CODEN JECAB4. ISSN 0022-0736 (print), 1532-8430 (electronic).

Wang:1998:CBI

[WWFW98] James Ze Wang, Gio Wiederhold, Oscar Firschein, and Sha Xin Wei. Contentbased image indexing and searching using Daubechies’ wavelets. International Journal on Digital Libraries, 1(4): 311–328, March 1998. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s007990050026.

Xu:2023:PKB

[XLY^{+}23] Jieren Xu, Yitong Li, Haizhao Yang, David Dunson, and Ingrid Daubechies. PiPs: a kernel-based optimization scheme for analyzing nonstationary 1D signals. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 66:1–17, 2023. ISSN 1063-5203 (print), 1096-603X (electronic).

Xie:2010:SWN

[XTZ10] Yanbin Xie, Ju Tang, and Qian Zhou. Suppressing white-noise in partial discharge measurements. Part 1: Construction of complex Daubechies wavelet and complex threshold. European Transactions on Electrical Power, 20(6):800–810, 2010. ISSN 1430-144X (print), 1546-3109 (electronic).

Xue:2007:SED

[XWL07] Yanbo Xue, Jinkuan Wang, and Zhigang Liu. Subbanding ESPRIT by Daubechies wavelets. International Journal of Information & Systems Sciences, 3(4):623–631, 2007. ISSN 1708-296X.

Xu:2016:RDB

[XYD16] Jieren Xu, Haizhao Yang, and Ingrid Daubechies. Recursive diffeomorphism-based regression for shape functions. arXiv.org, ??(??):1–31, October 12, 2016. CODEN ???? ISSN 2331-8422. URL https:/

/arxiv.org/abs/1610.03819.

Xu:2018:RDB

[XYD18] Jieren Xu, Haizhao Yang, and Ingrid Daubechies. Recursive diffeomorphism-based regression for shape functions. SIAM Journal on Mathematical Analysis, 50(1):5–32, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xu:2006:SAT

[XZCM06] Chang Fa Xu, Kai Zhang, Duan Chen, and Zhi Fang Min. δ-sequence approach to a twopoint boundary value problem using Daubechies wavelets. J. Huazhong Univ. Sci. Technol. Nat. Sci., 34(5):40–42, 2006. ISSN 1671-4512.

Yang:2012:CMS

[Yan12] Qixiang Yang. Characterization of multiplier spaces with Daubechies wavelets. Acta Mathematica Scientia. Series B. (English Edition), 32(6): 2315–2321, 2012. ISSN 0252-9602.

Yin:2016:RCA

[YCF^{+}16] Rujie Yin, Bruno Cornelis, Gabor Fodor, Noelle Ocon, David Dunson, and Ingrid Daubechies. Removing cradle artifacts in X-ray images of paintings. SIAM Journal on Imaging Sciences, 9(3):1247– 1272, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Yin:2016:DWB

[YD16] Rujie Yin and Ingrid Daubechies. Directional wavelet bases constructions with dyadic quincunx subsampling. arXiv.org, ??(??):1–36, February 17, 2016. CODEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/1602.05469.

Yin:2017:DWB

[YD17] Rujie Yin and Ingrid Daubechies. Directional wavelet bases constructions with dyadic quincunx subsampling. Journal of Fourier Analysis and Applications, ??(??):1–36, April 13, 2017. ISSN 1069-5869 (print), 1531-5851 (electronic).

Yin:2018:DWB

[YD18] Rujie Yin and Ingrid Daubechies. Directional wavelet bases constructions with dyadic quincunx subsampling. Journal of Fourier Analysis and Applications, 24(3):872–907, 2018. ISSN 1069-5869,1531-5851.

Yin:2014:DCR

[YDC^{+}14] Rujie Yin, David B. Dunson, Bruno Cornelis, Bill Brown, Noelle Ocon, and Ingrid Daubechies. Digital cradle removal in X-ray images of art paintings. In 2014 IEEE International Conference on Image Processing (ICIP), pages 4299– 4303. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 2014. ISBN 1-4799-5751-8. ISSN 1522-4880 (print), 2381-8549 (electronic).

Gao:2011:RAE

[yGjZsC11] Qiang ye Gao, Jian jiang Zhou, and Qun sheng Cao. Research and application to electromagnetic scattering of conformal MRTD method based on Daubechies scaling functions. Journal of Electronics & Information Technology, 33 (1):136–141, March 2011. ISSN 1009-5896.

Yin:2016:TTB

[YGLD16] Rujie Yin, Tingran Gao, Yue M. Lu, and Ingrid Daubechies. A tale of two bases: Local-nonlocal regularization on image patches with convolution framelets. arXiv.org, ??(??):1–51, June 4, 2016. CODEN ???? ISSN 2331-8422. URL https://

arxiv.org/abs/1606.01377.

Yin:2017:TTB

[YGLD17] Rujie Yin, Tingran Gao, Yue M. Lu, and Ingrid Daubechies. A tale of two bases: Local-nonlocal regularization on image patches with convolution framelets. SIAM Journal on Imaging Sciences, 10(2):711–750, ???? 2017. CO-DEN SJISBI. ISSN 1936-4954.

Yamada:2004:HSD

[YKIK04] Ioya Yamada, Hisashi Kubota, Akifumi Inui, and Yoshihiro Kawaguchi. High sensitive distinction of discharge in air by Daubechies wavelet transform. IEEJ Transactions on Power and Energy, 124(12): 1513–1519, 2004. ISSN 0385-4213 (print), 1348-8147 (electronic).

Yang:2015:QCW

[YLB^{+}15] Haizhao Yang, Jianfeng Lu, William P. Brown, Ingrid Daubechies, and Lexing Ying. Quantitative canvas weave analysis using 2-D synchrosqueezed transforms: Application of time-frequency analysis to art investigation. IEEE Signal Processing Magazine, 32(4):55– 63, July 2015. CODEN IS-PRE6. ISSN 1053-5888 (print), 1558-0792 (electronic).

Yin:2016:ORA

[YMH^{+}16] R. Yin, E. Monson, E. Honig, Ingrid Daubechies, and M. Maggioni. Object recognition in art drawings: Transfer of a neural network. In 2016 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), pages 2299– 2303. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, March 2016. ISBN 1-4799-9988-1. ISSN 1520-6149 (print), 2379-190X (electronic). LCCN ????

Yoshino:2010:ACA

[Yos10] Kunio Yoshino. Analytic continuation and applications of eigenvalues of Daubechies’ localization operator. Cubo, 12 (3):203–212, 2010. ISSN 0716-7776 (print), 0719-0646 (electronic).

Yoshino:2015:SAD

[Yos15] Kunio Yoshino. Spectral analysis of Daubechies localization operators. In Pseudodifferential operators and generalized functions. Selected papers of the 9th ISAAC congress, Cracow, Poland, August 5–9, 2013, volume 245 of Oper. Theory Adv. Appl., pages 285–290. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-14617-3 (hardcover); 3-319-14618-1 (e-book).

Zhang:2016:CRA

[ZCY16] Niaona Zhang, Congcong Cao, and Haifang Yu. The cycle recognition algorithm based on Daubechies wavelet and fuzzy C-means clustering. In 2016 35th Chinese Control Conference (CCC). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, July 2016.

Zeilberger:1993:ID

[Zei93] Doron Zeilberger. On an identity of Daubechies. American Mathematical Monthly, 100 (5):487, May 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/
stable/10.2307/2324306.

Zou:2004:TEA

[ZGSD04] Jing Zou, Anna Gilbert, Martin Strauss, and Ingrid Daubechies. Theoretical and experimental analysis of a randomized algorithm for sparse Fourier transform analysis. arXiv.org, ??(??):1– 21, November 4, 2004. CO-DEN ???? ISSN 2331-

                                        8422. URL https://arxiv.
                                        org/abs/math/0411102.

Zou:2006:TEA

[ZGSD06] Jing Zou, Anna Gilbert, Martin Strauss, and Ingrid Daubechies. Theoretical and experimental analysis of a randomized algorithm for Sparse Fourier Transform analysis. Journal of Computational Physics, 211(2):572–595, January 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0021999105002883.

Zhang:2000:DWS

[Zha00] Bo Jian Zhang. Daubechies wavelet solution on finite interval for the problem of acoustic scattering in a shallow ocean. J. Guangxi Univ. Nat. Sci. Ed., 25(3):182–186, 2000. ISSN 1001-7445.

Zhu:2017:LLD

[ZQH^{+}17] Wei Zhu, Qiang Qiu, Jiaji Huang, Robert Calderbank, Guillermo Sapiro, and Ingrid Daubechies. LDMNet: Low dimensional manifold regularized neural networks. arXiv.org, ?? (??):1–12, November 16, 2017. CODEN ???? ISSN 2331-

8422. URL https://arxiv.
org/abs/1711.06246.

Zhu:2019:SMD

[ZQW^{+}19] Wei Zhu, Qiang Qiu, Bao Wang, Jianfeng Lu, Guillermo Sapiro, and Ingrid Daubechies. Stop memorizing: a datadependent regularization framework for intrinsic pattern learning. SIAM Journal on Mathematics of Data Science, 1(3): 476–496, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.